### v1

# Predicting perturbation targets with causal differential networks

Abstract
Rationally identifying variables responsible for changes to a biological system can enable myriad applications in disease understanding and cell engineering. From a causality perspective, we are given two datasets generated by the same causal model, one observational (control) and one interventional (perturbed). The goal is to isolate the subset of measured variables (e.g. genes) that were the targets of the intervention, i.e. those whose conditional independencies have changed. Knowing the causal graph would limit the search space, allowing us to efficiently pinpoint these variables. However, current algorithms that infer causal graphs in the presence of unknown intervention targets scale poorly to the hundreds or thousands of variables in biological data, as they must jointly search the combinatorial spaces of graphs and consistent intervention targets. In this work, we propose a causality-inspired approach for predicting perturbation targets that decouples the two search steps. First, we use an amortized causal discovery model to separately infer causal graphs from the observational and interventional datasets. Then, we learn to map these paired graphs to the sets of variables that were intervened upon, in a supervised learning framework. This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets, each with thousands of measured variables. We also demonstrate significant improvements over six causal discovery algorithms in predicting intervention targets across a variety of tractable, synthetic datasets.
1 Introduction
Cells form the basis of biological systems, and they take on a multitude of dynamical states throughout their lifetime. In addition to natural factors like cell cycle, external perturbations (e.g. drugs, gene knockdown) can alter a cell’s state. While perturbations can affect numerous downstream variables, identifying the root causes, or targets, that drive these transitions has vast therapeutic implications, from cellular reprogramming (Cherry & Daley, 2012) to mechanism of action elucidation (Schenone et al., 2013). Large-scale experiments (Replogle et al., 2022) have attempted to map the effects of perturbing individual genes on single cells, and in-silico approaches (Roohani et al., 2023; Lotfollahi et al., 2023) have been designed for the “forward” inference task of predicting the post-perturbation expression of each gene. In principle, these models can be used within an active learning framework to suggest perturbation targets (Huang et al., 2023; Zhang et al., 2023a). However, the number of inference calls required scales exponentially with the size of the candidate set (e.g. drugs with off-target effects), rendering these approaches impractical for larger search spaces. These models’ predictive performance as oracles has also been called into question by subsequent works (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024), highlighting the difficulty of this forward approach. Finally, there are limited training data for combinatorial perturbations: the most widely used dataset only contains around a hundred pairs (Norman et al., 2019).
The “reverse” strategy – of directly predicting targets from perturbation data – alleviates these problems to an extent. Towards this end, it is common to assume (sometimes implicitly) that the data were generated by a structured causal model, and the perturbation targets can be directly inferred from the changes to this model. While the effects of perturbations are highly contextual, specific to the cell lines in which they are applied (Nadig et al., 2024), existing methods rely on non-specific, data-mined knowledge graphs as the backbones for their causal structures (Cosgrove et al., 2008; Gonzalez et al., 2024). It is hard to quantify the extent to which these graphs are relevant to each setting, as the quality and focus of assays have changed over time, and interactions present in one cell type may not exist in another (Huttlin et al., 2021).
Our goal is to predict the perturbation targets that translate one set of cells into another, while inferring the underlying causal structures directly from the data. We formalize this task within a causality framework. Specifically, given an observational dataset (e.g. gene expression of control cells) and an interventional dataset (e.g. perturbed cells) generated by the same system (e.g. cell type), we would like to identify the variables (e.g. genes) whose data-generating mechanisms are different between the two settings. Causal discovery algorithms have been designed to solve this exact problem, on interventional data with unknown targets (Squires et al., 2020; Brouillard et al., 2020; Hägele et al., 2023). These algorithms operate jointly over the spaces of graphs and intervention target sets. Due to this large search space, these algorithms are unable to scale to the thousands of variables and hundreds of thousands of observations in modern transcriptomics data, while simultaneously being robust to as few as tens of observations per interventional setting (Nadig et al., 2024).
To address unknown intervention target prediction in context of biological perturbations, we propose causal differential networks (Cdn). Drawing upon recent advances in scalable and data-efficient causal discovery (Wu et al., 2024), Cdn first “featurizes” the observational and interventional datasets in terms of their predicted causal graphs, using a pretrained causal discovery module. These paired, edge-level representations are then used as input to an axial attention-based classifier (Ho et al., 2020), which is supervised by ground truth intervention targets (on either real or synthetic data). We thoroughly evaluate Cdn on both real transcriptomic data and in synthetic settings. Cdn outperforms the state-of-the-art in perturbation modeling on the five largest Perturb-seq datasets at the time of publication (CRISPRi,  Replogle et al. (2022); Nadig et al. (2024)), and two Sci-Plex datasets (chemical perturbations, McFaline-Figueroa et al. (2024)), without using any external knowledge. In fact, Cdn is the only model that consistently ranks the ground truth perturbation targets higher than would be expected by random. Moreoever, Cdn is able to maintain decent performance even on unseen cell lines, demonstrating its potential to generalize. Finally, on synthetic settings, Cdn outperforms causal discovery algorithms for unknown intervention targets. To conclude, our contributions are three-fold.
  1. 1.
    We propose causal differential networks (Cdn), a causal discovery-based approach for perturbation target prediction on transcriptomic data, which achieves the state-of-the-art on genetic and chemical perturbation datasets.
  2. 2.
    We demonstrate that Cdn outperforms current causal discovery approaches for unknown intervention target prediction for both hard and soft interventions in diverse synthetic data.
  3. 3.
    Finally, we have prepared seven high-quality transcriptomics datasets, to serve as a benchmark for future machine learning studies.
Figure 1: (A) Biological applications. (B) Differences between observation dataset $D_{\text{obs}}$ and interventional dataset $D_{\text{int}}$ can be attributed to changes in the underlying causal model $G$. (C) We use an amortized causal discovery module to obtain features that represent $G_{\text{obs}}$ and $G_{\text{int}}$. Then we train a supervised classifier to predict the unknown intervention targets.
2 Background and related work
2.1 Identifying perturbation targets
The rise of large-scale perturbation screens (Replogle et al., 2022) has enabled machine learning approaches for predicting perturbation targets. One line of work focuses on active learning to reduce experimental costs (Zhang et al., 2023a; Huang et al., 2023), but these models require inference calls to an oracle model for each candidate perturbation. Given that human cells express over ten thousand genes, and drugs may impact multiple targets, the search space is enormous. Moreover, subsequent works have found that naive baselines (mean of the training set, linear regression) outperform these approaches at their intended task (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024). In contrast, perturbation targets can be predicted directly from data, using a variety of domain knowledge (Cosgrove et al., 2008; Gonzalez et al., 2024; Roohani et al., 2024). However, these external data were collected from highly inhomogenous sources, which may be inconsistent with the data at hand.
2.2 Causal discovery with unknown intervention targets
Perturbation target prediction can be formalized as a causal discovery problem, using data generated under interventions with unknown targets. Specifically, a causal graphical model is a directed graph $G=(V,E)$, where nodes $i\in V$ map to random variables $X_{i}\in X$, and edges $(i,j)\in E$ represent relationships from $X_{i}$ to $X_{j}$. There are a number of common assumptions that relate $G$ to the data distribution $P_{X}$, which we defer to works such as Spirtes et al. (2001); Yang et al. (2018); Zhang et al. (2023b), since the identifiability of any particular system is not the focus of this paper.
Data generated directly from $P_{X}$ are known as observational data. A causal graphical model allows us to perform interventions by assigning new conditionals
$P(X_{i}\mid X_{\pi_{i}})\leftarrow\tilde{P}(X_{i}\mid X_{\pi_{i}}),$ (1)
where $\pi_{i}$ denotes the parents of node $i$ in $G$. Hard interventions remove all dependence between $X_{i}$ and $\pi_{i}$, while soft interventions maintain the relationship with a different conditional. We denote the joint interventional distribution as $\tilde{P}_{X}$. Given an observational dataset $D_{\text{obs}}\sim P_{X}$ and an interventional dataset $D_{\text{int}}\sim\tilde{P}_{X}$, our goal is to predict the set of nodes $I$ for which
$P(X_{i}\mid X_{\pi_{i}})\neq\tilde{P}(X_{i}\mid X_{\pi_{i}}),\forall i\in I.$ (2)
This task has been well-studied in the causality literature, often as a sub-problem or setting while inferring graph $G$ from $D_{\text{obs}}$ and (multiple) $D_{\text{int}}$. Ke et al. (2019) first proposed a discovery algorithm for unknown interventions, on discrete data of up to 50 variables. Jaber et al. (2020) provides identifiability guarantees for soft interventions with unknown targets and proposes a constraint-based algorithm, which requires an exponential number of conditional independence tests with respect to the number of variables. Continuous causal discovery algorithms that are consistent despite unknown interventions include Dcdi (Brouillard et al., 2020) and BaCaDI (Hägele et al., 2023), which treat intervention target identities as learnable parameters of a generative model over $P_{X}$. Dcdi uses the Gumbel-Softmax approximation (Jang et al., 2017) to learn these categorical variables, while Bacadi takes a Bayesian approach, with gradient-based methods for efficient posterior estimation (Liu & Wang, 2016). Finally, Varici et al. (2022) and Yang et al. (2024) frame intervention target prediction as a separate task, similar to this paper. However, the former strictly assumes linearity, while the latter requires data from multiple environments, which are not available here.
2.3 Amortized causal discovery algorithms
The majority of causal discovery algorithms operate on one or more datasets $D$, which correspond to a single graph $G$ (Spirtes et al., 1995; Hauser & Bühlmann, 2012; Mooij et al., 2020; Brouillard et al., 2020). These algorithms must be fit or run from scratch for each data distribution – a potential challenge in low-resource scenarios, where there are too few observations of too many variables. Recently, amortized causal discovery algorithms have addressed causal structure learning as a supervised machine learning problem. Using large numbers of synthetic datasets, generated by (known) synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Ke et al., 2022; Lorch et al., 2022; Wu et al., 2024). Two immediate benefits are fast runtimes and robustness to aleatoric noise. In this work, we leverage the framework proposed by Wu et al. (2024), as it has been benchmarked on large, transcriptomics data (Chevalley et al., 2022).
3 Methods
Our model is composed of two modules: a frozen causal featurizer, and a learned differential network, which we refer to jointly as Cdn (Figure 2). The causal featurizer is a pretrained, amortized causal discovery model (Wu et al., 2024) that runs efficiently on up to a thousand nodes (Section 3.1), while the differential network is an axial-attention based classifier (Ho et al., 2020) that predicts the nodes whose data-generating mechanisms have changed (Section 3.2).
3.1 Causal featurizer
The causal featurizer takes as input datasets $D_{\text{obs}}$ and $D_{\text{int}}$ containing samples of $N$ variables, generated from observational and interventional settings. For example, $D_{\text{obs}}$ may correspond the gene expression matrix of non-targeting control cells; $D_{\text{int}}$ is the gene expression of the same type of cells, subject to a perturbation with an unknown target; and $N$ is the number of genes detected.
We use an amortized causal discovery model to obtain features $h_{\text{obs}},h_{\text{int}}\in\mathbb{R}^{N\times N\times d}$, which reflect the $N\times N$ adjacency matrices of causal graphs that may have generated each dataset. Specifically, to featurize each dataset, we follow the three steps proposed in Wu et al. (2024).
  1. 1.
    Given a dataset $D$, smaller batches are constructed by sub-sampling both examples and variables. Heuristics like pairwise correlation are used to select variables which are likely to have causal relationships, to minimize unnecessary computation.
  2. 2.
    Two sets of input features are generated from these batches: global summary statistics like correlation, computed over all variables, and marginal estimates, the outputs of classical causal discovery algorithms (Spirtes et al., 1995) run on small subsets of variables.
  3. 3.
    Finally, these two sets of features are input to a neural network (“aggregator”), which was trained to map them into causal graphs on synthetic datasets.
We run this procedure on $D_{\text{obs}}$ and $D_{\text{int}}$ independently, treating both as “observational” datasets. To obtain the graph representations for each dataset, we extract the aggregator’s last layer hidden representations before they are collapsed into binary edge predictions, yielding $h_{\text{obs}}$ and $h_{\text{int}}$. Finally, we concatenate $h_{\text{obs}}$ and $h_{\text{int}}$ along the hidden dimension, to obtain a paired graph representation of size $N\times N\times 2d$.
Figure 2: Model overview. The observational and interventional datasets are featurized using a pretrained, amortized causal discovery model. The pair of graph representations are concatenated and input to an axial attention model. After the 2D attention, representations are aggregated over incoming edges to obtain binary, node-level predictions. Cdn is trained with a classification objective, over synthetic and/or real data. Gray modules are frozen. Yellow modules are learned.
3.2 Differential network
Given the graph representation, the differential network predicts which nodes were intervened upon (Figure 2, right). Its architecture is composed of an axial-attention layer, followed by a linear projection. Following Sea, the axial attention layer attends separately along the rows and columns of the graph’s adjacency matrix. This operation is equivalent to self-attention along all nodes in the “outgoing edge” direction, with the “incoming edges” as a batch dimension, followed by the opposite. We use pre-layer normalization on each self-attention, followed by dropout and residual connections:
$h=h+\text{Dropout}(\text{Self-Attn}(\text{LayerNorm}(h))).$ (3)
After the axial attention, we mean over all “incoming edges” and linearly project to binary node-level predictions, where a normalized 1 indicates an intervention target, and 0 indicates otherwise.
3.3 Implementation and training details
For the causal featurizer, we used a frozen, pretrained aggregator from (Wu et al., 2024), which corresponded to the FCI-based marginal estimates (Spirtes et al., 1995). The differential network was implemented with one axial attention layer, following the same architecture as the pretrained aggregator, with twice the model dimension (paired graphs). We swept over the number of layers (1-4) on synthetic data and found that a single layer performed the best. In terms of model design, we ablated replacing the attention architecture with a multi-layer perceptron (no communication between nodes, Table 4), as well as other components (Table 7) and summary statistics (Table 8). Due to the large graph sizes (adjacency matrices up to $1000^{2}$), standard 1D attention was intractable. Following Sea, we used inverse covariance as the global statistic on synthetic data, and correlation on transcriptomics data.
We trained separate differential networks on synthetic and real data. On synthetic data, we used a batch size of 16 and the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 1e-4 (identical to Sea). On real data, we started from the synthetic Cdn checkpoint, and then used a batch size of 2 with a learning rate of 1e-5, due to the larger graph sizes. All models took around 4-10 hours to train on a single A6000 GPU.
3.4 Theoretical context
This paper primarily describes a causality-inspired approach towards perturbation target prediction. Contrary to causal discovery algorithms that are self-contained and provably sound under standard assumptions, we cannot “prove” that a pretrained model extracts correct graphs on real data, as the true graphs are unknown. Our evaluation focuses solely on perturbation target prediction, as the ground truth is known, by (experimental) design. To contextualize the predictions of our model, we provide brief sketches with regards to our model’s capacity to map paired graph features to intervention targets. We defer formal claims to Appendix A.
Claim 3.1 (Hard interventions, informal).
Given a causal graphical model $G=(V,E)$ and intervention targets $I\subseteq V$, let $E^{\prime}$ denote the adjacency matrix of mutilated graph $G^{\prime}$ after hard intervention on $I$. An axial attention layer is well-specified to map $(E,E^{\prime})$ to $I$.
Claim 3.2 (Soft interventions, informal).
Let $G$ be a graphical causal model associated with data distribution $P_{X}$, and let $\tilde{P}_{X}$ be the interventional distribution after soft intervention on $I\subsetneq V$. Let $R,R^{\prime}$ denote the correlation matrices and let $\Sigma,\Sigma^{\prime}$ denote the covariance matrices of $X,\tilde{X}$. An axial attention layer is well-specified to map $(R,R^{\prime},\Sigma,\Sigma^{\prime})$ to $I$.
4 Experiments and results
Table 1: Final biological dataset statistics. “gw” and “es” are two independent screens with different perturbations.
Type Cell line ##\## Perts ##\## Cells
Genetic K562 gw 1767 492,096
K562 es 1060 213,552
RPE1 961 179,696
HepG2 627 79,309
Jurkat 1014 174,698
Chemical A172 3 18,196
T98G 3 13,126
Figure 3: Data illustration and splitting procedure.
We evaluated Cdn on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1). Note that these baselines all use various domain knowledge to inform their predictions. For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
4.1 Biological experiments
Datasets
We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al. (2022) and Nadig et al. (2024); as well as two Sci-Plex (Srivatsan et al., 2020) datasets (chemical perturbations) from McFaline-Figueroa et al. (2024). Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$ (Figure 3A). In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target.
To ensure high quality labels, we filtered perturbations to those that induced over 10 differentially-expressed genes (statistically significant change, compared to control), of which the true target should be present. This is to exclude perturbations with insufficient cells (low statistical power), with minimal to no effect (uninteresting), and those that did not achieve the desired effect (CRISPR efficacy is not guaranteed). For evaluation, we limited the set of candidate targets to the top 1000 differentially expressed genes, ranked by log-fold change per perturbation. If fewer genes were differentially expressed, we randomly sampled additional candidates until we reached a minimum of 100 genes. Finally, we also stratify genetic perturbations based on whether the target is trivially identifiable as the gene with the largest log-fold change (“trivial” vs. “non-trivial”). This is because genetic perturbations are highly specific by design, but also constitute the largest single-cell perturbation datasets. The final data statistics are shown in Table 1, with additional details in Appendix B.1.
Evaluation
We consider two splits: seen and unseen cell lines (Figure 3B). In the former, models may be trained on approximately half of the perturbations from each cell line, and are evaluated on the unseen perturbations. In the latter, we hold out one cell line at a time, and models may be trained on data from the remaining cell lines. Note that not all baselines can be evaluated on unseen cell lines. To ensure that our train and test splits are sufficiently distinct, we cluster perturbations based on their log-fold change and assign each cluster to the same split.
While perturbation target prediction appears to be a simple classification task, there are two aspects of the data that render standard metrics less meaningful. Not all genes can be mapped to the baselines’ domain knowledge, so their effects as perturbations cannot be predicted. In addition, due to genetic redundancy, it is common for multiple perturbations to elicit similar responses (Kernfeld et al., 2023). Therefore, an “incorrectly” predicted target (e.g. out of 1000 genes) may not necessarily indicate poor performance. Therefore, we propose the following metrics.
  • We report the rank of the ground truth target, normalized by the number of candidate genes. Here, 1 indicates that the ground truth target is at the top of the list (best), while 0 indicates that the ground truth is at the bottom (worst).
  • To emulate a “virtual screening” setting, we report recall at k (the proportion of targets recovered within the top $k$ candidates). This value ranges from 0 to 1, where 1 is best.
  • The Pearson correlation $r$ measures the similarity between the (ground truth) mean log-fold change of the top prediction that was observed as a perturbation, and that of the actual perturbation (given as input). These values range from -1 to 1, where 1 is best.
Baselines
We compare to two classes of algorithms related to biological perturbations. Forward inference models predict the effects of perturbations, and perturbation targets can be ranked by comparing the inferred effect of each candidate to the interventional data.
  • Gears (Roohani et al., 2023) is a graph neural network that predicts the effects of unseen genetic perturbations, using the gene ontology knowledge graph (Ashburner et al., 2000) to model the relationship between genes. We trained Gears on each cell line separately, and predicted the effects of perturbing every expressed gene. Then we ranked candidates based on cosine similarity to the interventional data.
  • GenePt (Chen & Zou, 2023) is a set of natural language-based embeddings that have been shown to achieve state-of-the-art performance on unseen perturbation effect prediction with simple downstream models (Märtens et al., 2024). We concatenated each perturbation’s gene embedding to the log-fold change of the top 5000 highly-variable genes (Wolf et al., 2018) and trained a logistic regression model on each cell line to predict true vs. decoy perturbations. Candidates were ranked based on predicted probability.
Predictive models directly predict perturbation targets from data as a classification task.
  • Linear and Mlp take as input the mean expression of all perturbation targets, plus the top 2000 highly-variable genes (Wolf et al., 2018). They are trained to predict a binary label for each gene on each cell line independently.
  • PDGrapher (Gonzalez et al., 2024) is a graph neural network that predicts the perturbation targets of genetic or chemical perturbations, on seen or unseen cell lines, with the human reference interactome (Luck et al., 2020) as the knowledge graph.
All baselines were run with their official implementations and/or latest releases. For more details, please see Appendix C.1.
Table 2: Results on 5 Perturb-seq datasets. Top: “trivial” perturbations. Bottom: “non-trivial” perturbations. Primary setting: train all cell lines jointly, test on unseen perturbations. Suffix -cl: leave out the one cell line from training, test on the same set of unseen perturbations in the unseen cell line. Metrics, from left to right: normalized rank of ground truth, top 20 recall, and top 1 Pearson correlation. For ablations, runtimes, and uncertainties, see Appendix D.
K562 gw K562 es RPE1 HepG2 Jurkat
Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
Random $0.50$ $0.13$ $0.16$ $0.50$ $0.09$ $0.17$ $0.50$ $0.09$ $0.36$ $0.50$ $0.06$ $0.34$ $0.50$ $0.08$ $0.22$
Linear $0.49$ $0.16$ $0.21$ $0.50$ $0.13$ $0.31$ $0.48$ $0.09$ $0.45$ $0.48$ $0.07$ $0.43$ $0.48$ $0.08$ $0.29$
Mlp $0.48$ $0.21$ $0.20$ $0.47$ $0.13$ $0.21$ $0.54$ $0.19$ $0.37$ $0.51$ $0.10$ $0.42$ $0.48$ $0.13$ $0.32$
GenePt $0.60$ $0.39$ $0.39$ $0.47$ $0.11$ $0.26$ $0.57$ $0.33$ $0.55$ $0.44$ $0.12$ $0.41$ $0.55$ 0.32 0.53
Gears $0.49$ $0.14$ $0.18$ $0.51$ $0.09$ $0.27$ $0.49$ $0.09$ $0.40$ $0.48$ $0.09$ $0.40$ $0.47$ $0.06$ $0.24$
PdG $0.49$ $0.20$ $0.26$ $0.52$ $0.18$ $0.34$ $0.51$ $0.11$ $0.42$ $0.46$ $0.06$ $0.40$ $0.49$ $0.11$ $0.35$
Cdn 0.77 0.52 0.49 0.68 0.32 0.45 0.74 0.37 0.55 0.68 0.20 0.50 0.67 $0.27$ $0.45$
PdG-Cl $0.40$ $0.04$ $0.15$ $0.41$ $0.02$ $0.24$ $0.48$ $0.02$ $0.37$ $0.47$ $0.03$ $0.41$ $0.53$ $0.06$ $0.33$
Cdn-Cl 0.71 0.43 0.40 0.66 0.27 0.42 0.62 0.24 0.53 0.66 0.17 0.46 0.65 0.26 0.47
Random $0.50$ $0.11$ $0.17$ $0.50$ $0.04$ $0.17$ $0.50$ $0.04$ $0.44$ $0.50$ $0.04$ $0.33$ $0.50$ $0.06$ $0.21$
Linear $0.51$ $0.18$ $0.23$ $0.48$ $0.01$ $0.25$ $0.49$ $0.03$ $0.52$ $0.51$ $0.04$ 0.43 $0.50$ $0.06$ $0.23$
Mlp $0.45$ $0.19$ $0.21$ $0.49$ $0.04$ $0.11$ $0.53$ $0.09$ $0.41$ $0.48$ 0.07 $0.40$ $0.45$ $0.07$ $0.29$
GenePt $0.52$ $0.24$ $0.32$ $0.42$ $0.03$ $0.19$ $0.41$ $0.06$ $0.46$ $0.31$ $0.03$ $0.40$ $0.51$ 0.15 $0.36$
Gears $0.56$ $0.16$ $0.22$ $0.45$ $0.06$ $0.20$ $0.49$ $0.04$ $0.44$ $0.54$ $0.03$ $0.40$ $0.50$ $0.11$ $0.26$
PdG $0.54$ $0.24$ $0.29$ $0.49$ $0.03$ $0.23$ $0.48$ $0.05$ $0.52$ $0.54$ $0.05$ 0.43 $0.52$ $0.10$ $0.35$
Cdn 0.72 0.37 0.32 0.56 0.15 0.35 0.59 0.12 0.53 0.55 0.07 $0.40$ 0.62 $0.10$ 0.39
PdG-Cl $0.39$ $0.02$ $0.13$ $0.44$ $0.01$ 0.28 0.54 $0.02$ 0.48 $0.48$ $0.02$ 0.41 $0.57$ $0.06$ $0.35$
Cdn-Cl 0.68 0.25 0.29 0.55 0.06 $0.27$ $0.49$ 0.07 0.48 0.59 0.05 $0.39$ 0.60 0.18 0.38
Figure 4: Recall at $k$ on K562 datasets.
Table 3: Rank of drug targets on chemical perturbation datasets (McFaline-Figueroa et al., 2024). A172 and T98G are unseen cell lines. For details and full drug names, please see Appendix B.1.
A172 T98G
Model infig. nint. palb. doxo. palb. vola.
PdG $0.719$ $0.346$ $0.395$ $0.055$ $0.345$ $0.420$
Cdn 0.883 0.436 0.582 0.059 0.543 0.647
Results
Table 2 reports our results on five Perturb-seq datasets. In the majority of cases, Cdn outperforms baselines, both in the seen and unseen cell line settings. This is more evident in Figure 4, in which Cdn achieves higher recall at $k$ at nearly all points on the curves. While no baseline is as consistent as Cdn, GenePt is surprisingly competitive against more complex baselines. This may indicate that natural language induces embedding spaces with favorable geometries with respect to biological function. On the two chemical perturbation datasets, Cdn ranks the ground truth targets higher than PdGrapher on all six drugs (Table 3). However, the mutually low performance on doxorubicin (doxo.) in T98G cells may indicate a failure mode of both models.
4.2 Synthetic experiments
There are two considerations that motivate further experiments in controlled settings. First, while it would be ideal to evaluate all algorithms on real data, current causal discovery algorithms that support unknown interventions are not tractable on datasets with more than tens of variables. In fact, many baselines require hours on even $N=10$ datasets, and they do not scale favorably (Table 9). Our transcriptomics datasets contain hundreds of genes, even after filtering to those that are differentially expressed (Figures 6 and 7). On the other hand, existing models for biological perturbations all rely on some form of domain knowledge, and their performance is inseparable from the choice and quality of these external data. Here, synthetic settings allow us to assess the model’s capacity to predict intervention targets in isolation.
Datasets
We generated 120 synthetic test datasets with hard and soft interventions. To generate observational data, we sampled Erdős-Rényi graphs with $N=10,20$ nodes and $E=N,2N$ expected edges; causal mechanism parameters; and observations of each variable, in topological order. We sampled $3N$ distinct subsets of 1-3 nodes ($N$ each) as intervention targets. For hard interventions, we set $x\leftarrow z$, where $z$ is uniform. To emulate transcriptomics data, in which perturbation effects are measured in fold change, we introduce soft interventions: $x\leftarrow cf(\pi_{x})$, where $c$ is a positive scaling factor, with equal probability $c\lessgtr 1$. For our synthetic training set, we generated approximately 4000 datasets following the training set of Sea, with hard interventions only. Full details are available in Appendix B.2.
Baselines
We compare against discrete and continuous causal discovery algorithms for unknown interventions. Ut-Igsp (Squires et al., 2020) infers causal graphs and unknown targets by greedily selecting the permutation of variables that minimizes their proposed score function. Dcdi (Brouillard et al., 2020) and Bacadi (Hägele et al., 2023) are continuous causal discovery algorithms that fit generative models to the data, where the causal graph and intervention targets are model parameters. The -G and -Dsf suffixes on Dcdi correspond to Gaussian and deep sigmoidal flow parametrizations of the likelihood. Bacadi is evaluated here using the fully-connected implementation (better performance), with the linear version in Table 11. Its -E and -M suffixes indicate empirical (standard) and mixture (bootstrap) variants.
We also compare against the difference causal inference (Dci) algorithm with stability selection (Belyaeva et al., 2021), a discrete optimization algorithm that aims to detect edge-level differences between two causal graphs. On this node-centric task, we take each node’s proportion of changed edges as its likelihood of being an intervention target. While Dci was also motivated by biological applications, it only scales to around a hundred nodes at most, so we evaluate it alongside other causal discovery methods here.
Evaluation
In the synthetic case, we are not constrained by biological redundancy or incomplete predictions, so we report standard classification metrics.
  • Mean average precision (mAP) is the average precision over a continuum of binarization thresholds, computed independently for each variable, averaged over all regimes of the same number of targets. This ranges from 0 to 1 (perfect), where the random baseline depends on the positive rate.
  • Area under the receiver operating curve (AUC) is also computed independently for each variable, averaged over the same set. This ranges from 0 to 1 (perfect), where the random baseline is 0.5 (per edge).
Results
Table 4: Intervention target prediction results on synthetic datasets with $N=10,E=10$. Uncertainty is standard deviation over 5 distinct datasets. Top: hard interventions; bottom: soft interventions (“scale”). Number in parentheses indicates number of intervention targets. Runtimes are documented in Table 9, and full results are available in Table 11.
Linear (1) Linear (3) Polynomial (1) Polynomial (3)
[…]
Hard Ut-Igsp $0.19\scriptstyle\pm.03$ $0.56\scriptstyle\pm.05$ $0.38\scriptstyle\pm.04$ $0.57\scriptstyle\pm.05$ $0.27\scriptstyle\pm.05$ $0.63\scriptstyle\pm.03$ $0.38\scriptstyle\pm.03$ 0.55±.03plus-or-minus0.55.030.$5$03
Dcdi-G $0.38\scriptstyle\pm.12$ $0.59\scriptstyle\pm.11$ $0.47\scriptstyle\pm.03$ $0.53\scriptstyle\pm.03$ $0.27\scriptstyle\pm.09$ $0.47\scriptstyle\pm.07$ $0.45\scriptstyle\pm.04$ $0.49\scriptstyle\pm.05$
Dcdi-Dsf $0.32\scriptstyle\pm.05$ $0.58\scriptstyle\pm.10$ $0.43\scriptstyle\pm.03$ $0.48\scriptstyle\pm.03$ $0.25\scriptstyle\pm.04$ $0.42\scriptstyle\pm.05$ $0.46\scriptstyle\pm.03$ $50$05
Dci $0.42\scriptstyle\pm.07$ $0.75\scriptstyle\pm.03$ $0.63\scriptstyle\pm.11$ $80$08 $0.39\scriptstyle\pm.10$ $0.72\scriptstyle\pm.06$ $0.53\scriptstyle\pm.09$ $0.69\scriptstyle\pm.07$
BaCaDI-E $0.16\scriptstyle\pm.04$ $0.61\scriptstyle\pm.09$ $0.36\scriptstyle\pm.04$ $60$06 $0.13\scriptstyle\pm.02$ $0.62\scriptstyle\pm.05$ $0.36\scriptstyle\pm.03$ $0.60\scriptstyle\pm.04$
BaCaDI-M $0.12\scriptstyle\pm.01$ $0.57\scriptstyle\pm.05$ 0.33±.02plus-or-minus0.33.020.$3$02 $0.56\scriptstyle\pm.04$ $0.12\scriptstyle\pm.01$ $0.57\scriptstyle\pm.04$ 0.33±.02plus-or-minus0.33.020.$3$02 0.55±.04plus-or-minus0.55.040.$5$04
[…]
Soft Ut-Igsp $20$02 $0.71\scriptstyle\pm.04$ $0.37\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $0.24\scriptstyle\pm.03$ $0.72\scriptstyle\pm.03$ $0.39\scriptstyle\pm.04$ $0.58\scriptstyle\pm.05$
Dcdi-G $0.38\scriptstyle\pm.11$ $0.58\scriptstyle\pm.11$ $0.48\scriptstyle\pm.05$ $0.50\scriptstyle\pm.07$ $0.25\scriptstyle\pm.06$ $0.49\scriptstyle\pm.08$ $0.48\scriptstyle\pm.05$ $0.53\scriptstyle\pm.06$
Dcdi-Dsf $0.37\scriptstyle\pm.09$ $0.59\scriptstyle\pm.11$ $0.45\scriptstyle\pm.05$ $0.51\scriptstyle\pm.06$ $0.32\scriptstyle\pm.09$ $0.53\scriptstyle\pm.13$ $0.43\scriptstyle\pm.04$ $0.47\scriptstyle\pm.04$
Dci $0.45\scriptstyle\pm.10$ $\textbf{0.73}\scriptstyle\pm.07$ $0.54\scriptstyle\pm.04$ $\textbf{0.68}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.03$ $0.69\scriptstyle\pm.03$ $0.45\scriptstyle\pm.03$ $0.62\scriptstyle\pm.01$
BaCaDI-E $0.34\scriptstyle\pm.16$ $0.63\scriptstyle\pm.10$ $0.49\scriptstyle\pm.14$ $0.62\scriptstyle\pm.15$ 0.22±.03plus-or-minus0.22.030.$2$03 $0.64\scriptstyle\pm.05$ $0.38\scriptstyle\pm.06$ $0.58\scriptstyle\pm.06$
BaCaDI-M $0.28\scriptstyle\pm.11$ $0.62\scriptstyle\pm.10$ 0.44±.14plus-or-minus0.44.140.$4$14 $0.59\scriptstyle\pm.15$ $0.21\scriptstyle\pm.02$ $0.64\scriptstyle\pm.05$ $0.38\scriptstyle\pm.06$ $0.58\scriptstyle\pm.06$
Cdn (Mlp) $0.29\scriptstyle\pm.04$ $0.53\scriptstyle\pm.05$ 0.44±.04plus-or-minus0.44.040.$4$04 $0.47\scriptstyle\pm.02$ 0.33±.09plus-or-minus0.33.090.$3$09 $0.50\scriptstyle\pm.09$ $0.52\scriptstyle\pm.06$ $0.56\scriptstyle\pm.05$
[…]
On synthetic data, Cdn achieves high performance across intervention types and data-generating mechanisms (Table 4), while running in seconds (Table 9). As an ablation study, we investigated replacing the differential network with a multi-layer perceptron, which does not model interactions between edges. On the hard interventions, the Mlp is sufficient for predicting intervention targets, perhaps by abusing marginal variances (Reisach et al., 2021) or other node-level artifacts of synthetic data. Once we consider soft interventions, however, the original Cdn formulation significantly outperforms the Mlp variant – indicating that graph-level information is essential.
Dci is the best baseline, which is encouraging, as it was designed for the edge-centric version of our task, with the intuition that detecting edge differences is easier than reproducing the entire causal graph. It performs particularly well on linear data, which align with its assumptions, and less well on non-linear data. Surprisingly, most other baselines perform poorly at recovering intervention targets. In the case of Dcdi and BaCaDI, this may be because it is hard to select a single sparsity threshold for varying sizes of intervention sets, and to balance the regularizer with the generative modeling objective.
5 Conclusion
We presented Cdn, a causal discovery-powered approach for predicting perturbation targets in the context of single-cell transcriptomics. Cdn uses an amortized causal discovery algorithm to represent a pair of observational and interventional datasets in terms of their data-generating mechanisms, and then trains a classifier to identify variables whose conditional independencies have changed. Our approach achieves the state-of-the-art in predicting perturbation targets over seven transcriptomics datasets, compared to a variety of perturbation modeling baselines. Cdn also surpasses causal discovery algorithms for predicting unknown targets across synthetic settings, while only requiring a fraction of the runtime. Nonetheless, there is substantial space for improvement, so we hope that this work and the associated datasets will inform future method development for these biological applications.
Reproducibility statement
All source code will be made publicly available. Hyperparameter and implementation choices are described in Section 3.3. Details regarding biological data pre-processing can be found in Section 4.1 and Appendix B.1. The raw data are freely available via the accession codes listed in Table 5. Synthetic data generation is discussed in Appendix B.2. Details required to reproduce baselines are described in Sections 4.1 and 4.2, as well as Appendix C.1.
Ethics statement
We use publicly available datasets which do not involve human subjects. While our work has applications to biological understanding, our work does not directly concern the design of molecules or other potentially harmful agents.
Acknowledgements
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 1745302. We would like to acknowledge support from the NSF Expeditions grant (award 1918839: Collaborative Research: Understanding the World Through Code), Machine Learning for Pharmaceutical Discovery and Synthesis (MLPDS) consortium, and the Abdul Latif Jameel Clinic for Machine Learning in Health.
Appendix A Theoretical intuition
While this paper focuses on biological applications, we would like to show that our model is well-specified, i.e. it has the capacity to predict correct intervention targets, given certain inputs.
Preliminary note
Continuous causal discovery works typically rely on the universality of their architectures for consistency (Brouillard et al., 2020), so this analysis also focuses on computational capacity. Theoretically, it is impossible to “guarantee” that high-dimensional representations from learned model contain or do not contain certain information. This can be tested empirically through probing, e.g. querying for grammatical structure in language models (Vulić et al., 2020), but cannot be “proven” in the traditional sense. Thus, we believe that the empirical performance of our model on real datasets is its primary contribution, but provide these explanations for interested readers. In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$. The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2024). We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
Axial attention architecture
Our differential network is implemented using an axial attention layer, which is composed of two self-attention layers (one along each axis of the adjacency matrix) and a feed-forward network. For simplicity, we follow prior work Yun et al. (2019) and ignore layer normalization and dropout.
Our inputs $h\in\mathbb{R}^{2d\times N\times N}$ are $2d$-dimension features, which represent a pair of $N\times N$ causal graphs. We use $h_{\cdot,j}$ to denote a length $N$ row for a fixed column $j$, and $h_{i,\cdot}$ to denote a length $N$ column for a fixed row $i$. The axial attention layer implements:
$\text{Attn}_{\text{row}}(h_{\cdot,j})$ $=h_{\cdot,j}+W_{O}W_{V}h_{\cdot,j}\cdot\sigma\left[(W_{K}h_{\cdot,j})^{T}W_{Q}h_{\cdot,j}\right],$
$\text{Attn}_{\text{col}}(h_{i,\cdot})$ $=h_{i,\cdot}+W_{O}W_{V}h_{i,\cdot}\cdot\sigma\left[(W_{K}h_{i,\cdot})^{T}W_{Q}h_{i,\cdot}\right],$
$\text{FFN}(h)$ $=h+W_{2}\cdot\text{ReLU}(W_{1}\cdot h+b_{1})+b_{2},$
where $W_{O}\in\mathbb{R}^{2d\times 2d},W_{V},W_{K},W_{Q}\in\mathbb{R}^{2d\times 2d},W_{2}\in\mathbb{R}^{2d\times m},W_{1}\in\mathbb{R}^{m\times 2d},b_{2}\in\mathbb{R}^{2d},b_{1}\in\mathbb{R}^{m}$, and $m$ is the FFN hidden dimension. We have omitted the $i$ and $j$ subscripts on the $W$s, but they use separate parameters. Any self attention can take on the identity mapping by setting $W_{O},W_{V},W_{K},W_{Q}$ to $2d\times 2d$ matrices of zeros.
Hard interventions
Let $G=(V,E)$ be a causal graphical model associated with data distribution $P_{X}$. Let $G^{\prime}=(V,E^{\prime})$ and $\tilde{P}_{X}$ denote the causal graph and data distribution after an unknown intervention, with ground truth targets $I\subsetneq V$. For convenience, we use $E,E^{\prime}$ both to denote sets of edges, as well as the equivalent adjacency matrices.
In the case of perfect interventions,
$f(x_{i})\leftarrow z_{i},\forall i\in I$ (4)
where $z_{i}\perp\!\!\!\perp X$ are independent random variables. $\tilde{P}_{X}$ is associated with mutilated graph $E^{\prime}$, where
$E^{\prime}=E\setminus\bigcup_{i\in I}\{(j,i)\}_{(j,i)\in E}.$ (5)
In terms of the associated adjacency matrices, $E-E^{\prime}$ has 1s in each column $i\in I$ and 0s elsewhere.
Here, the axial attention layer should implement $h-h^{\prime}$, so that when we collapse over the incoming edges, the output is non-zero only at the edge differences. Suppose the first dimension of the $2d$ feature stores $E$, and the second dimension stores $E^{\prime}$. The row self-attention implements the identity (in the first two dimensions). Then we can set $W_{K,Q}$ to zero, $W_{V}$ to the identity, and $W_{O}$ to
$W_{O}=\left[\begin{array}[]{cc}1&-1\\ 0&0\end{array}\right]-\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right]$ (6)
to account for the residual. The FFN implements the identity, so that when we take the mean over along the rows, we recover non-zero elements at all nodes whose incoming edges were removed.
Soft interventions
We also study soft interventions the context of causal models with non-multiplicative noise, in which intervention targets are scaled by constant factors,
$f(x_{i})\leftarrow c_{i}f(x),c_{i}>0$ (7)
where $c_{i}$ are sampled at random per synthetic dataset. This choice is inspired by the fact that biological perturbation effects are measured in fold-change. Here, the adjacency matrices are the same, but global statistics differ. In particular, we focus on two statistics: the correlation matrix $R$ and the covariance matrix $\Sigma$. Note that while these inputs differ slightly from our main experiments (reasons discussed in Appendix D), we show that they still achieve reasonable performance in Table 8.
Suppose $x$ is an intervention target.
  • $R-R^{\prime}$ is non-zero in all entries $i,j$ and $j,i$ where $i$ is a descendent of $x$, and $j$ is any node for which $R_{i,j}\neq 0$ (e.g. ancestors, descendants, and $x$, if $P_{X}$ is faithful to $G$).
  • $\Sigma-\Sigma^{\prime}$ is non-zero in all entries $i,j$ and $j,i$ where $i$ is a descendent of $x$ or $i=x$, and $j$ is any node for which $\Sigma_{i,j}\neq 0$ (e.g. ancestors, descendants, and $x$, if $P_{X}$ is faithful to $G$).
All descendants are always affected, due to the non-multiplicative noise term. These two differ in the row and column that correspond to $x$ since
$\text{Corr}(c\cdot x,y)$ $=\text{Corr}(x,y)$ (8)
$\text{Cov}(c\cdot x,y)$ $=c\cdot\text{Cov}(x,y).$ (9)
Therefore, to identify $x$, we should find the index in which $\Sigma$ differs but not $R$. Suppose that dimensions 3-6 of $h$ encode $R,R^{\prime},\Sigma,\Sigma^{\prime}$. Following the same strategy as the hard interventions, we can use the row attention to compute $R-R^{\prime},\Sigma-\Sigma^{\prime}$ and store them in dimensions 3, 4. Then we use the column attention to filter out variables that are independent from $x$ by storing the sum of each column in dimensions 5, 6. While not strictly impossible, it is unlikely that a variable dependent on $x$ would result in a column that sums to exactly 0. Thus, all columns with non-zero sums are either ancestors, descendants, or $x$. The feedforward network implements
$\text{FFN}(h_{\cdot,3-6})=\begin{cases}1&h_{\cdot,3}=0,h_{\cdot,4}\neq 0,h_{\cdot,5}\neq 0\\ 0&\text{otherwise}.\end{cases}$ (10)
This results in 1s in the rows and columns where $\Delta R$ and $\Delta\Sigma$ differ. After collapsing over incoming edges and normalizing to probabilities, the maximum probabilities can be found at the intervention targets.
Supporting both intervention types
Recall that the final output layer is a linear projection from $2d$ to 1. If this layer implements a simple summation over all $2d$, the predicted intervention targets are consistent with both hard and soft interventions. For soft interventions, $E=E^{\prime}$, so the hard intervention dimensions will be 0. Likewise, for hard interventions, both $R$ and $\Sigma$ will differ as the same locations, as the underlying variable has changed, so the soft intervention dimensions will be 0. Since the two techniques produce mutually exclusive predictions, this means that both hard and soft interventions can co-exist and be detected on different nodes.
Finally, we re-iterate that Cdn operates over high-dimensional features, rather than the inputs and outputs themselves. Thus, these sketches only serve to provide context regarding the axial attention framework’s computational capacity, rather than as a blueprint for the computations it performs.
Appendix B Datasets
B.1 Biological datasets
We converted all single cell datasets to LogTP10K + 1 expression values (log-normalized, transcripts per 10,000 UMIs). Perturb-seq dataset variables represented genes that were mapped and filtered by the authors. Sci-Plex dataset variables represented genes that appeared in at least 5,000 cells (threshold chosen to achieve a similar number of genes). We performed differential expression analysis via the scanpy package (Wolf et al., 2018), using the Wilcoxon signed-rank test (Wilcoxon, 1945) with Benjamini-Hochberg p-value correction and a threshold of adjusted p-value $<0.05$. Table 5 reports statistics of the raw, unprocessed datasets.
For Perturb-seq datasets, we kept perturbations with $>10$ differentially-expressed genes (DEGs), and clustered them using k-means with $k=200$, chosen heuristically based on log-fold change heatmaps (Figure 5). These clusters were used to inform data splits for seen cell lines, where the largest cluster was allocated to the training set, and all remaining clusters were split equally among train and test. The largest cluster(s) appear to contain perturbations with smaller effects. For Sci-Plex datasets, only 6 drug perturbations across 2 cell lines resulted in differential expression of their known protein targets (Supplementary Table 8 from McFaline-Figueroa et al. (2024)). Therefore, we used these exclusively as test sets. Table 6 reports statistics of the final, processed datasets. Figures 6 and 7 plot the full distribution of number of cells and DEGs per perturbation.
The full names of the Sci-Plex drugs are as follows.
  • Infigratinib (“infig”)
  • Nintedanib (“nint”)
  • Palbociclib (“palb”)
  • Doxorubicin (“doxo”)
  • Volasertib (“vola”)
Table 5: Extended biological dataset statistics (raw).
Type Source Accession Cell line ##\## Perts ##\## Genes ##\## NTCs ##\## Cells
Genetic Replogle et al. (2022) Figshare 20029387 K562 gw 9,866 8,248 75,328 1,989,578
K562 es 2,057 8,563 10,691 310,385
RPE1 2,393 8,749 11,485 247,914
Nadig et al. (2024) GSE220095 HepG2 2,393 9,624 4,976 145,473
Jurkat 2,393 8,882 12,013 262,956
Chemical McFaline-Figueroa et al. (2024) GSM7056151 A172 23 8,393 8,660 58,347
T98G 23 8,393 6,921 58,347
Table 6: Extended biological dataset statistics (processed).
Perturbations Genes Cells
Dataset Train Test Trivial Non-trivial Unique ##\## DE Median ##\## DE
K562 gw 1089 678 587 91 7,378 81 492,096
K562 es 640 420 348 72 8,492 226 213,552
RPE1 564 397 233 164 8,641 399 179,696
HepG2 364 263 162 101 9,282 271 79,309
Jurkat 679 333 262 71 8,432 162 174,698
A172 3 445 324 18,196
T98G 3 1,644 508 13,126
Figure 5: Heatmap of correlation between log-fold change, sorted by cluster.
Figure 6: Perturb-seq dataset statistics, after processing (Table 6).
Figure 7: Sci-Plex dataset statistics, after processing. There are 3 chemical perturbations per cell line.
Figure 8: Percentage of differentially expression genes (per perturbation) that were not in the GEARS knowledge graph, and whose effects could not be predicted.
B.2 Synthetic data
Synthetic data were generated using code modified from Dcdi (Brouillard et al., 2020) for soft interventions. We used the following implementations for causal mechanisms $f(x)$, where $x$ is the variable in question, $M$ is a binary mask for the parents of $x$, $X$ contains measurements of all variables, $E$ is independent Gaussian noise, and $W$ is a random weight matrix.
  • Linear: $f(x)=MXW+E$.
  • Polynomial: $f(x)=W_{0}+MXW_{1}+MX^{2}W_{2}+E$
Root causal mechanisms are uniform. For hard interventions, we set $f(x)\leftarrow z$, where $z\sim\text{Uniform}(-1,1)$. For soft interventions, we set
$f(x)$ $\leftarrow z_{1}^{\text{Sign}(z)}\cdot f(x)$
$z_{1}$ $\sim\text{Uniform}(2,4)$
$z$ $\sim\text{Uniform}(-1,1).$
That is, we multiply $f(x)$ by a scaling factor that is equal probability $\lessgtr 1$ (constant across all observations).
Appendix C Implementation details
C.1 Baselines
We used the latest releases of all baselines. For GenePt, this corresponds to the v2 March 2024 update, which used newer models and additional protein data, compared to their initial paper (GenePT_gene_protein_embedding_model_3_text).
Due to the varying coverage of genes in their domain knowledge, several baselines failed to make predictions for certain genes. Of the 2,842 unique genetic perturbations, only 2,819 ($99.1\%$) mapped to GenePt embeddings. The remainder used the mean gene embedding (within the dataset) as the language-based embedding.
Figure 8 depicts the low, but non-trivial proportion of differentially expressed genes for which Gears was unable to make a prediction, due to lack of node coverage in their processed gene ontology graph (Ashburner et al., 2000). These genes were not considered in the rankings or evaluations for Gears.
PdGrapher was published on the union of three distinct knowledge graphs, but their harmonized graphs were not available publicly, and the authors did not respond to requests for data sharing. As a result, we relied on solely the human reference interactome (Luck et al., 2020) graph, as it was the only one of the three that could be easily processed. All test perturbation targets could be inferred through PdGrapher.
Appendix D Additional analyses
Table 7: Ablation studies. We re-train Cdn without pretrained aggregator weights, marginal estimates, or global statistics.
K562 gw K562 es RPE1 HepG2 Jurkat
Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
Cdn 0.72 0.37 0.32 0.56 0.15 0.35 0.59 0.12 0.53 0.55 0.07 0.40 0.62 0.10 0.39
--pretrain $0.50$ $0.14$ $0.20$ $0.51$ $0.01$ $0.23$ $0.49$ $0.03$ $0.50$ $0.52$ $0.03$ $0.37$ $0.54$ $0.07$ $0.30$
--marginal $0.55$ $0.15$ $0.19$ $0.49$ $0.06$ $0.25$ $0.50$ $0.05$ $0.47$ $0.49$ $0.05$ $0.39$ $0.58$ $0.04$ $0.31$
--global $0.52$ $0.14$ $0.20$ $0.49$ $0.04$ $0.27$ $0.47$ $0.04$ $0.49$ $0.49$ $0.04$ $0.35$ $0.46$ $0.06$ $0.25$
Table 8: Ablation study on global statistics over synthetic datasets. Mlp and Axial use inverse covariance (same as pretraining). Corr was finetuned for correlation, and Corr+Cov was finetuned for correlation and covariance (concatenated).
$N$ $E$ Cdn Variant Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft)
[…]
Axial $\textbf{0.73}\scriptstyle\pm.05$ $\textbf{0.85}\scriptstyle\pm.03$ $\textbf{0.48}\scriptstyle\pm.06$ $\textbf{0.64}\scriptstyle\pm.05$ $\textbf{0.62}\scriptstyle\pm.09$ 0.77±.07plus-or-minus0.77.070.$7$07 $\textbf{0.73}\scriptstyle\pm.07$ $\textbf{0.82}\scriptstyle\pm.06$
+Corr $0.71\scriptstyle\pm.07$ $0.84\scriptstyle\pm.04$ $0.32\scriptstyle\pm.05$ $0.47\scriptstyle\pm.07$ $0.58\scriptstyle\pm.08$ $0.74\scriptstyle\pm.05$ $0.27\scriptstyle\pm.04$ $0.34\scriptstyle\pm.04$
+Corr+Cov $0.68\scriptstyle\pm.04$ $0.82\scriptstyle\pm.02$ $0.39\scriptstyle\pm.03$ $0.50\scriptstyle\pm.03$ $0.58\scriptstyle\pm.07$ $0.71\scriptstyle\pm.07$ $0.48\scriptstyle\pm.05$ $0.65\scriptstyle\pm.05$
[…]
+Corr $\textbf{0.78}\scriptstyle\pm.03$ 0.88±.02plus-or-minus0.88.020.$8$02 $0.35\scriptstyle\pm.02$ $0.52\scriptstyle\pm.02$ $0.73\scriptstyle\pm.06$ $0.84\scriptstyle\pm.03$ $0.34\scriptstyle\pm.05$ $0.41\scriptstyle\pm.06$
+Corr+Cov $0.75\scriptstyle\pm.03$ $0.84\scriptstyle\pm.02$ $0.34\scriptstyle\pm.03$ $0.47\scriptstyle\pm.06$ $0.69\scriptstyle\pm.07$ $0.79\scriptstyle\pm.05$ $0.53\scriptstyle\pm.04$ $\textbf{0.68}\scriptstyle\pm.04$
20 20 Mlp $\textbf{0.67}\scriptstyle\pm.05$ $\textbf{0.87}\scriptstyle\pm.02$ $0.28\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $60$06 $\textbf{0.84}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.04$ $0.51\scriptstyle\pm.04$
Axial $0.65\scriptstyle\pm.03$ $0.86\scriptstyle\pm.01$ $\textbf{0.49}\scriptstyle\pm.05$ $\textbf{0.72}\scriptstyle\pm.03$ $0.58\scriptstyle\pm.05$ $0.83\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.81}\scriptstyle\pm.02$
+Corr $0.62\scriptstyle\pm.03$ $0.83\scriptstyle\pm.01$ $0.25\scriptstyle\pm.05$ $0.47\scriptstyle\pm.05$ $0.58\scriptstyle\pm.05$ $0.81\scriptstyle\pm.02$ $0.21\scriptstyle\pm.02$ $0.37\scriptstyle\pm.03$
+Corr+Cov $0.60\scriptstyle\pm.03$ $0.81\scriptstyle\pm.01$ $0.31\scriptstyle\pm.03$ $0.58\scriptstyle\pm.01$ $0.58\scriptstyle\pm.04$ $0.80\scriptstyle\pm.02$ $0.40\scriptstyle\pm.03$ $0.65\scriptstyle\pm.03$
[…]
Axial $0.74\scriptstyle\pm.05$ $0.90\scriptstyle\pm.02$ $\textbf{0.38}\scriptstyle\pm.03$ $\textbf{0.65}\scriptstyle\pm.01$ $0.68\scriptstyle\pm.08$ $0.87\scriptstyle\pm.04$ $\textbf{0.64}\scriptstyle\pm.03$ $\textbf{0.79}\scriptstyle\pm.03$
[…]
+Corr+Cov $0.73\scriptstyle\pm.05$ $0.87\scriptstyle\pm.02$ $0.36\scriptstyle\pm.03$ $0.58\scriptstyle\pm.03$ $0.62\scriptstyle\pm.07$ $0.82\scriptstyle\pm.03$ $0.43\scriptstyle\pm.04$ $0.67\scriptstyle\pm.02$
Ablation studies
We run ablation studies on the model architecture (Table 7) by re-training Cdn without pretrained aggregator weights, global statistics, or marginal estimates. We find that pretrained weights and global correlation are essential to model performance. The marginal estimates are of mixed importance: they are beneficial on the genome-wide (gw) dataset and Jurkat, but less on the others.
Table 8 ablates variants of Cdn trained with different global statistics (but initialized with the inverse covariance pretrained weights) on synthetic data. The default summary statistic (Mlp, Axial) is inverse covariance. Corr denotes the correlation version, later finetuned on Perturb-seq data (since correlation is much easier to compute on large graphs, compared to inverse covariance). The Corr+Cov version aligns exactly with our theoretical intuition, and performs well in most settings (polynomial soft is the exception). This model modified the global layer to take an input of size $N\times N\times 2$, where each embedding was initialized to a distinct copy of the original. However, the inverse covariance version still performed the best, perhaps since there is no mismatch between the pretraining and finetuning statistics.
Runtime
We also compared the runtimes of various algorithms on the Perturb-seq and synthetic datasets. On the Perturb-seq data, all models finished running within minutes with the exception of Gears (Figure 9). Runtimes of the various causal algorithms varied significantly (Table 9). The slowest method was Dcdi, with an average runtime of around 10 hours on $N=20$ datasets, while the fastest were Ut-Igsp and the Mlp variant of Cdn. All models were benchmarked on equivalent hardware (A6000 GPU, 1 CPU core).
Perturb-seq uncertainty quantification
Due to space limitations, we report uncertainty estimates in Table 10. Multiple baselines produce deterministic results (Linear, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells. Specifically, for each perturbation with $M$ cells, we sample
$M^{\prime}=\min(M,\max(50,0.8M))$ (11)
cells uniformly at random, repeated 5 times.
Additional synthetic results
Table 11 reports results on all synthetic test datasets, averaging over all interventions for each dataset.
Figure 9: Inference runtimes on Perturb-seq datasets. K562 datasets are reported together. All models run on a single A6000 GPU, no constraint on memory (up to 500G). Only GEARS required over $\sim$20G of memory. To the best of our ability, we normalized batch size to 1.
Table 9: Runtimes on synthetic datasets (sec).
Nodes Model Min Max Mean Std
10 Ut-Igsp 1 98 26 31
Dcdi-G 31 27908 18731 9857
Dcdi-Dsf 7118 44440 23798 5688
Dci 41 1438 404 365
Bacadi 4005 9483 6284 1584
Bacadi-l 1076 1403 1272 76
Cdn (Mlp) 1 5 1 1
Cdn 17 147 39 27
20 Ut-Igsp 1 78 19 26
Dcdi-G 45 46281 35032 16068
Dcdi-Dsf 23181 55406 30076 7200
Dci 325 21414 4415 4707
Bacadi 15082 50737 26020 9044
Bacadi-l 2862 3819 3461 252
Cdn (Mlp) 1 3 1 0
Cdn 36 86 54 13
Table 10: Uncertainty quantification on Perturb-seq datasets by sub-sampling to $80\%$ of cells per perturbation (or 50, whichever is higher). Standard deviation reported over 5 rounds of sub-sampling. Cdn is also the most robust to sub-sampling, compared to baselines. GenePt performance is highly variable.
K562 gw K562 es RPE1
Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
Linear $0.50\scriptstyle\pm.00$ $0.17\scriptstyle\pm.01$ $0.20\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $10$01 $0.28\scriptstyle\pm.01$ $0.48\scriptstyle\pm.00$ $0.07\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$
Mlp $0.48\scriptstyle\pm.00$ $0.21\scriptstyle\pm.00$ $0.20\scriptstyle\pm.00$ $0.48\scriptstyle\pm.00$ 0.11±.00plus-or-minus0.11.000.$1$00 $0.19\scriptstyle\pm.00$ $0.53\scriptstyle\pm.00$ $0.15\scriptstyle\pm.00$ $0.39\scriptstyle\pm.00$
GenePt $0.54\scriptstyle\pm.09$ $0.28\scriptstyle\pm.06$ $0.29\scriptstyle\pm.08$ $0.50\scriptstyle\pm.08$ $0.21\scriptstyle\pm.09$ $0.39\scriptstyle\pm.10$ $0.51\scriptstyle\pm.10$ $0.20\scriptstyle\pm.08$ $0.51\scriptstyle\pm.07$
Gears $0.50\scriptstyle\pm.01$ $0.13\scriptstyle\pm.01$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.51\scriptstyle\pm.00$ $0.07\scriptstyle\pm.00$ $0.23\scriptstyle\pm.00$ $0.46\scriptstyle\pm.00$ $0.07\scriptstyle\pm.01$ $0.41\scriptstyle\pm.02$
PdG $0.49\scriptstyle\pm.01$ $0.21\scriptstyle\pm.01$ $0.25\scriptstyle\pm.00$ $0.50\scriptstyle\pm.01$ $0.16\scriptstyle\pm.01$ $0.32\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.08\scriptstyle\pm.01$ $0.47\scriptstyle\pm.01$
[…]
HepG2 Jurkat
rank $R_{20}$ $r$ rank $R_{20}$ $r$
Linear $0.49\scriptstyle\pm.00$ $0.06\scriptstyle\pm.01$ $0.42\scriptstyle\pm.01$ $0.50\scriptstyle\pm.00$ $0.08\scriptstyle\pm.00$ $0.27\scriptstyle\pm.01$
Mlp $0.50\scriptstyle\pm.00$ $0.09\scriptstyle\pm.00$ $0.41\scriptstyle\pm.00$ $0.47\scriptstyle\pm.00$ $0.12\scriptstyle\pm.00$ $0.31\scriptstyle\pm.00$
GenePt $0.50\scriptstyle\pm.13$ $\textbf{0.19}\scriptstyle\pm.08$ $0.45\scriptstyle\pm.06$ $0.51\scriptstyle\pm.07$ 0.22±.09plus-or-minus0.22.090.$2$09 $0.40\scriptstyle\pm.11$
Gears $0.51\scriptstyle\pm.00$ 0.11±.00plus-or-minus0.11.000.$1$00 $0.45\scriptstyle\pm.00$ $0.52\scriptstyle\pm.01$ $0.08\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$
PdG $0.49\scriptstyle\pm.00$ $0.05\scriptstyle\pm.01$ $0.41\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ 0.11±.01plus-or-minus0.11.010.$1$01 $0.34\scriptstyle\pm.00$
Cdn $\textbf{0.65}\scriptstyle\pm.01$ $0.17\scriptstyle\pm.01$ $\textbf{0.46}\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.01$ $\textbf{0.24}\scriptstyle\pm.02$ $\textbf{0.43}\scriptstyle\pm.01$
Table 11: Intervention target prediction results on synthetic datasets, extended results. Uncertainty is standard deviation over 5 distinct datasets. Metrics are averaged over all $3N$ perturbations for a given dataset (1-3 targets).
$N$ $E$ Model Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft)
[…]
10 10 Ut-Igsp $0.29\scriptstyle\pm.02$ $0.57\scriptstyle\pm.03$ $0.30\scriptstyle\pm.01$ $0.63\scriptstyle\pm.02$ 0.33±.04plus-or-minus0.33.040.$3$04 $0.59\scriptstyle\pm.03$ $0.32\scriptstyle\pm.01$ $0.64\scriptstyle\pm.02$
Dcdi-G $0.39\scriptstyle\pm.04$ $0.52\scriptstyle\pm.04$ $40$04 $0.51\scriptstyle\pm.04$ $0.38\scriptstyle\pm.03$ $0.49\scriptstyle\pm.02$ $0.36\scriptstyle\pm.03$ $0.49\scriptstyle\pm.03$
Dcdi-Dsf $0.36\scriptstyle\pm.02$ $0.51\scriptstyle\pm.04$ $0.40\scriptstyle\pm.05$ $0.53\scriptstyle\pm.05$ $0.37\scriptstyle\pm.02$ $0.50\scriptstyle\pm.01$ $0.38\scriptstyle\pm.02$ $0.49\scriptstyle\pm.03$
Dci 0.55±.08plus-or-minus0.55.080.$5$08 $0.79\scriptstyle\pm.04$ $\textbf{0.50}\scriptstyle\pm.08$ $\textbf{0.70}\scriptstyle\pm.05$ $0.47\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ $0.41\scriptstyle\pm.02$ $0.67\scriptstyle\pm.01$
BaCaDI-E $0.26\scriptstyle\pm.03$ $0.61\scriptstyle\pm.05$ $0.42\scriptstyle\pm.13$ $0.63\scriptstyle\pm.11$ $0.25\scriptstyle\pm.03$ $0.61\scriptstyle\pm.05$ $0.30\scriptstyle\pm.05$ $0.61\scriptstyle\pm.04$
BaCaDI-M $0.23\scriptstyle\pm.01$ $0.56\scriptstyle\pm.03$ $0.37\scriptstyle\pm.11$ $0.62\scriptstyle\pm.11$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.56\scriptstyle\pm.04$ $0.30\scriptstyle\pm.05$ $0.61\scriptstyle\pm.04$
[…]
Cdn (Axial) $\textbf{0.73}\scriptstyle\pm.05$ $\textbf{0.85}\scriptstyle\pm.03$ $0.48\scriptstyle\pm.06$ $0.64\scriptstyle\pm.05$ $\textbf{0.62}\scriptstyle\pm.09$ 0.77±.07plus-or-minus0.77.070.$7$07 $\textbf{0.73}\scriptstyle\pm.07$ $\textbf{0.82}\scriptstyle\pm.06$
10 20 Ut-Igsp $0.28\scriptstyle\pm.02$ $0.56\scriptstyle\pm.03$ $0.26\scriptstyle\pm.01$ $0.59\scriptstyle\pm.03$ $0.28\scriptstyle\pm.03$ $0.58\scriptstyle\pm.04$ $0.27\scriptstyle\pm.01$ $0.60\scriptstyle\pm.01$
Dcdi-G $0.42\scriptstyle\pm.02$ $0.54\scriptstyle\pm.02$ $0.40\scriptstyle\pm.03$ $0.51\scriptstyle\pm.04$ $0.40\scriptstyle\pm.06$ $0.53\scriptstyle\pm.05$ $0.39\scriptstyle\pm.02$ $0.51\scriptstyle\pm.03$
Dcdi-Dsf $0.41\scriptstyle\pm.03$ $0.52\scriptstyle\pm.03$ $0.39\scriptstyle\pm.05$ $0.52\scriptstyle\pm.05$ $0.39\scriptstyle\pm.02$ $0.51\scriptstyle\pm.02$ $0.39\scriptstyle\pm.03$ $50$05
[…]
BaCaDI-E $0.34\scriptstyle\pm.04$ $0.68\scriptstyle\pm.03$ $0.53\scriptstyle\pm.08$ $0.71\scriptstyle\pm.04$ 0.33±.03plus-or-minus0.33.030.$3$03 $0.72\scriptstyle\pm.04$ $0.64\scriptstyle\pm.07$ $0.78\scriptstyle\pm.05$
BaCaDI-M $0.27\scriptstyle\pm.02$ $0.63\scriptstyle\pm.03$ $0.48\scriptstyle\pm.09$ $0.71\scriptstyle\pm.04$ $0.27\scriptstyle\pm.04$ $0.64\scriptstyle\pm.08$ $0.59\scriptstyle\pm.09$ 0.77±.06plus-or-minus0.77.060.$7$06
[…]
20 20 Ut-Igsp $0.15\scriptstyle\pm.01$ $0.54\scriptstyle\pm.01$ $0.18\scriptstyle\pm.01$ $0.65\scriptstyle\pm.02$ $20$02 $0.60\scriptstyle\pm.02$ $0.21\scriptstyle\pm.02$ $0.67\scriptstyle\pm.02$
Dcdi-G $0.24\scriptstyle\pm.02$ $0.49\scriptstyle\pm.02$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.02$ 0.22±.03plus-or-minus0.22.030.$2$03 $0.50\scriptstyle\pm.03$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.03$
Dcdi-Dsf $0.24\scriptstyle\pm.02$ $0.50\scriptstyle\pm.02$ $0.23\scriptstyle\pm.02$ $0.50\scriptstyle\pm.03$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.50\scriptstyle\pm.02$ $0.20\scriptstyle\pm.01$ $0.48\scriptstyle\pm.02$
Dci $0.43\scriptstyle\pm.03$ 0.77±.01plus-or-minus0.77.010.$7$01 $0.45\scriptstyle\pm.08$ $\textbf{0.75}\scriptstyle\pm.04$ $0.48\scriptstyle\pm.05$ $0.76\scriptstyle\pm.03$ $0.50\scriptstyle\pm.02$ $0.79\scriptstyle\pm.02$
BaCaDI-E $0.14\scriptstyle\pm.01$ $0.61\scriptstyle\pm.03$ $0.25\scriptstyle\pm.10$ $60$06 0.22±.06plus-or-minus0.22.060.$2$06 $0.76\scriptstyle\pm.06$ $0.30\scriptstyle\pm.06$ $0.74\scriptstyle\pm.05$
BaCaDI-M $0.12\scriptstyle\pm.00$ $0.58\scriptstyle\pm.02$ $0.17\scriptstyle\pm.03$ $0.60\scriptstyle\pm.05$ $0.14\scriptstyle\pm.02$ 0.66±.05plus-or-minus0.66.050.$6$05 0.22±.05plus-or-minus0.22.050.$2$05 $0.71\scriptstyle\pm.06$
Cdn (Mlp) $\textbf{0.67}\scriptstyle\pm.05$ $\textbf{0.87}\scriptstyle\pm.02$ $0.28\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $60$06 $\textbf{0.84}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.04$ $0.51\scriptstyle\pm.04$
Cdn (Axial) $0.65\scriptstyle\pm.03$ $0.86\scriptstyle\pm.01$ $\textbf{0.49}\scriptstyle\pm.05$ $0.72\scriptstyle\pm.03$ $0.58\scriptstyle\pm.05$ $0.83\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.81}\scriptstyle\pm.02$
20 40 Ut-Igsp $0.14\scriptstyle\pm.01$ $0.57\scriptstyle\pm.01$ $0.14\scriptstyle\pm.01$ $0.61\scriptstyle\pm.02$ $0.20\scriptstyle\pm.01$ $0.62\scriptstyle\pm.02$ $0.19\scriptstyle\pm.01$ $0.65\scriptstyle\pm.01$
Dcdi-G 0.22±.03plus-or-minus0.22.030.$2$03 $0.49\scriptstyle\pm.03$ $0.21\scriptstyle\pm.02$ $0.48\scriptstyle\pm.03$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.02$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.50\scriptstyle\pm.02$
Dcdi-Dsf $0.21\scriptstyle\pm.02$ $0.46\scriptstyle\pm.03$ $0.21\scriptstyle\pm.02$ $0.47\scriptstyle\pm.02$ $0.23\scriptstyle\pm.02$ $0.50\scriptstyle\pm.03$ $0.21\scriptstyle\pm.01$ $0.49\scriptstyle\pm.03$
[…]
BaCaDI-E $0.23\scriptstyle\pm.05$ $0.70\scriptstyle\pm.04$ 0.33±.10plus-or-minus0.33.100.$3$10 $0.67\scriptstyle\pm.04$ $0.29\scriptstyle\pm.08$ $0.83\scriptstyle\pm.04$ $0.45\scriptstyle\pm.18$ $0.78\scriptstyle\pm.06$
BaCaDI-M $0.15\scriptstyle\pm.01$ $0.64\scriptstyle\pm.02$ $0.19\scriptstyle\pm.04$ $0.63\scriptstyle\pm.02$ $0.17\scriptstyle\pm.02$ $0.70\scriptstyle\pm.05$ $0.31\scriptstyle\pm.10$ 0.77±.07plus-or-minus0.77.070.$7$07
[…]
Cdn (Axial) $0.74\scriptstyle\pm.05$ $0.90\scriptstyle\pm.02$ $0.38\scriptstyle\pm.03$ $0.65\scriptstyle\pm.01$ $0.68\scriptstyle\pm.08$ $0.87\scriptstyle\pm.04$ $\textbf{0.64}\scriptstyle\pm.03$ $0.79\scriptstyle\pm.03$

### v2

@@
-Predicting perturbation targets with causal differential networks
+Identifying biological perturbation targets through causal differential networks
 Abstract
-Rationally identifying variables responsible for changes to a biological system can enable myriad applications in disease understanding and cell engineering.
-From a causality perspective, we are given two datasets generated by the same causal model, one observational (control) and one interventional (perturbed).
-The goal is to isolate the subset of measured variables (e.g.
-genes) that were the targets of the intervention, i.e.
-those whose conditional independencies have changed.
-Knowing the causal graph would limit the search space, allowing us to efficiently pinpoint these variables.
-However, current algorithms that infer causal graphs in the presence of unknown intervention targets scale poorly to the hundreds or thousands of variables in biological data, as they must jointly search the combinatorial spaces of graphs and consistent intervention targets.
-In this work, we propose a causality-inspired approach for predicting perturbation targets that decouples the two search steps.
-First, we use an amortized causal discovery model to separately infer causal graphs from the observational and interventional datasets.
-Then, we learn to map these paired graphs to the sets of variables that were intervened upon, in a supervised learning framework.
-This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets, each with thousands of measured variables.
-We also demonstrate significant improvements over six causal discovery algorithms in predicting intervention targets across a variety of tractable, synthetic datasets.
+Identifying variables responsible for changes to a biological system enables applications in drug target discovery and cell engineering.
+Given a pair of observational and interventional datasets, the goal is to isolate the subset of observed variables that were the targets of the intervention.
+Directly applying causal discovery algorithms is challenging: the data may contain thousands of variables with as few as tens of samples per intervention, and biological systems do not adhere to classical causality assumptions.
+We propose a causality-inspired approach to address this practical setting.
+First, we infer noisy causal graphs from the observational and interventional data.
+Then, we learn to map the differences between these graphs, along with additional statistical features, to sets of variables that were intervened upon.
+Both modules are jointly trained in a supervised framework, on simulated and real data that reflect the nature of biological interventions.
+This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets.
+We also demonstrate significant improvements over current causal discovery methods for predicting soft and hard intervention targets across a variety of synthetic data.
+perturbation, transcriptomics, causality
 1 Introduction
@@
 drugs, gene knockdown) can alter a cell’s state.
-While perturbations can affect numerous downstream variables, identifying the root causes, or targets, that drive these transitions has vast therapeutic implications, from cellular reprogramming (Cherry & Daley, 2012) to mechanism of action elucidation (Schenone et al., 2013).
-Large-scale experiments (Replogle et al., 2022) have attempted to map the effects of perturbing individual genes on single cells, and in-silico approaches (Roohani et al., 2023; Lotfollahi et al., 2023) have been designed for the “forward” inference task of predicting the post-perturbation expression of each gene.
-In principle, these models can be used within an active learning framework to suggest perturbation targets (Huang et al., 2023; Zhang et al., 2023a).
-However, the number of inference calls required scales exponentially with the size of the candidate set (e.g.
-drugs with off-target effects), rendering these approaches impractical for larger search spaces.
-These models’ predictive performance as oracles has also been called into question by subsequent works (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024), highlighting the difficulty of this forward approach.
-Finally, there are limited training data for combinatorial perturbations: the most widely used dataset only contains around a hundred pairs (Norman et al., 2019).
-The “reverse” strategy – of directly predicting targets from perturbation data – alleviates these problems to an extent.
-Towards this end, it is common to assume (sometimes implicitly) that the data were generated by a structured causal model, and the perturbation targets can be directly inferred from the changes to this model.
-While the effects of perturbations are highly contextual, specific to the cell lines in which they are applied (Nadig et al., 2024), existing methods rely on non-specific, data-mined knowledge graphs as the backbones for their causal structures (Cosgrove et al., 2008; Gonzalez et al., 2024).
-It is hard to quantify the extent to which these graphs are relevant to each setting, as the quality and focus of assays have changed over time, and interactions present in one cell type may not exist in another (Huttlin et al., 2021).
-Our goal is to predict the perturbation targets that translate one set of cells into another, while inferring the underlying causal structures directly from the data.
-We formalize this task within a causality framework.
-Specifically, given an observational dataset (e.g.
-gene expression of control cells) and an interventional dataset (e.g.
-perturbed cells) generated by the same system (e.g.
-cell type), we would like to identify the variables (e.g.
-genes) whose data-generating mechanisms are different between the two settings.
-Causal discovery algorithms have been designed to solve this exact problem, on interventional data with unknown targets (Squires et al., 2020; Brouillard et al., 2020; Hägele et al., 2023).
-These algorithms operate jointly over the spaces of graphs and intervention target sets.
-Due to this large search space, these algorithms are unable to scale to the thousands of variables and hundreds of thousands of observations in modern transcriptomics data, while simultaneously being robust to as few as tens of observations per interventional setting (Nadig et al., 2024).
-To address unknown intervention target prediction in context of biological perturbations, we propose causal differential networks (Cdn).
-Drawing upon recent advances in scalable and data-efficient causal discovery (Wu et al., 2024), Cdn first “featurizes” the observational and interventional datasets in terms of their predicted causal graphs, using a pretrained causal discovery module.
-These paired, edge-level representations are then used as input to an axial attention-based classifier (Ho et al., 2020), which is supervised by ground truth intervention targets (on either real or synthetic data).
-We thoroughly evaluate Cdn on both real transcriptomic data and in synthetic settings.
-Cdn outperforms the state-of-the-art in perturbation modeling on the five largest Perturb-seq datasets at the time of publication (CRISPRi,  Replogle et al.
-(2022); Nadig et al.
-(2024)), and two Sci-Plex datasets (chemical perturbations, McFaline-Figueroa et al.
-(2024)), without using any external knowledge.
-In fact, Cdn is the only model that consistently ranks the ground truth perturbation targets higher than would be expected by random.
-Moreoever, Cdn is able to maintain decent performance even on unseen cell lines, demonstrating its potential to generalize.
-Finally, on synthetic settings, Cdn outperforms causal discovery algorithms for unknown intervention targets.
-To conclude, our contributions are three-fold.
-  1. 1.
-    We propose causal differential networks (Cdn), a causal discovery-based approach for perturbation target prediction on transcriptomic data, which achieves the state-of-the-art on genetic and chemical perturbation datasets.
-  2. 2.
-    We demonstrate that Cdn outperforms current causal discovery approaches for unknown intervention target prediction for both hard and soft interventions in diverse synthetic data.
-  3. 3.
-    Finally, we have prepared seven high-quality transcriptomics datasets, to serve as a benchmark for future machine learning studies.
+While perturbations can affect numerous downstream variables, identifying the root causes, or targets, that induce these transitions has vast therapeutic implications, from cellular reprogramming (Cherry & Daley, 2012) to mechanism of action elucidation (Schenone et al., 2013).
+Machine learning methods have primarily been designed to infer the effects of perturbations on cells, with the goal of generalizing to unseen perturbations (Roohani et al., 2023), or unseen cell distributions (Lotfollahi et al., 2023; Bunne et al., 2023).
+In principle, these models can be used within an active learning framework to discover intervention targets (Huang et al., 2023; Zhang et al., 2023a).
+However, the number of candidate sets scales exponentially with the number of variables under consideration, and there are limited training data for combinatorial perturbations ($<150$ pairs in Norman et al.
+(2019)).
+Thus, this option may be impractical for larger search spaces.
+Alternatively, some methods predict targets by comparing pairs of “before” and “after” cell states, similar to this paper.
+They do so by attributing observed differences to sparse, mechanistic changes.
+Specifically, relationships between variables are represented through a graph, e.g.
+a gene regulatory network (Babu et al., 2004).
+Existing methods often use data-mined relations as priors for this graph, and search for modifications to the set of edges that best explain the interventional data (Cosgrove et al., 2008; Gonzalez et al., 2024).
+However, biological knowledge graphs are incomplete and noisy priors.
+They contain heterogeneous, potentially conflicting information, since the reported findings were observed under diverse contexts (Ashburner et al., 2000).
+Moreover, different cell populations may exhibit distinct gene regulatory behavior (Huttlin et al., 2021), leading to incomplete knowledge, especially of less well-characterized systems.
+In this paper, we propose causal differential networks (Cdn): a causality-inspired approach to identify variables that drive desired shifts in cell state, while estimating their mechanistic structure directly from data.
+Given a pair of observational and interventional datasets, we train a causal structure learner to predict causal graphs that could have generated each dataset.
+The pair of graphs is input to an attention-based classifier, which predicts whether each variable was subject to intervention.
+To address the challenges of data noise and sparsity, and to simulate the structure of biological interventions, these models are trained jointly in a supervised (amortized) framework (Ke et al., 2022; Petersen et al., 2023), over thousands of synthetic or real datasets.
+We evaluate Cdn on real transcriptomic data and synthetic settings.
+Cdn outperforms the state-of-the-art in perturbation modeling (deep learning and statistical approaches) on the five largest Perturb-seq datasets at the time of publication  (Replogle et al., 2022; Nadig et al., 2024), without using any external knowledge.
+Furthermore, Cdn generalizes with minimal performance drop to unseen cell lines, which have different supports (genes), causal mechanisms (gene regulatory networks), and data distributions.
+Finally, on synthetic settings, Cdn outperforms causal discovery approaches for estimating unknown intervention targets.
 Figure 1: (A) Biological applications.
-(B) Differences between observation dataset $D_{\text{obs}}$ and interventional dataset $D_{\text{int}}$ can be attributed to changes in the underlying causal model $G$.
-(C) We use an amortized causal discovery module to obtain features that represent $G_{\text{obs}}$ and $G_{\text{int}}$.
-Then we train a supervised classifier to predict the unknown intervention targets.
+(B) Differences between observation dataset $D_{\text{obs}}$ and interventional dataset $D_{\text{int}}$ can be attributed to changes in the underlying causal mechanisms.
+(C) We use an amortized causal discovery model to predict $G_{\text{obs}}$ and $G_{\text{int}}$, whose hidden representations are input to a perturbation target classifier.
+Both modules are jointly trained.
 2 Background and related work
-2.1 Identifying perturbation targets
-The rise of large-scale perturbation screens (Replogle et al., 2022) has enabled machine learning approaches for predicting perturbation targets.
-One line of work focuses on active learning to reduce experimental costs (Zhang et al., 2023a; Huang et al., 2023), but these models require inference calls to an oracle model for each candidate perturbation.
-Given that human cells express over ten thousand genes, and drugs may impact multiple targets, the search space is enormous.
-Moreover, subsequent works have found that naive baselines (mean of the training set, linear regression) outperform these approaches at their intended task (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024).
-In contrast, perturbation targets can be predicted directly from data, using a variety of domain knowledge (Cosgrove et al., 2008; Gonzalez et al., 2024; Roohani et al., 2024).
-However, these external data were collected from highly inhomogenous sources, which may be inconsistent with the data at hand.
-2.2 Causal discovery with unknown intervention targets
-Perturbation target prediction can be formalized as a causal discovery problem, using data generated under interventions with unknown targets.
-Specifically, a causal graphical model is a directed graph $G=(V,E)$, where nodes $i\in V$ map to random variables $X_{i}\in X$, and edges $(i,j)\in E$ represent relationships from $X_{i}$ to $X_{j}$.
-There are a number of common assumptions that relate $G$ to the data distribution $P_{X}$, which we defer to works such as Spirtes et al.
-(2001); Yang et al.
-(2018); Zhang et al.
-(2023b), since the identifiability of any particular system is not the focus of this paper.
-Data generated directly from $P_{X}$ are known as observational data. A causal graphical model allows us to perform interventions by assigning new conditionals
+2.1 Problem statement
+We consider a system of $N$ random variables $X_{i}\in X$, whose relationships are represented by directed graph $G=(V,E)$, where nodes $i\in V$ map to variables $X_{i}$, and edges $(i,j)\in E$ to relationships from $X_{i}$ to $X_{j}$.
+Samples from the data distribution $P_{X}$ are known as observational.
+We can perform interventions by assigning new conditionals
 $P(X_{i}\mid X_{\pi_{i}})\leftarrow\tilde{P}(X_{i}\mid X_{\pi_{i}}),$ (1)
@@
 $P(X_{i}\mid X_{\pi_{i}})\neq\tilde{P}(X_{i}\mid X_{\pi_{i}}),\forall i\in I.$ (2)
-This task has been well-studied in the causality literature, often as a sub-problem or setting while inferring graph $G$ from $D_{\text{obs}}$ and (multiple) $D_{\text{int}}$.
-Ke et al.
-(2019) first proposed a discovery algorithm for unknown interventions, on discrete data of up to 50 variables.
-Jaber et al.
-(2020) provides identifiability guarantees for soft interventions with unknown targets and proposes a constraint-based algorithm, which requires an exponential number of conditional independence tests with respect to the number of variables.
-Continuous causal discovery algorithms that are consistent despite unknown interventions include Dcdi (Brouillard et al., 2020) and BaCaDI (Hägele et al., 2023), which treat intervention target identities as learnable parameters of a generative model over $P_{X}$.
-Dcdi uses the Gumbel-Softmax approximation (Jang et al., 2017) to learn these categorical variables, while Bacadi takes a Bayesian approach, with gradient-based methods for efficient posterior estimation (Liu & Wang, 2016).
-Finally, Varici et al.
+In the context of molecular biology, a perturbation is a hard or soft intervention, and $I$ is the set of perturbation targets, where $\|I\|\geq 1$.
+2.2 Modeling perturbations
+Perturbation experiments result in datasets of the form $D_{\text{obs}},\{(I^{k},D_{\text{int}}^{k})\}$, where $D_{\text{obs}}\sim P_{X}$, $D_{\text{int}}^{k}\sim\tilde{P}^{k}_{X}$.
+In this paper, we focus on the transcriptomic data modality, where variables $X_{i}$ represent the levels of gene $i$; the number of variables $N$ ranges from hundreds to thousands; $D_{\text{obs}}$ contains thousands of samples; and each $D_{\text{int}}$ contains tens to hundreds of samples (Replogle et al., 2022).
+Each experiment is conducted in a single cell line, which represents a distinct set of variables $X$, distribution $P_{X}$, and graph $G$.
+These three objects may share similarities across cell lines, as certain genes or pathways are essential for survival.
+However, the empirical discrepancy is quite large between experiments.
+For example, Nadig et al.
+(2024) reports that the median correlation in log-fold change of genes under the same perturbation, across two experiments in the same cell line, by the same lab, was 0.16.
+There are two directions for modeling perturbations.
+To reduce experimental costs, which scale with the number of cell lines and perturbations, one option is to infer the post-intervention distribution of cells.
+That is, given $D_{\text{obs}}\sim P_{X}$ and $I$, the goal is to infer $\tilde{P}_{X}$.
+Works such as scGen (Lotfollahi et al., 2019), CPA (Lotfollahi et al., 2023), ChemCPA (Hetzel et al., 2022), and CellOT (Bunne et al., 2023) aim to generalize to new settings $P_{X}$, e.g.
+unseen cell lines or patients.
+Alternatively, models like GEARS (Roohani et al., 2023), AttentionPert (Bai et al., 2024), and GIM (Schneider et al., 2024) are designed to generalize to unseen perturbations.
+These approaches do not explicitly predict targets, but if the space of perturbations $\mathcal{I}$ can be efficiently enumerated, these models are useful for comparing different $I$ (Zhang et al., 2023a; Huang et al., 2023).
+For example, given an oracle $\phi:D_{\text{obs}},I\to\tilde{P}$, the $I^{*}$ associated with $\tilde{P}^{*}$ is
+$I^{*}=\arg\min_{I\in\mathcal{I}}d(\tilde{P}^{*},\phi(D_{\text{obs}},I))$ (3)
+for some dissimilarity $d$.
+However, recent works have reported that the predictive ability of current models does not exceed that of naive baselines (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024), reflecting potential barriers to this approach.
+Perturbation datasets can also be used to train models for predicting targets of interest, as is the focus of this work.
+A common strategy is to relate observed changes (e.g.
+disease, gene levels) to mechanistic explanations (e.g.
+pathways), in which the targets are involved.
+To draw this connection, current approaches use external information such as knowledge graphs (Cosgrove et al., 2008; Gonzalez et al., 2024) and natural language (Roohani et al., 2024).
+However, these information are collected from highly inhomogenous sources (Ashburner et al., 2000), and it is difficult to quantify their correctness and completeness in each new setting.
+In contrast to these approaches, we choose to infer the mechanistic structure directly from data.
+2.3 Causal discovery
+The task of inferring causal mechanisms from data is known as causal discovery, or causal structure learning (Spirtes et al., 2001).
+In the simplest setting, the goal is to infer $G$ from $D_{\text{obs}}$.
+The earliest works were discrete search algorithms, which operated over the combinatorial space of potential graphs (Spirtes et al., 1995; Tian & Pearl, 2001; Chickering, 2002; Shimizu et al., 2006).
+More recently, a number of works have reframed the task as a constrained optimization over continuous adjacency matrices (Zheng et al., 2018; Lachapelle et al., 2020; Sanchez et al., 2023).
+If data from multiple environments are available, they can provide additional identifiability regarding the system (Jaber et al., 2020; Zhang et al., 2023b).
+Differences between these environments, e.g.
+intervention targets $I$, are either given as input or part of the inference task.
+As in the observational case, discovery algorithms can be categorized into discrete search (Hauser & Bühlmann, 2012; Ghassami et al., 2018; Ke et al., 2019; Huang et al., 2020; Mooij et al., 2020) or continuous optimization (Brouillard et al., 2020; Hägele et al., 2023).
+Additionally, Varici et al.
 (2022) and Yang et al.
-(2024) frame intervention target prediction as a separate task, similar to this paper.
-However, the former strictly assumes linearity, while the latter requires data from multiple environments, which are not available here.
-2.3 Amortized causal discovery algorithms
-The majority of causal discovery algorithms operate on one or more datasets $D$, which correspond to a single graph $G$ (Spirtes et al., 1995; Hauser & Bühlmann, 2012; Mooij et al., 2020; Brouillard et al., 2020).
-These algorithms must be fit or run from scratch for each data distribution – a potential challenge in low-resource scenarios, where there are too few observations of too many variables.
-Recently, amortized causal discovery algorithms have addressed causal structure learning as a supervised machine learning problem.
-Using large numbers of synthetic datasets, generated by (known) synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Ke et al., 2022; Lorch et al., 2022; Wu et al., 2024).
-Two immediate benefits are fast runtimes and robustness to aleatoric noise.
-In this work, we leverage the framework proposed by Wu et al.
-(2024), as it has been benchmarked on large, transcriptomics data (Chevalley et al., 2022).
+(2024) frame intervention target prediction as its own task, similar to this paper.
+However, the former strictly assumes linearity, while the latter requires at least as many environments as $\|I\|$ (we only consider before and after perturbation).
+Finally, Dci (Belyaeva et al., 2021) infers edge-level differences between linear Gaussian models, rather than identifying the full causal structure.
+Our work is also motivated by the idea that differences between summary statistics (e.g.
+regression coefficients in Dci) are informative of changes between systems.
+Thus, Cdn can be viewed as implementing similar operations within an end-to-end, supervised framework.
+2.4 Amortized causal inference
+While the majority of causal discovery algorithms must be fit or run from scratch for each data distribution, amortized causal discovery algorithms have addressed this task as a supervised machine learning problem.
+Using large numbers of synthetic datasets, generated by synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Li et al., 2020; Ke et al., 2022; Lorch et al., 2022; Petersen et al., 2023; Wu et al., 2024).
+Compared to traditional approaches, these algorithms trade explicit assumptions for implicit characteristics of the data simulation process (Montagna et al., 2024).
+These models are also fast and robust to noise.
+In this work, we simulate interventions that reflect biological perturbations, and our model architecture builds upon that of Wu et al.
+(2024), as it scales easily to hundreds of variables.
 3 Methods
-Our model is composed of two modules: a frozen causal featurizer, and a learned differential network, which we refer to jointly as Cdn (Figure 2).
-The causal featurizer is a pretrained, amortized causal discovery model (Wu et al., 2024) that runs efficiently on up to a thousand nodes (Section 3.1), while the differential network is an axial-attention based classifier (Ho et al., 2020) that predicts the nodes whose data-generating mechanisms have changed (Section 3.2).
-3.1 Causal featurizer
-The causal featurizer takes as input datasets $D_{\text{obs}}$ and $D_{\text{int}}$ containing samples of $N$ variables, generated from observational and interventional settings.
-For example, $D_{\text{obs}}$ may correspond the gene expression matrix of non-targeting control cells; $D_{\text{int}}$ is the gene expression of the same type of cells, subject to a perturbation with an unknown target; and $N$ is the number of genes detected.
-We use an amortized causal discovery model to obtain features $h_{\text{obs}},h_{\text{int}}\in\mathbb{R}^{N\times N\times d}$, which reflect the $N\times N$ adjacency matrices of causal graphs that may have generated each dataset.
-Specifically, to featurize each dataset, we follow the three steps proposed in Wu et al.
-(2024).
+We introduce causal differential networks (Cdn), a causality inspired method for predicting intervention targets.
+On a high level, Cdn is motivated by the idea that differences in observations $D_{\text{obs}}$ and $D_{\text{int}}$ can be explained by the changes to their underlying causal mechanisms (Figure 1B).
+Our model is composed of two modules: a causal structure learner (Section 3.1), and a differential network classifier (Section 3.2).
+Given a single dataset $D$, the causal structure learner predicts a putative $G$.
+The pair of datasets $D_{\text{obs}}$ and $D_{\text{int}}$ yields graphs $G_{\text{obs}}$ and $G_{\text{int}}$, whose features are input to the differential network, to predict targets $I$ (Figure 1C).
+Both modules are trained jointly, supervised by classification losses for the graph and predicted targets.
+Finally, individual samples are known to be noisy in single-cell biology and are often analyzed in “pseudo-bulk” (Blake et al., 2003).
+Thus, instead of operating over raw data, we also represent datasets in terms of their summary statistics.
+Figure 2: Our model is composed of a causal structure learner and a differential network.
+Learned components in yellow (causal structure learner) and green (differential network).
+Both modules are trained jointly to optimize $\mathcal{L}=\mathcal{L}_{G}+\mathcal{L}_{I}$.
+3.1 Causal structure learner
+The causal structure learner takes as input dataset $D$ and outputs the predicted $G$.
+In principle, this module can be implemented with any data-efficient causal discovery model.
+Here, we based the architecture on Sea (Wu et al., 2024), as its runtime complexity does not explicitly depend on the number of samples (which can be large for $D_{\text{obs}}$).
+Input featurization  We represent dataset $D$ through two types of statistics.
+First, we compute global statistics $\rho\in\mathbb{R}^{N\times N}$, like pairwise correlation, over all variables.
+Second, we obtain estimates of local causal structure by running classical causal discovery algorithms like FCI (Spirtes et al., 1995) over small subsets of variables.
+We denote the set of estimates as $E^{\prime}\in\mathcal{E}^{k\times k\times T}$, where $\mathcal{E}$ is the space of possible edge types (e.g.
+,$\leftrightarrow$,-$\rightarrow$ , - , →), and $k,T$ are hyperparameters that trade off accuracy vs.
+runtime.
+To compare $\rho$ and $E^{\prime}$ in the same space, we linearly project elements of $\rho$ from $\mathbb{R}$ to $\mathbb{R}^{d}$, and map categorical edge types into $\mathbb{R}^{d}$ using a learned embedding, i.e.
+like token embeddings in language models (Devlin et al., 2019).
+Thus, the dataset $D$ is summarized into a matrix of size $\mathbb{R}^{N\times N\times T\times d}$.
+Model architecture  Both the causal structure learner and differential network are implemented with a series of attention-based blocks.
+Since scaled dot-product attention scales quadratically as sequence length (Vaswani et al., 2017), we attend over one dimension at a time.
+For example, given a $N^{2}$ entries in an adjacency matrix, naive self-attention would cost $O(N^{4})$ operations.
+Instead, we apply self-attention along all nodes in the “outgoing edge” direction, with the “incoming edges” as a batch dimension, followed by the opposite – resulting in two operations, each of cost $O(N\times N^{2})$.
+Following Rao et al.
+(2021), we use pre-layer normalization on each self-attention, followed by dropout and residual connections:
+$h\leftarrow h+\text{Dropout}(\text{Self-Attn}(\text{LayerNorm}(h)))$ (4)
+where $h\in\mathbb{R}^{\dots\times d}$ denotes an intermediate hidden representation.
+This design takes after Transformer models for higher dimensional data in other domains, e.g.
+Axial Transformers for videos (Ho et al., 2020) and MSA Transformers for aligned protein sequences (Rao et al., 2021).
+A key difference is the graph modality: when making edge or node-level predictions, the model should be equivariant to node labeling.
+This is accomplished by adding randomly-permuted “positional” embeddings to each node and edge representation.
+For more details, please see Appendix A.
+Edge-level outputs  Let $h_{\text{obs}},h_{\text{int}}\in\mathbb{R}^{N\times N\times d}$ denote the final layer representations of each dataset, after pooling over the $T$ dimension.
+These representations are input to the differential network.
+In addition, we introduce a graph prediction head,
+$z_{i,j}$ $=\text{FFN}([h_{i,j},h_{j,i}])$ (logits) (5)
+$P(E_{i,j}=e)$ $=\sigma(z_{i,j})_{e}$ (softmax) (6)
+where $\sigma$ denotes softmax over the set of output edge types (forwards, backwards, or no edge).
+The graph prediction loss $\mathcal{L}_{G}$ is simply the cross entropy loss between the predicted and true edge types, summed over the entire graph.
+3.2 Differential network
+Given a pair of graph representations $h_{\text{obs}},h_{\text{int}}$, the differential network predicts $I$, the set of intervention targets.
+Input featurization  In addition to pairwise representations $h$, we compute node-level statistics (mean, variance), which are mapped to $\mathbb{R}^{d}$ via a feed-forward network, and concatenated along “incoming edge” dimension.
+This results in graph representation $h^{\prime}\in\mathbb{R}^{N\times(N+1)\times d}$.
+We combine the graph representations either through concatenation or subtraction:
+$h_{\text{cat}}$ $=[h^{\prime}_{\text{obs}},h^{\prime}_{\text{int}}]\in\mathbb{R}^{N\times(N+1)\times 2d}$ (7)
+$h_{\text{diff}}$ $=h^{\prime}_{\text{int}}-h^{\prime}_{\text{obs}}\in\mathbb{R}^{N\times(N+1)\times d}.$ (8)
+The former can express richer relationships between the two graphs, while the latter is more efficient on large graphs.
+Model architecture  The differential network closely follows the architecture of the causal structure learner.
+However, since paired graph representation lacks the $T$ dimension (introduced by marginal estimates $E^{\prime}$), the model only needs to attend over two “length” dimensions (rows and columns of adjacency matrix).
+After the attention layers, we mean over all “incoming edges” and linearly project to binary node-level predictions, where a normalized 1 indicates an intervention target, and 0 indicates otherwise.
+The target prediction loss $\mathcal{L}_{I}$ is the binary cross entropy loss between the predicted and true targets, computed independently for each node.
+This is to accommodate interventions with multiple targets, e.g.
+drugs with off-target effects.
+Remark  We take an amortized causal discovery approach because it is hard to specify what assumptions a biological system ought to follow.
+This makes it difficult to provide identifiability guarantees without simplifying aspects of either the data or modeling framework.
+Instead, in Appendix C, we show that the differential network is a well-specified class of models.
+That is, there exist parametrizations of its attention architecture that map predicted graphs and global statistics (for hard and soft interventions) to the set of targets.
+3.3 Implementation details
+Training procedure  The overall training loss is $\mathcal{L}=\mathcal{L}_{G}+\mathcal{L}_{I}$ with L2 weight regularization.
+Both $\mathcal{L}_{G}$ and $\mathcal{L}_{I}$ supervise the causal structure learner, while only the latter supervises the differential network.
+We use the synthetic and real datasets as follows.
   1. 1.
-    Given a dataset $D$, smaller batches are constructed by sub-sampling both examples and variables.
-Heuristics like pairwise correlation are used to select variables which are likely to have causal relationships, to minimize unnecessary computation.
+    Initialize the causal structure learner with pretrained weights from Wu et al. (2024) (except ablations).
   2. 2.
-    Two sets of input features are generated from these batches: global summary statistics like correlation, computed over all variables, and marginal estimates, the outputs of classical causal discovery algorithms (Spirtes et al., 1995) run on small subsets of variables.
+    Train both modules jointly on synthetic data.
   3. 3.
-    Finally, these two sets of features are input to a neural network (“aggregator”), which was trained to map them into causal graphs on synthetic datasets.
-We run this procedure on $D_{\text{obs}}$ and $D_{\text{int}}$ independently, treating both as “observational” datasets.
-To obtain the graph representations for each dataset, we extract the aggregator’s last layer hidden representations before they are collapsed into binary edge predictions, yielding $h_{\text{obs}}$ and $h_{\text{int}}$.
-Finally, we concatenate $h_{\text{obs}}$ and $h_{\text{int}}$ along the hidden dimension, to obtain a paired graph representation of size $N\times N\times 2d$.
-Figure 2: Model overview.
-The observational and interventional datasets are featurized using a pretrained, amortized causal discovery model.
-The pair of graph representations are concatenated and input to an axial attention model.
-After the 2D attention, representations are aggregated over incoming edges to obtain binary, node-level predictions.
-Cdn is trained with a classification objective, over synthetic and/or real data.
-Gray modules are frozen.
-Yellow modules are learned.
-3.2 Differential network
-Given the graph representation, the differential network predicts which nodes were intervened upon (Figure 2, right).
-Its architecture is composed of an axial-attention layer, followed by a linear projection.
-Following Sea, the axial attention layer attends separately along the rows and columns of the graph’s adjacency matrix.
-This operation is equivalent to self-attention along all nodes in the “outgoing edge” direction, with the “incoming edges” as a batch dimension, followed by the opposite.
-We use pre-layer normalization on each self-attention, followed by dropout and residual connections:
-$h=h+\text{Dropout}(\text{Self-Attn}(\text{LayerNorm}(h))).$ (3)
-After the axial attention, we mean over all “incoming edges” and linearly project to binary node-level predictions, where a normalized 1 indicates an intervention target, and 0 indicates otherwise.
-3.3 Implementation and training details
-For the causal featurizer, we used a frozen, pretrained aggregator from (Wu et al., 2024), which corresponded to the FCI-based marginal estimates (Spirtes et al., 1995).
-The differential network was implemented with one axial attention layer, following the same architecture as the pretrained aggregator, with twice the model dimension (paired graphs).
-We swept over the number of layers (1-4) on synthetic data and found that a single layer performed the best.
-In terms of model design, we ablated replacing the attention architecture with a multi-layer perceptron (no communication between nodes, Table 4), as well as other components (Table 7) and summary statistics (Table 8).
-Due to the large graph sizes (adjacency matrices up to $1000^{2}$), standard 1D attention was intractable.
-Following Sea, we used inverse covariance as the global statistic on synthetic data, and correlation on transcriptomics data.
-We trained separate differential networks on synthetic and real data.
-On synthetic data, we used a batch size of 16 and the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 1e-4 (identical to Sea).
-On real data, we started from the synthetic Cdn checkpoint, and then used a batch size of 2 with a learning rate of 1e-5, due to the larger graph sizes.
-All models took around 4-10 hours to train on a single A6000 GPU.
-3.4 Theoretical context
-This paper primarily describes a causality-inspired approach towards perturbation target prediction.
-Contrary to causal discovery algorithms that are self-contained and provably sound under standard assumptions, we cannot “prove” that a pretrained model extracts correct graphs on real data, as the true graphs are unknown.
-Our evaluation focuses solely on perturbation target prediction, as the ground truth is known, by (experimental) design.
-To contextualize the predictions of our model, we provide brief sketches with regards to our model’s capacity to map paired graph features to intervention targets.
-We defer formal claims to Appendix A.
-Claim 3.1 (Hard interventions, informal).
-Given a causal graphical model $G=(V,E)$ and intervention targets $I\subseteq V$, let $E^{\prime}$ denote the adjacency matrix of mutilated graph $G^{\prime}$ after hard intervention on $I$.
-An axial attention layer is well-specified to map $(E,E^{\prime})$ to $I$.
-Claim 3.2 (Soft interventions, informal).
-Let $G$ be a graphical causal model associated with data distribution $P_{X}$, and let $\tilde{P}_{X}$ be the interventional distribution after soft intervention on $I\subsetneq V$.
-Let $R,R^{\prime}$ denote the correlation matrices and let $\Sigma,\Sigma^{\prime}$ denote the covariance matrices of $X,\tilde{X}$.
-An axial attention layer is well-specified to map $(R,R^{\prime},\Sigma,\Sigma^{\prime})$ to $I$.
-4 Experiments and results
-Table 1: Final biological dataset statistics. “gw” and “es” are two independent screens with different perturbations.
-Type Cell line ##\## Perts ##\## Cells
-Genetic K562 gw 1767 492,096
-K562 es 1060 213,552
-RPE1 961 179,696
-HepG2 627 79,309
-Jurkat 1014 174,698
-Chemical A172 3 18,196
-T98G 3 13,126
-Figure 3: Data illustration and splitting procedure.
+    Finetune both modules jointly with only $\mathcal{L}_{I}$ (no ground truth graphs), on each training split of the real datasets.
+For the synthetic training data, we generated 8640 datasets with hard and soft interventions, which vary in causal mechanism, intervention type, and graph topology.
+To emulate transcriptomics data, in which perturbation effect sizes are quantified in (log) fold change, we introduced soft interventions: “shift” $x\leftarrow f(\pi_{x})+c$, where $c\in\mathbb{R}$, and “scale” $x\leftarrow cf(\pi_{x})$, where $c$ is a positive scaling factor, with equal probability $c\lessgtr 1$.
+We also sampled interventions with multiple targets, to simulate off-target effects.
+Details are available in Section 4.2 and Appendix D.2.
+Model parameters  Marginal estimates $E^{\prime}$ are computed using the classic FCI algorithm (Spirtes et al., 1995), and global statistic $\rho$ is pairwise correlation, as it is fast to compute and numerically stable.
+We swept over the number of differential network layers (Figure 4) on synthetic data, and we used 3 layers for $h_{\text{cat}}$ and 2 layers for $h_{\text{diff}}$.
+Following Sea, we adopted hidden dimension $d=64$, the AdamW optimizer (Loshchilov & Hutter, 2019), learning rate 1e-4, batch size 16, and weight decay 1e-5.
+On the real data, where $N=1000$, we changed to a batch size of 1, decreased the learning rate to 5e-6, and finetuned the models with half precision (FP16).
+We trained both $h_{\text{cat}}$ and $h_{\text{diff}}$ models on synthetic data, but only finetuned the $h_{\text{diff}}$ architecture on the real data (memory constraint).
+4 Experiments
 We evaluated Cdn on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1).
-Note that these baselines all use various domain knowledge to inform their predictions.
 For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
 4.1 Biological experiments
-Datasets
-We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al.
+Datasets  We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al.
 (2022) and Nadig et al.
@@
 (2024).
-Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$ (Figure 3A).
-In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target.
-To ensure high quality labels, we filtered perturbations to those that induced over 10 differentially-expressed genes (statistically significant change, compared to control), of which the true target should be present.
-This is to exclude perturbations with insufficient cells (low statistical power), with minimal to no effect (uninteresting), and those that did not achieve the desired effect (CRISPR efficacy is not guaranteed).
-For evaluation, we limited the set of candidate targets to the top 1000 differentially expressed genes, ranked by log-fold change per perturbation.
-If fewer genes were differentially expressed, we randomly sampled additional candidates until we reached a minimum of 100 genes.
-Finally, we also stratify genetic perturbations based on whether the target is trivially identifiable as the gene with the largest log-fold change (“trivial” vs.
-“non-trivial”).
-This is because genetic perturbations are highly specific by design, but also constitute the largest single-cell perturbation datasets.
-The final data statistics are shown in Table 1, with additional details in Appendix B.1.
-Evaluation
-We consider two splits: seen and unseen cell lines (Figure 3B).
+Each dataset is a real-valued matrix of gene expression levels: the number of examples $M$ is the number of cells, the number of variables $N$ is the number of genes, and each entry is a log-normalized count of how many copies of gene $j$ was measured from cell $i$.
+In Perturb-seq datasets, we aim to recover the gene whose promoter was targeted by the CRISPR guide, and in Sci-Plex datasets, we aim to identify the gene that corresponds to the drug’s intended target (discussion in Appendix D.1).
+Biologically, many perturbations do not affect the cell’s production of mRNA (they may affect other factors like protein activity, which transcriptomics does not measure).
+To ensure that $P_{\text{obs}}$ and $P_{\text{int}}$ are distinct, we filtered perturbations to those that induced over 10 differentially-expressed genes (statistically significant change, compared to control), of which the true target should be present.
+This excludes perturbations with insufficient cells (low statistical power), with minimal to no effect (uninteresting), and those that did not achieve the desired effect (low CRISPR efficiency).
+In the Sci-Plex data, several drugs have known off-targets, but we did not observe differential expression of any of them (adjusted p-value $\approx 1$).
+Evaluation  We consider two splits: seen and unseen cell lines.
 In the former, models may be trained on approximately half of the perturbations from each cell line, and are evaluated on the unseen perturbations.
 In the latter, we hold out one cell line at a time, and models may be trained on data from the remaining cell lines.
-Note that not all baselines can be evaluated on unseen cell lines.
-To ensure that our train and test splits are sufficiently distinct, we cluster perturbations based on their log-fold change and assign each cluster to the same split.
-While perturbation target prediction appears to be a simple classification task, there are two aspects of the data that render standard metrics less meaningful.
-Not all genes can be mapped to the baselines’ domain knowledge, so their effects as perturbations cannot be predicted.
+To ensure that our train and test splits are sufficiently distinct, we cluster perturbations based on their log-fold change and assign each cluster to the same split (Figure 5).
+Since the Sci-Plex data are limited, we use them only as test sets without finetuning.
+We limited the set of candidate targets to the top 1000 differentially expressed genes by log-fold change, per perturbation.
+If too few genes passed the significance threshold ($p<0.05$), we added additional candidates with the largest log-fold change until we reached a minimum of 100 genes.
+Genetic perturbations have highly specific effects compared to chemical perturbations, so we focused on the subset for which the intended target was not trivially identifiable as the gene with the largest log-fold change.
+[^1] Cdn predicts trivial perturbations nearly perfectly (Table 8).
+While perturbation target prediction appears to be a simple classification task, standard metrics are not well-suited for these real data.
+Not all genes are present in the baselines’ domain knowledge (e.g.
+isolated node in graph), so their effects as perturbations cannot be predicted.
 In addition, due to genetic redundancy, it is common for multiple perturbations to elicit similar responses (Kernfeld et al., 2023).
-Therefore, an “incorrectly” predicted target (e.g.
-out of 1000 genes) may not necessarily indicate poor performance.
-Therefore, we propose the following metrics.
-  • We report the rank of the ground truth target, normalized by the number of candidate genes.
-Here, 1 indicates that the ground truth target is at the top of the list (best), while 0 indicates that the ground truth is at the bottom (worst).
-  • To emulate a “virtual screening” setting, we report recall at k (the proportion of targets recovered within the top $k$ candidates). This value ranges from 0 to 1, where 1 is best.
-  • The Pearson correlation $r$ measures the similarity between the (ground truth) mean log-fold change of the top prediction that was observed as a perturbation, and that of the actual perturbation (given as input).
-These values range from -1 to 1, where 1 is best.
-Baselines
-We compare to two classes of algorithms related to biological perturbations.
-Forward inference models predict the effects of perturbations, and perturbation targets can be ranked by comparing the inferred effect of each candidate to the interventional data.
-  • Gears (Roohani et al., 2023) is a graph neural network that predicts the effects of unseen genetic perturbations, using the gene ontology knowledge graph (Ashburner et al., 2000) to model the relationship between genes.
-We trained Gears on each cell line separately, and predicted the effects of perturbing every expressed gene.
-Then we ranked candidates based on cosine similarity to the interventional data.
-  • GenePt (Chen & Zou, 2023) is a set of natural language-based embeddings that have been shown to achieve state-of-the-art performance on unseen perturbation effect prediction with simple downstream models (Märtens et al., 2024).
-We concatenated each perturbation’s gene embedding to the log-fold change of the top 5000 highly-variable genes (Wolf et al., 2018) and trained a logistic regression model on each cell line to predict true vs.
-decoy perturbations.
-Candidates were ranked based on predicted probability.
-Predictive models directly predict perturbation targets from data as a classification task.
+An “incorrect” top 1 prediction may not necessarily reflect poor performance.
+Thus, we propose the following metrics.
+  • We compute the rank of the true target, normalized by the number of candidate genes. Rank ranges from 0 (worst, bottom of list) to 1 (best, top of list).
+  • We measure the similarity between the perturbation effect (log-fold change) of the top 1 predicted target and the true target with the Spearman rank correlation $\rho$ and the Pearson correlation $r$.
+These range from -1 to 1, where 1 is best.
+  • To emulate “virtual screening,” we plot recall at p: the fraction of targets recovered within the $p\in(0,1)$ candidates. Recall ranges from 0 to 1, where 1 is best.
+Baselines  We compare to both algorithms that infer the effects of perturbations on cells (used as scoring functions, Equation 3), and algorithms that directly predict targets.
+All baselines were run with their official implementations and/or latest releases.
+For details, please see Appendix B.2.
+Gears (Roohani et al., 2023) and GenePt (Chen & Zou, 2023) predict the effects of unseen genetic perturbations.
+Gears is a graph neural network that regresses log-fold change upon perturbation, using the Gene Ontology knowledge graph as an undirected backbone.
+GenePt is a set of large language model-derived embeddings that have been shown to achieve state-of-the-art performance (Märtens et al., 2024) with simple downstream models like logistic regression (used here).
+Both are trained per cell line.
+PDGrapher (Gonzalez et al., 2024) is a graph neural network that predicts the perturbation targets of genetic or chemical perturbations, on seen or unseen cell lines, with the Human Reference Interactome (Luck et al., 2020) as the knowledge graph.
+To the best of our knowledge, PDGrapher is the only published deep learning baseline for the target prediction task on transcriptomic data.
+Therefore, we compare against additional statistical and naive baselines.
+Dge stands for differential gene expression (Love et al., 2014), i.e.
+the process of identifying statistically significant changes in genes between settings.
+Specifically, we run the Wilcoxon ranked-sum test with Benjamini-Hochberg correction (Wilcoxon, 1945; Benjamini & Hochberg, 2000) between control and perturbed cells, as implemented by Wolf et al.
+(2018).
+Genes are ranked by adjusted p-value.
 Linear and Mlp take as input the mean expression of all perturbation targets, plus the top 2000 highly-variable genes (Wolf et al., 2018).
 They are trained to predict a binary label for each gene on each cell line independently.
-  • PDGrapher (Gonzalez et al., 2024) is a graph neural network that predicts the perturbation targets of genetic or chemical perturbations, on seen or unseen cell lines, with the human reference interactome (Luck et al., 2020) as the knowledge graph.
-All baselines were run with their official implementations and/or latest releases. For more details, please see Appendix C.1.
 Table 1: Results on 5 Perturb-seq datasets.
-Top: “trivial” perturbations.
-Bottom: “non-trivial” perturbations.
-Primary setting: train all cell lines jointly, test on unseen perturbations.
-Suffix -cl: leave out the one cell line from training, test on the same set of unseen perturbations in the unseen cell line.
-Metrics, from left to right: normalized rank of ground truth, top 20 recall, and top 1 Pearson correlation.
-For ablations, runtimes, and uncertainties, see Appendix D.
-K562 gw K562 es RPE1 HepG2 Jurkat
-Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
-Random $0.50$ $0.13$ $0.16$ $0.50$ $0.09$ $0.17$ $0.50$ $0.09$ $0.36$ $0.50$ $0.06$ $0.34$ $0.50$ $0.08$ $0.22$
-Linear $0.49$ $0.16$ $0.21$ $0.50$ $0.13$ $0.31$ $0.48$ $0.09$ $0.45$ $0.48$ $0.07$ $0.43$ $0.48$ $0.08$ $0.29$
-Mlp $0.48$ $0.21$ $0.20$ $0.47$ $0.13$ $0.21$ $0.54$ $0.19$ $0.37$ $0.51$ $0.10$ $0.42$ $0.48$ $0.13$ $0.32$
-GenePt $0.60$ $0.39$ $0.39$ $0.47$ $0.11$ $0.26$ $0.57$ $0.33$ $0.55$ $0.44$ $0.12$ $0.41$ $0.55$ 0.32 0.53
-Gears $0.49$ $0.14$ $0.18$ $0.51$ $0.09$ $0.27$ $0.49$ $0.09$ $0.40$ $0.48$ $0.09$ $0.40$ $0.47$ $0.06$ $0.24$
-PdG $0.49$ $0.20$ $0.26$ $0.52$ $0.18$ $0.34$ $0.51$ $0.11$ $0.42$ $0.46$ $0.06$ $0.40$ $0.49$ $0.11$ $0.35$
-Cdn 0.77 0.52 0.49 0.68 0.32 0.45 0.74 0.37 0.55 0.68 0.20 0.50 0.67 $0.27$ $0.45$
-PdG-Cl $0.40$ $0.04$ $0.15$ $0.41$ $0.02$ $0.24$ $0.48$ $0.02$ $0.37$ $0.47$ $0.03$ $0.41$ $0.53$ $0.06$ $0.33$
-Cdn-Cl 0.71 0.43 0.40 0.66 0.27 0.42 0.62 0.24 0.53 0.66 0.17 0.46 0.65 0.26 0.47
-Random $0.50$ $0.11$ $0.17$ $0.50$ $0.04$ $0.17$ $0.50$ $0.04$ $0.44$ $0.50$ $0.04$ $0.33$ $0.50$ $0.06$ $0.21$
-Linear $0.51$ $0.18$ $0.23$ $0.48$ $0.01$ $0.25$ $0.49$ $0.03$ $0.52$ $0.51$ $0.04$ 0.43 $0.50$ $0.06$ $0.23$
-Mlp $0.45$ $0.19$ $0.21$ $0.49$ $0.04$ $0.11$ $0.53$ $0.09$ $0.41$ $0.48$ 0.07 $0.40$ $0.45$ $0.07$ $0.29$
-GenePt $0.52$ $0.24$ $0.32$ $0.42$ $0.03$ $0.19$ $0.41$ $0.06$ $0.46$ $0.31$ $0.03$ $0.40$ $0.51$ 0.15 $0.36$
-Gears $0.56$ $0.16$ $0.22$ $0.45$ $0.06$ $0.20$ $0.49$ $0.04$ $0.44$ $0.54$ $0.03$ $0.40$ $0.50$ $0.11$ $0.26$
-PdG $0.54$ $0.24$ $0.29$ $0.49$ $0.03$ $0.23$ $0.48$ $0.05$ $0.52$ $0.54$ $0.05$ 0.43 $0.52$ $0.10$ $0.35$
-Cdn 0.72 0.37 0.32 0.56 0.15 0.35 0.59 0.12 0.53 0.55 0.07 $0.40$ 0.62 $0.10$ 0.39
-PdG-Cl $0.39$ $0.02$ $0.13$ $0.44$ $0.01$ 0.28 0.54 $0.02$ 0.48 $0.48$ $0.02$ 0.41 $0.57$ $0.06$ $0.35$
-Cdn-Cl 0.68 0.25 0.29 0.55 0.06 $0.27$ $0.49$ 0.07 0.48 0.59 0.05 $0.39$ 0.60 0.18 0.38
-Figure 4: Recall at $k$ on K562 datasets.
-Table 3: Rank of drug targets on chemical perturbation datasets (McFaline-Figueroa et al., 2024). A172 and T98G are unseen cell lines. For details and full drug names, please see Appendix B.1.
+Top: Train on all cell lines jointly.
+Bottom: Leave one cell line out of training.
+Test sets (unseen perturbations) are identical in both settings.
+Cdn (synthetic) and Dge do not require training on real Perturb-seq data.
+Metrics, from left to right: normalized rank of ground truth, top 1 Spearman and Pearson correlations.
+Uncertainty quantification in Table 7.
+Runtimes in Table 11.
+K562 (gw) K562 (es) RPE1 HepG2 Jurkat
+Setting Model rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$
+Seen cell line Linear $0.51$ $0.18$ $0.23$ $0.48$ $0.22$ $0.25$ $0.49$ $0.42$ $0.52$ $0.51$ $0.35$ $0.43$ $0.50$ $0.20$ $0.23$
+Mlp $0.45$ $0.18$ $0.21$ $0.49$ $0.08$ $0.11$ $0.53$ $0.32$ $0.41$ $0.48$ $0.33$ $0.40$ $0.45$ $0.24$ $0.29$
+GenePt $0.52$ $0.29$ $0.32$ $0.42$ $0.16$ $0.19$ $0.41$ $0.37$ $0.46$ $0.31$ $0.32$ $0.40$ $0.51$ $0.31$ $0.36$
+Gears $0.56$ $0.19$ $0.22$ $0.45$ $0.18$ $0.20$ $0.49$ $0.35$ $0.44$ $0.54$ $0.32$ $0.40$ $0.50$ $0.22$ $0.26$
+PdG $0.54$ $0.25$ $0.29$ $0.49$ $0.20$ $0.23$ $0.48$ $0.41$ $0.52$ $0.54$ $0.37$ $0.43$ $0.52$ $0.31$ $0.35$
+Cdn (Perturb-seq) 0.92 0.72 0.74 0.95 0.81 0.82 0.92 0.67 0.76 0.94 0.69 0.75 0.98 0.83 0.84
+Unseen cell line Dge $0.82$ $0.50$ $0.54$ $0.89$ $0.60$ $0.62$ $0.79$ $0.56$ $0.64$ $0.86$ $0.51$ $0.51$ $0.84$ $0.45$ $0.50$
+PdG $0.39$ $0.10$ $0.13$ $0.44$ $0.24$ $0.28$ $0.54$ $0.38$ $0.48$ $0.48$ $0.34$ $0.41$ $0.57$ $0.31$ $0.35$
+Cdn (synthetic) $0.79$ $0.46$ $0.49$ $0.80$ $0.59$ $0.62$ $0.82$ $0.61$ $0.71$ $0.84$ $0.56$ $0.63$ $0.87$ $0.64$ $0.67$
+Cdn (Perturb-seq) 0.91 0.69 0.71 0.95 0.81 0.82 0.92 0.65 0.74 0.93 0.67 0.72 0.97 0.84 0.85
+Figure 3: Recall at $p$ (normalized over number of candidate genes on 5 Perturb-seq datasets. For Cdn, (P) and (S) denote “Perturb-seq” and “synthetic” training data.
+Table 2: Rank of drug targets on chemical perturbation datasets (McFaline-Figueroa et al., 2024).
+A172 and T98G are unseen cell lines.
+Highlighted: Within top 100.
+Drug names and details in Appendix D.1.
 A172 T98G
 Model infig. nint. palb. doxo. palb. vola.
-PdG $0.719$ $0.346$ $0.395$ $0.055$ $0.345$ $0.420$
-Cdn 0.883 0.436 0.582 0.059 0.543 0.647
-Results
-Table 2 reports our results on five Perturb-seq datasets.
-In the majority of cases, Cdn outperforms baselines, both in the seen and unseen cell line settings.
-This is more evident in Figure 4, in which Cdn achieves higher recall at $k$ at nearly all points on the curves.
-While no baseline is as consistent as Cdn, GenePt is surprisingly competitive against more complex baselines.
-This may indicate that natural language induces embedding spaces with favorable geometries with respect to biological function.
-On the two chemical perturbation datasets, Cdn ranks the ground truth targets higher than PdGrapher on all six drugs (Table 3).
-However, the mutually low performance on doxorubicin (doxo.) in T98G cells may indicate a failure mode of both models.
+Dge $0.12$ $0.64$ $0.28$ $0.94$ $0.24$ $0.83$
+PdG $0.72$ $0.35$ $0.39$ $0.06$ $0.34$ $0.42$
+Cdn $0.14$ $0.81$ $0.01$ $0.89$ $0.93$ $0.53$
+Cdn (no $\mu$) $0.88$ $0.44$ $0.58$ $0.06$ $0.54$ $0.65$
+Results  Cdn consistently outperforms all baselines on the five Perturb-seq datasets, in both the seen and unseen cell line settings (Table 1).
+There is minimal drop in between seen and unseen cell lines, suggesting that the functional characteristics of genetic perturbations are highly transferrable, even to different data distributions.
+Notably, no other deep learning method exceeds simple statistical tests (Dge).
+This is more evident in Figure 3, in which Cdn achieves higher recall at $p$ at nearly all points.
+Finally, if we ablate the finetuning step (“synthetic” vs.
+“Perturb-seq”), we find that finetuning leads to significant improvements.
+Chemical perturbations are much less specific than genetic perturbations (Figures 6 and 7), and it is less clear whether drug effects should be reflected in mRNA levels (via feedback mechanisms), as drugs act upon their protein products.
+Still, on the two chemical perturbation datasets, Cdn finds the true target within the top 100 candidates for 3/6 cases (Table 2).
+Due to the low performance of Dge on several cases, we hypothesized that gene-level changes may not be as predictive.
+Indeed, an ablation of Cdn, trained without node-level statistics, improves where Dge fails (“no $\mu$”).
 4.2 Synthetic experiments
-There are two considerations that motivate further experiments in controlled settings.
-First, while it would be ideal to evaluate all algorithms on real data, current causal discovery algorithms that support unknown interventions are not tractable on datasets with more than tens of variables.
+While it would be ideal to evaluate all algorithms on real data, current causal discovery algorithms that support unknown interventions are not tractable on datasets with more than tens of variables.
 In fact, many baselines require hours on even $N=10$ datasets, and they do not scale favorably (Table 11).
-Our transcriptomics datasets contain hundreds of genes, even after filtering to those that are differentially expressed (Figures 6 and 7).
-On the other hand, existing models for biological perturbations all rely on some form of domain knowledge, and their performance is inseparable from the choice and quality of these external data.
-Here, synthetic settings allow us to assess the model’s capacity to predict intervention targets in isolation.
-Datasets
-We generated 120 synthetic test datasets with hard and soft interventions.
+Our transcriptomics datasets contain hundreds of genes, even after filtering to those that are differentially expressed (Figure 8).
+At the same time, existing models for biological perturbations all rely on some form of domain knowledge, and their performance is inseparable from the choice and quality of these external data.
+Synthetic data allow us to assess the model’s capacity to predict intervention targets in isolation.
+Datasets  Each evaluation setting consists of 5 unseen, i.i.d.
+sampled graphs.
+To assess the generalization capacity of our framework, we tested the model on seen (linear) and unseen causal mechanisms (polynomial, sigmoid), and unseen soft interventions.
 To generate observational data, we sampled Erdős-Rényi graphs with $N=10,20$ nodes and $E=N,2N$ expected edges; causal mechanism parameters; and observations of each variable, in topological order.
 We sampled $3N$ distinct subsets of 1-3 nodes ($N$ each) as intervention targets.
-For hard interventions, we set $x\leftarrow z$, where $z$ is uniform.
-To emulate transcriptomics data, in which perturbation effects are measured in fold change, we introduce soft interventions: $x\leftarrow cf(\pi_{x})$, where $c$ is a positive scaling factor, with equal probability $c\lessgtr 1$.
-For our synthetic training set, we generated approximately 4000 datasets following the training set of Sea, with hard interventions only.
-Full details are available in Appendix B.2.
-Baselines
-We compare against discrete and continuous causal discovery algorithms for unknown interventions.
+Hard interventions set $x\leftarrow z$, where $z$ is uniform.
+For soft interventions (Section 3.3), we trained on shift and tested on scale for synthetic experiments (see Table 9 for explanation).
+Real experiments trained on all interventions.
+Table 3: Intervention target prediction results on synthetic datasets with $N=20,E=40$.
+Top: hard interventions (uniform); bottom: soft interventions (scale).
+Neither polynomial nor “scale” soft interventions were seen during training.
+Number in parentheses indicates number of intervention targets.
+Uncertainty is standard deviation over 5 i.i.d.
+datasets.
+Runtimes in Table 11.
+Extended results in Table 10.
+Linear (1) Linear (3) Polynomial (1) Polynomial (3)
+[…]
+Hard Ut-Igsp $10$01 $0.63\scriptstyle\pm.03$ $0.18\scriptstyle\pm.01$ $0.52\scriptstyle\pm.03$ $0.19\scriptstyle\pm.02$ $0.70\scriptstyle\pm.02$ $0.23\scriptstyle\pm.03$ $0.56\scriptstyle\pm.03$
+Dcdi-G $0.16\scriptstyle\pm.02$ $0.50\scriptstyle\pm.02$ $0.27\scriptstyle\pm.05$ $0.50\scriptstyle\pm.04$ $0.15\scriptstyle\pm.03$ $0.45\scriptstyle\pm.04$ $0.26\scriptstyle\pm.02$ $0.51\scriptstyle\pm.04$
+Dcdi-Dsf $0.15\scriptstyle\pm.02$ $0.46\scriptstyle\pm.03$ $0.26\scriptstyle\pm.03$ $0.47\scriptstyle\pm.04$ $0.18\scriptstyle\pm.05$ $0.48\scriptstyle\pm.08$ $0.28\scriptstyle\pm.04$ $0.52\scriptstyle\pm.05$
+Bacadi-E $0.18\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ $0.28\scriptstyle\pm.02$ $0.69\scriptstyle\pm.03$ $0.20\scriptstyle\pm.06$ $0.84\scriptstyle\pm.04$ $0.37\scriptstyle\pm.07$ $0.82\scriptstyle\pm.05$
+Bacadi-M $0.10\scriptstyle\pm.02$ 0.66±.05plus-or-minus0.66.050.$6$05 $0.21\scriptstyle\pm.01$ $0.63\scriptstyle\pm.02$ $0.09\scriptstyle\pm.02$ $0.71\scriptstyle\pm.05$ $0.24\scriptstyle\pm.03$ $0.70\scriptstyle\pm.05$
+Dci $0.58\scriptstyle\pm.07$ $0.85\scriptstyle\pm.04$ $0.63\scriptstyle\pm.06$ $0.84\scriptstyle\pm.03$ 0.55±.08plus-or-minus0.55.080.$5$08 $0.82\scriptstyle\pm.04$ $0.56\scriptstyle\pm.08$ $0.79\scriptstyle\pm.05$
+Mb+Ci $0.47\scriptstyle\pm.12$ $0.68\scriptstyle\pm.08$ $0.58\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ 0.77±.05plus-or-minus0.77.050.$7$05 0.88±.03plus-or-minus0.88.030.$8$03 $0.79\scriptstyle\pm.04$ 0.88±.02plus-or-minus0.88.020.$8$02
+[…]
+Cdn (no $\mathcal{L}_{G}$) $0.65\scriptstyle\pm.07$ $0.86\scriptstyle\pm.06$ $0.72\scriptstyle\pm.09$ $0.84\scriptstyle\pm.07$ $0.81\scriptstyle\pm.05$ $0.95\scriptstyle\pm.02$ $0.86\scriptstyle\pm.03$ $0.94\scriptstyle\pm.01$
+Soft Ut-Igsp $10$01 $0.69\scriptstyle\pm.03$ $0.19\scriptstyle\pm.01$ $0.56\scriptstyle\pm.03$ $0.18\scriptstyle\pm.02$ 0.77±.02plus-or-minus0.77.020.$7$02 0.22±.01plus-or-minus0.22.010.$2$01 $0.60\scriptstyle\pm.01$
+Dcdi-G $0.17\scriptstyle\pm.03$ $50$05 $0.23\scriptstyle\pm.01$ 0.44±.03plus-or-minus0.44.030.$4$03 $0.14\scriptstyle\pm.03$ $0.48\scriptstyle\pm.03$ $0.30\scriptstyle\pm.04$ $0.54\scriptstyle\pm.04$
+Dcdi-Dsf $0.17\scriptstyle\pm.04$ $0.46\scriptstyle\pm.04$ $0.24\scriptstyle\pm.02$ $0.47\scriptstyle\pm.03$ $0.16\scriptstyle\pm.02$ $50$05 $0.26\scriptstyle\pm.02$ $0.48\scriptstyle\pm.02$
+Bacadi-E $0.25\scriptstyle\pm.07$ $0.65\scriptstyle\pm.03$ $0.40\scriptstyle\pm.13$ $0.68\scriptstyle\pm.08$ $0.43\scriptstyle\pm.22$ $0.81\scriptstyle\pm.09$ $0.49\scriptstyle\pm.15$ 0.77±.05plus-or-minus0.77.050.$7$05
+Bacadi-M 0.11±.01plus-or-minus0.11.010.$1$01 $0.61\scriptstyle\pm.02$ $0.26\scriptstyle\pm.07$ $0.64\scriptstyle\pm.06$ $0.25\scriptstyle\pm.10$ $0.80\scriptstyle\pm.09$ $0.38\scriptstyle\pm.10$ $0.76\scriptstyle\pm.06$
+Dci $0.48\scriptstyle\pm.02$ $0.78\scriptstyle\pm.02$ $0.53\scriptstyle\pm.06$ 0.77±.04plus-or-minus0.77.040.$7$04 $0.62\scriptstyle\pm.09$ $0.89\scriptstyle\pm.04$ 0.55±.04plus-or-minus0.55.040.$5$04 $0.80\scriptstyle\pm.04$
+Mb+Ci $0.46\scriptstyle\pm.10$ $0.65\scriptstyle\pm.06$ $0.49\scriptstyle\pm.04$ $0.64\scriptstyle\pm.03$ $\textbf{0.97}\scriptstyle\pm.02$ $0.98\scriptstyle\pm.01$ $0.92\scriptstyle\pm.05$ $0.94\scriptstyle\pm.03$
+[…]
+[…]
+Baselines  We compare against discrete and continuous causal discovery algorithms for unknown interventions.
 Ut-Igsp (Squires et al., 2020) infers causal graphs and unknown targets by greedily selecting the permutation of variables that minimizes their proposed score function.
 Dcdi (Brouillard et al., 2020) and Bacadi (Hägele et al., 2023) are continuous causal discovery algorithms that fit generative models to the data, where the causal graph and intervention targets are model parameters.
-The -G and -Dsf suffixes on Dcdi correspond to Gaussian and deep sigmoidal flow parametrizations of the likelihood.
-Bacadi is evaluated here using the fully-connected implementation (better performance), with the linear version in Table 11.
-Its -E and -M suffixes indicate empirical (standard) and mixture (bootstrap) variants.
-We also compare against the difference causal inference (Dci) algorithm with stability selection (Belyaeva et al., 2021), a discrete optimization algorithm that aims to detect edge-level differences between two causal graphs.
-On this node-centric task, we take each node’s proportion of changed edges as its likelihood of being an intervention target.
-While Dci was also motivated by biological applications, it only scales to around a hundred nodes at most, so we evaluate it alongside other causal discovery methods here.
-Evaluation
-In the synthetic case, we are not constrained by biological redundancy or incomplete predictions, so we report standard classification metrics.
-  • Mean average precision (mAP) is the average precision over a continuum of binarization thresholds, computed independently for each variable, averaged over all regimes of the same number of targets.
-This ranges from 0 to 1 (perfect), where the random baseline depends on the positive rate.
-  • Area under the receiver operating curve (AUC) is also computed independently for each variable, averaged over the same set.
-This ranges from 0 to 1 (perfect), where the random baseline is 0.5 (per edge).
-Results
-Table 4: Intervention target prediction results on synthetic datasets with $N=10,E=10$.
-Uncertainty is standard deviation over 5 distinct datasets.
-Top: hard interventions; bottom: soft interventions (“scale”).
-Number in parentheses indicates number of intervention targets.
-Runtimes are documented in Table 9, and full results are available in Table 11.
-Linear (1) Linear (3) Polynomial (1) Polynomial (3)
-[…]
-Hard Ut-Igsp $0.19\scriptstyle\pm.03$ $0.56\scriptstyle\pm.05$ $0.38\scriptstyle\pm.04$ $0.57\scriptstyle\pm.05$ $0.27\scriptstyle\pm.05$ $0.63\scriptstyle\pm.03$ $0.38\scriptstyle\pm.03$ 0.55±.03plus-or-minus0.55.030.$5$03
-Dcdi-G $0.38\scriptstyle\pm.12$ $0.59\scriptstyle\pm.11$ $0.47\scriptstyle\pm.03$ $0.53\scriptstyle\pm.03$ $0.27\scriptstyle\pm.09$ $0.47\scriptstyle\pm.07$ $0.45\scriptstyle\pm.04$ $0.49\scriptstyle\pm.05$
-Dcdi-Dsf $0.32\scriptstyle\pm.05$ $0.58\scriptstyle\pm.10$ $0.43\scriptstyle\pm.03$ $0.48\scriptstyle\pm.03$ $0.25\scriptstyle\pm.04$ $0.42\scriptstyle\pm.05$ $0.46\scriptstyle\pm.03$ $50$05
-Dci $0.42\scriptstyle\pm.07$ $0.75\scriptstyle\pm.03$ $0.63\scriptstyle\pm.11$ $80$08 $0.39\scriptstyle\pm.10$ $0.72\scriptstyle\pm.06$ $0.53\scriptstyle\pm.09$ $0.69\scriptstyle\pm.07$
-BaCaDI-E $0.16\scriptstyle\pm.04$ $0.61\scriptstyle\pm.09$ $0.36\scriptstyle\pm.04$ $60$06 $0.13\scriptstyle\pm.02$ $0.62\scriptstyle\pm.05$ $0.36\scriptstyle\pm.03$ $0.60\scriptstyle\pm.04$
-BaCaDI-M $0.12\scriptstyle\pm.01$ $0.57\scriptstyle\pm.05$ 0.33±.02plus-or-minus0.33.020.$3$02 $0.56\scriptstyle\pm.04$ $0.12\scriptstyle\pm.01$ $0.57\scriptstyle\pm.04$ 0.33±.02plus-or-minus0.33.020.$3$02 0.55±.04plus-or-minus0.55.040.$5$04
-[…]
-[…]
-Soft Ut-Igsp $20$02 $0.71\scriptstyle\pm.04$ $0.37\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $0.24\scriptstyle\pm.03$ $0.72\scriptstyle\pm.03$ $0.39\scriptstyle\pm.04$ $0.58\scriptstyle\pm.05$
-Dcdi-G $0.38\scriptstyle\pm.11$ $0.58\scriptstyle\pm.11$ $0.48\scriptstyle\pm.05$ $0.50\scriptstyle\pm.07$ $0.25\scriptstyle\pm.06$ $0.49\scriptstyle\pm.08$ $0.48\scriptstyle\pm.05$ $0.53\scriptstyle\pm.06$
-Dcdi-Dsf $0.37\scriptstyle\pm.09$ $0.59\scriptstyle\pm.11$ $0.45\scriptstyle\pm.05$ $0.51\scriptstyle\pm.06$ $0.32\scriptstyle\pm.09$ $0.53\scriptstyle\pm.13$ $0.43\scriptstyle\pm.04$ $0.47\scriptstyle\pm.04$
-Dci $0.45\scriptstyle\pm.10$ $\textbf{0.73}\scriptstyle\pm.07$ $0.54\scriptstyle\pm.04$ $\textbf{0.68}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.03$ $0.69\scriptstyle\pm.03$ $0.45\scriptstyle\pm.03$ $0.62\scriptstyle\pm.01$
-BaCaDI-E $0.34\scriptstyle\pm.16$ $0.63\scriptstyle\pm.10$ $0.49\scriptstyle\pm.14$ $0.62\scriptstyle\pm.15$ 0.22±.03plus-or-minus0.22.030.$2$03 $0.64\scriptstyle\pm.05$ $0.38\scriptstyle\pm.06$ $0.58\scriptstyle\pm.06$
-BaCaDI-M $0.28\scriptstyle\pm.11$ $0.62\scriptstyle\pm.10$ 0.44±.14plus-or-minus0.44.140.$4$14 $0.59\scriptstyle\pm.15$ $0.21\scriptstyle\pm.02$ $0.64\scriptstyle\pm.05$ $0.38\scriptstyle\pm.06$ $0.58\scriptstyle\pm.06$
-Cdn (Mlp) $0.29\scriptstyle\pm.04$ $0.53\scriptstyle\pm.05$ 0.44±.04plus-or-minus0.44.040.$4$04 $0.47\scriptstyle\pm.02$ 0.33±.09plus-or-minus0.33.090.$3$09 $0.50\scriptstyle\pm.09$ $0.52\scriptstyle\pm.06$ $0.56\scriptstyle\pm.05$
-[…]
-On synthetic data, Cdn achieves high performance across intervention types and data-generating mechanisms (Table 4), while running in seconds (Table 9).
-As an ablation study, we investigated replacing the differential network with a multi-layer perceptron, which does not model interactions between edges.
-On the hard interventions, the Mlp is sufficient for predicting intervention targets, perhaps by abusing marginal variances (Reisach et al., 2021) or other node-level artifacts of synthetic data.
-Once we consider soft interventions, however, the original Cdn formulation significantly outperforms the Mlp variant – indicating that graph-level information is essential.
-Dci is the best baseline, which is encouraging, as it was designed for the edge-centric version of our task, with the intuition that detecting edge differences is easier than reproducing the entire causal graph.
-It performs particularly well on linear data, which align with its assumptions, and less well on non-linear data.
+Dcdi’s -G and -Dsf suffixes correspond to Gaussian and deep sigmoidal flow parametrizations of the likelihood.
+Bacadi’s -E and -M suffixes indicate empirical (standard) and mixture (bootstrap) variants.
+Mb+Ci is inspired by Huang et al.
+(2020).
+We introduce a domain index (indicator of environment), compute its Markov boundary using graph lasso, and use conditional mutual information to identify true children of the domain index (intervention targets).
+Dci with stability selection (Belyaeva et al., 2021) predicts edge-level differences between two causal graphs.
+We take each node’s proportion of changed edges as its likelihood of being an intervention target.
+While Dci was also motivated by biological applications, it only scales to around a hundred variables at most, so we evaluate it alongside other causal discovery methods here.
+Evaluation  In the synthetic case, we are not constrained by biological redundancy or incomplete predictions, so we report standard classification metrics: mean average precision (mAP) and area under the ROC curve (AUC).
+Both metrics are computed independently for each variable and averaged over all regimes of the same number of targets.
+Their values range from 0 to 1 (perfect).
+The mAP random baseline depends on the positive rate, while the AUC random baseline is 0.5 (per edge).
+Results  On synthetic data, Cdn achieves high performance across intervention types and data-generating mechanisms (Table 3), while running in seconds (Table 11).
+As an ablation study, we investigated removing the graph loss and the pretrained weights, with no change to the graph-centric architecture.
+We found that supervising based on the graph is almost always helpful (“no $\mathcal{L}_{G}$” row), and pretraining leads to more stable training dynamics (Figure 10).
+Dci and Mb+Ci are the best baselines, which is encouraging, as they are designed with the intuition that detecting differences is easier than reproducing the entire causal graph.
 Surprisingly, most other baselines perform poorly at recovering intervention targets.
-In the case of Dcdi and BaCaDI, this may be because it is hard to select a single sparsity threshold for varying sizes of intervention sets, and to balance the regularizer with the generative modeling objective.
-5 Conclusion
-We presented Cdn, a causal discovery-powered approach for predicting perturbation targets in the context of single-cell transcriptomics.
-Cdn uses an amortized causal discovery algorithm to represent a pair of observational and interventional datasets in terms of their data-generating mechanisms, and then trains a classifier to identify variables whose conditional independencies have changed.
-Our approach achieves the state-of-the-art in predicting perturbation targets over seven transcriptomics datasets, compared to a variety of perturbation modeling baselines.
-Cdn also surpasses causal discovery algorithms for predicting unknown targets across synthetic settings, while only requiring a fraction of the runtime.
-Nonetheless, there is substantial space for improvement, so we hope that this work and the associated datasets will inform future method development for these biological applications.
-Reproducibility statement
-All source code will be made publicly available.
-Hyperparameter and implementation choices are described in Section 3.3.
-Details regarding biological data pre-processing can be found in Section 4.1 and Appendix B.1.
-The raw data are freely available via the accession codes listed in Table 5.
-Synthetic data generation is discussed in Appendix B.2.
-Details required to reproduce baselines are described in Sections 4.1 and 4.2, as well as Appendix C.1.
-Ethics statement
-We use publicly available datasets which do not involve human subjects.
-While our work has applications to biological understanding, our work does not directly concern the design of molecules or other potentially harmful agents.
+In the case of Dcdi and BaCaDI, this may be because it is hard to select a single sparsity threshold for varying sizes of intervention sets, and to balance the sparsity regularizer with the generative modeling objective.
+Impact statement
+Our work proposes a machine learning method to improve understanding of large-scale perturbation experiments, in the context of molecular biology.
+Towards broader impact, this work may facilitate better understanding of drug mechanisms; inform protocol design for cellular reprogramming; or identify causal targets of disease.
+While this work has numerous applications to the life sciences, it does not directly concern the design of molecules or other potentially harmful agents.
 Acknowledgements
@@
 We would like to acknowledge support from the NSF Expeditions grant (award 1918839: Collaborative Research: Understanding the World Through Code), Machine Learning for Pharmaceutical Discovery and Synthesis (MLPDS) consortium, and the Abdul Latif Jameel Clinic for Machine Learning in Health.
-Appendix A Theoretical intuition
-While this paper focuses on biological applications, we would like to show that our model is well-specified, i.e. it has the capacity to predict correct intervention targets, given certain inputs.
+Appendix A Model architecture
+A.1 Attention layers
+To balance expressivity and efficiency, our model is designed around series of “axial” attention layers, similar to Ho et al.
+(2020); Rao et al.
+(2021); Wu et al.
+(2024).
+The causal structure learner closely resembles the architecture of Sea (Wu et al., 2024), while the differential network is slightly different (since the input space is different).
+Briefly, the input to the causal structure learner is a matrix of shape $N\times N\times T\times d$.
+However, the (dense) global statistic does not span the $T$ dimension, and the local graph estimates are generally sparse.
+Thus, the causal graph learner separately attends over the $N\times N\times d$ and $K\times T\times d$ components ($K$ is the number of unique edges in sparse format), with “messages” to align the representations between each attention layer.
+In contrast, the differential network only operates over the dense $N\times(N+1)\times d$ (or $2d$) representation.
+Thus, it omits the message passing layers and directly stacks dense attention layers.
+A.2 Edge embeddings
+Since attention itself is invariant to input ordering, Transformer-style architectures use absolute (Devlin et al., 2019) or relative (Su et al., 2024) positional embeddings to distinguish input elements.
+Images and sequences have a natural ordering among their elements.
+This structure inherently induces distances between positions, which provides signal for learning models.
+In contrast, models over graphs should be equivariant to the labeling of nodes.
+That is, if $1\to 2$ is an edge, while $2$ and $3$ are not connected, then re-labeling the nodes $(1,2,3)\mapsto(2,3,1)$ should result in edge $2\to 3$ and no edge between $3$ and $1$.
+If we assign an arbitrary ordering to each graphx, the Transformer will be able to distinguish between nodes.
+However, there are several practical issues.
+Since synthetic data are generated in the topological ordering of the nodes, always assigning the root nodes to low positions may lead to unintended information leakage.
+Furthermore, if the model is trained on graphs of up to $N=100$, but tested on larger graphs, the positional embeddings associated with $N>100$ will be entirely random and out of distribution.
+Following Wu et al.
+(2024), we randomly sample a permutation of the maximum graph size to each graph.
+Each edge embedding is the concatenation of its constituent nodes’ embeddings, combined via a feed-forward network:
+$\text{Embed}(i,j)=\text{FFN}([\text{Embed}(i),\text{Embed}(j)])$ (9)
+where $i,j$ are arbitrarily assigned, but consistent within the graph. This embedding is added to the hidden representation $h_{i,j}$ before the first layer of attention.
+In the differential network, the additional “column” associated with node-level statistics is not subject to any embedding (which implicitly differentiates it from the remaining inputs), while the “row” identifies the node these statistics were computed for.
+Appendix B Implementation details
+B.1 Cdn training
+We trained until convergence (no improvement for 50 epochs on synthetic; 10 epochs on Perturb-seq) and selected the best model based on validation mAP (Figure 4).
+The validation distribution was identical to training and did not contain test causal mechanisms.
+During training, we used 15 CPU workers (primarily for local graph estimates) and 1 A6000 GPU.
+Cdn took around 6 hours to train on synthetic data, and 1 hour to finetune on real data.
+Unless otherwise noted, the causal structure learner was initialized from the Sea inverse covariance, FCI pretrained weights (Wu et al., 2024).
+Figure 4: Hyperparameter search: number of differential network layers. We selected 3 layers based on validation metrics.
+B.2 Baselines
+We used the latest releases of all baselines.
+Gears is a graph neural network that predicts the effects of unseen genetic perturbations, based on the Gene Ontology knowledge graph (Ashburner et al., 2000).
+We trained Gears on each cell line separately, and predicted the effects of perturbing every expressed gene.
+Then we ranked candidates based on cosine similarity to the interventional data.
+Figure 9 depicts the low, but non-trivial proportion of differentially expressed genes for which Gears was unable to make a prediction, due to lack of node coverage in their processed gene ontology graph (Ashburner et al., 2000).
+These genes were not considered in the rankings or evaluations for Gears.
+GenePt utilized the v2 March 2024 update, which used newer models and additional protein data, compared to their initial paper (GenePT_gene_protein_embedding_model_3_text).
+We concatenated each perturbation’s gene embedding to the log-fold change of the top 5000 highly-variable genes (Wolf et al., 2018) and trained a logistic regression model on each cell line to predict true vs.
+decoy perturbations.
+Candidates were ranked based on predicted probability.
+Of the 2,842 unique genetic perturbations, only 2,819 ($99.1\%$) mapped to GenePt embeddings.
+The remainder used the mean gene embedding (within the dataset) as the language-based embedding.
+PdGrapher was published on the union of three distinct knowledge graphs, but their harmonized graphs were not available publicly, and the authors did not respond to requests for data sharing.
+As a result, we relied on solely the human reference interactome (Luck et al., 2020) graph, as it was the only one of the three that could be easily processed.
+All test perturbation targets could be inferred through PdGrapher.
+Ut-Igsp used the official implementation from commit 7349396, since the repository is not back-compatible.
+Bacadi is evaluated using the fully-connected implementation, since it performed better than the linear version.
+However, since the linear version is significantly faster, we include its runtime in Table 11.
+Mb+Ci was implemented using scikit-learn (Pedregosa et al., 2011).
+Appendix C Theoretical motivation
+While this paper focuses on biological applications, we show that the differential network is well-specified as a class of models, i.e.
+a correct solution is contained within the space of possible parametrizations.
 Preliminary note
-Continuous causal discovery works typically rely on the universality of their architectures for consistency (Brouillard et al., 2020), so this analysis also focuses on computational capacity.
-Theoretically, it is impossible to “guarantee” that high-dimensional representations from learned model contain or do not contain certain information.
-This can be tested empirically through probing, e.g.
-querying for grammatical structure in language models (Vulić et al., 2020), but cannot be “proven” in the traditional sense.
-Thus, we believe that the empirical performance of our model on real datasets is its primary contribution, but provide these explanations for interested readers.
 In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$.
@@
 We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
-Axial attention architecture
-Our differential network is implemented using an axial attention layer, which is composed of two self-attention layers (one along each axis of the adjacency matrix) and a feed-forward network.
-For simplicity, we follow prior work Yun et al.
-(2019) and ignore layer normalization and dropout.
+Attention-based architecture
+Our differential network is implemented using an attention layer, which is composed of two self-attention layers (one along each axis of the adjacency matrix) and a feed-forward network.
+For simplicity, we follow prior work (Yun et al., 2019) and ignore layer normalization and dropout.
 Our inputs $h\in\mathbb{R}^{2d\times N\times N}$ are $2d$-dimension features, which represent a pair of $N\times N$ causal graphs.
 We use $h_{\cdot,j}$ to denote a length $N$ row for a fixed column $j$, and $h_{i,\cdot}$ to denote a length $N$ column for a fixed row $i$.
-The axial attention layer implements:
+The attention layer implements:
 $\text{Attn}_{\text{row}}(h_{\cdot,j})$ $=h_{\cdot,j}+W_{O}W_{V}h_{\cdot,j}\cdot\sigma\left[(W_{K}h_{\cdot,j})^{T}W_{Q}h_{\cdot,j}\right],$
@@
 In terms of the associated adjacency matrices, $E-E^{\prime}$ has 1s in each column $i\in I$ and 0s elsewhere.
-Here, the axial attention layer should implement $h-h^{\prime}$, so that when we collapse over the incoming edges, the output is non-zero only at the edge differences.
+Here, the attention layer should implement $h-h^{\prime}$, so that when we collapse over the incoming edges, the output is non-zero only at the edge differences.
 Suppose the first dimension of the $2d$ feature stores $E$, and the second dimension stores $E^{\prime}$.
@@
 In particular, we focus on two statistics: the correlation matrix $R$ and the covariance matrix $\Sigma$.
-Note that while these inputs differ slightly from our main experiments (reasons discussed in Appendix D), we show that they still achieve reasonable performance in Table 8.
+Note that while we do not explicitly provide the covariance matrix as input, we do provide marginal variances, from which the same information can be derived.
 Suppose $x$ is an intervention target.
@@
 Since the two techniques produce mutually exclusive predictions, this means that both hard and soft interventions can co-exist and be detected on different nodes.
-Finally, we re-iterate that Cdn operates over high-dimensional features, rather than the inputs and outputs themselves.
-Thus, these sketches only serve to provide context regarding the axial attention framework’s computational capacity, rather than as a blueprint for the computations it performs.
-Appendix B Datasets
-B.1 Biological datasets
-We converted all single cell datasets to LogTP10K + 1 expression values (log-normalized, transcripts per 10,000 UMIs).
-Perturb-seq dataset variables represented genes that were mapped and filtered by the authors.
+Appendix D Datasets
+D.1 Biological datasets
+We converted all single cell datasets to log-normalized, transcripts per 10,000 UMIs.
+Perturb-seq genes were mapped to standard identifiers and filtered by the authors.
 Sci-Plex dataset variables represented genes that appeared in at least 5,000 cells (threshold chosen to achieve a similar number of genes).
@@
 Table 5 reports statistics of the final, processed datasets.
-Figures 6 and 7 plot the full distribution of number of cells and DEGs per perturbation.
-The full names of the Sci-Plex drugs are as follows.
-  • Infigratinib (“infig”)
-  • Nintedanib (“nint”)
-  • Palbociclib (“palb”)
-  • Doxorubicin (“doxo”)
-  • Volasertib (“vola”)
+Figure 8 plots the full distribution of number of cells and DEGs per perturbation.
+Figures 6 and 7 depict the differences in gene co-expression matrices between control and perturbed cells.
+Genetic perturbations tend to have much clearer phenotypes compared to chemical perturbations, whose effects are more diffuse.
+The full names of the Sci-Plex drugs and their targets are as follows. These targets were collated by the authors (McFaline-Figueroa et al., 2024), and are each are well-documented in the literature.
+  • Infigratinib (“infig”) binds target FGFR1 with nano-molar affinities (Knox et al., 2024).
+  • Nintedanib (“nint”) and target FGFR1 have been probed through structural studies (Hilberg et al., 2008).
+  • Palbociclib (“palb”) has been co-crystallized (PDB 5L2I) with its target CDK6.
+  • Doxorubicin (“doxo”) and target TOP2A are associated in a number of works (Knox et al., 2024).
+  • Volasertib (“vola”) has been co-crystallized (PDB 3FC2) with PLK1 and a number of works confirm specificity for PLK2 (Chen et al., 2021).
+It is important to note that these drugs are also associated with known off-targets.
+However, we did not observe differential expression of any off-target, adjusted $p\approx 1$, perhaps because the impact on their mRNA expression is negligible.
+Protein activity would be more suitable for quantifying the impact of a drug on a cell.
+However, this cannot be measured at scale.
+In fact, even protein abundance is difficult to measure at single-cell resolution, as it relies on mass spectrometry to differentiate proteins (Newman et al., 2006), while mRNA can be exhaustively enumerated and sequenced (based on the known human genome).
+Therefore, this chemical perturbation study is primarily a proof of concept, to probe whether targets can be inferred from mRNA expression, and if so, to what extent.
 Table 4: Extended biological dataset statistics (raw).
@@
 T98G 3 1,644 508 13,126
-Figure 5: Heatmap of correlation between log-fold change, sorted by cluster.
-Figure 6: Perturb-seq dataset statistics, after processing (Table 6).
-Figure 7: Sci-Plex dataset statistics, after processing. There are 3 chemical perturbations per cell line.
+Figure 5: Heatmap of correlation between log-fold change, sorted by cluster, used for data splits.
+Figure 6: Difference between observational and interventional correlation matrices on K562 genome-wide. In some cases, the changes are minimal and specific; in others, they are more diffuse.
+Figure 7: Difference between observational and interventional correlation matrices on Sci-Plex datasets. Effects are much more diffuse than in genetic perturbations (Figure 6).
+Figure 8: Transcriptomics dataset statistics, after processing. See Table 5 for details.
 Figure 9: Percentage of differentially expression genes (per perturbation) that were not in the GEARS knowledge graph, and whose effects could not be predicted.
-B.2 Synthetic data
+D.2 Synthetic data
 Synthetic data were generated using code modified from Dcdi (Brouillard et al., 2020) for soft interventions.
@@
   • Linear: $f(x)=MXW+E$.
+  • Neural network, non-additive $f(x)=\text{Tanh}((X,E)W_{\text{in}})W_{\text{out}}$
+  • Neural network, additive: $f(x)=\text{Tanh}(XW_{\text{in}})W_{\text{out}}+E$
   • Polynomial: $f(x)=W_{0}+MXW_{1}+MX^{2}W_{2}+E$
-Root causal mechanisms are uniform. For hard interventions, we set $f(x)\leftarrow z$, where $z\sim\text{Uniform}(-1,1)$. For soft interventions, we set
+  • Sigmoid: $f(x)=\sum_{i=1}^{d}W_{i}\cdot\text{sigmoid}(X_{i})+E$
+Linear and the neural network variants were used for training. Polynomial and sigmoid were only used for testing.
+Root causal mechanisms are uniform. For hard interventions, we set $f(x)\leftarrow z$, where $z\sim\text{Uniform}(-1,1)$. For soft “scale” interventions, we set
 $f(x)$ $\leftarrow z_{1}^{\text{Sign}(z)}\cdot f(x)$
@@
 That is, we multiply $f(x)$ by a scaling factor that is equal probability $\lessgtr 1$ (constant across all observations).
-Appendix C Implementation details
-C.1 Baselines
-We used the latest releases of all baselines.
-For GenePt, this corresponds to the v2 March 2024 update, which used newer models and additional protein data, compared to their initial paper (GenePT_gene_protein_embedding_model_3_text).
-Due to the varying coverage of genes in their domain knowledge, several baselines failed to make predictions for certain genes.
-Of the 2,842 unique genetic perturbations, only 2,819 ($99.1\%$) mapped to GenePt embeddings.
-The remainder used the mean gene embedding (within the dataset) as the language-based embedding.
-Figure 8 depicts the low, but non-trivial proportion of differentially expressed genes for which Gears was unable to make a prediction, due to lack of node coverage in their processed gene ontology graph (Ashburner et al., 2000).
-These genes were not considered in the rankings or evaluations for Gears.
-PdGrapher was published on the union of three distinct knowledge graphs, but their harmonized graphs were not available publicly, and the authors did not respond to requests for data sharing.
-As a result, we relied on solely the human reference interactome (Luck et al., 2020) graph, as it was the only one of the three that could be easily processed.
-All test perturbation targets could be inferred through PdGrapher.
-Appendix D Additional analyses
-Table 7: Ablation studies. We re-train Cdn without pretrained aggregator weights, marginal estimates, or global statistics.
-K562 gw K562 es RPE1 HepG2 Jurkat
-Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
-Cdn 0.72 0.37 0.32 0.56 0.15 0.35 0.59 0.12 0.53 0.55 0.07 0.40 0.62 0.10 0.39
---pretrain $0.50$ $0.14$ $0.20$ $0.51$ $0.01$ $0.23$ $0.49$ $0.03$ $0.50$ $0.52$ $0.03$ $0.37$ $0.54$ $0.07$ $0.30$
---marginal $0.55$ $0.15$ $0.19$ $0.49$ $0.06$ $0.25$ $0.50$ $0.05$ $0.47$ $0.49$ $0.05$ $0.39$ $0.58$ $0.04$ $0.31$
---global $0.52$ $0.14$ $0.20$ $0.49$ $0.04$ $0.27$ $0.47$ $0.04$ $0.49$ $0.49$ $0.04$ $0.35$ $0.46$ $0.06$ $0.25$
-Table 8: Ablation study on global statistics over synthetic datasets.
-Mlp and Axial use inverse covariance (same as pretraining).
-Corr was finetuned for correlation, and Corr+Cov was finetuned for correlation and covariance (concatenated).
-$N$ $E$ Cdn Variant Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft)
-[…]
-[…]
-Axial $\textbf{0.73}\scriptstyle\pm.05$ $\textbf{0.85}\scriptstyle\pm.03$ $\textbf{0.48}\scriptstyle\pm.06$ $\textbf{0.64}\scriptstyle\pm.05$ $\textbf{0.62}\scriptstyle\pm.09$ 0.77±.07plus-or-minus0.77.070.$7$07 $\textbf{0.73}\scriptstyle\pm.07$ $\textbf{0.82}\scriptstyle\pm.06$
-+Corr $0.71\scriptstyle\pm.07$ $0.84\scriptstyle\pm.04$ $0.32\scriptstyle\pm.05$ $0.47\scriptstyle\pm.07$ $0.58\scriptstyle\pm.08$ $0.74\scriptstyle\pm.05$ $0.27\scriptstyle\pm.04$ $0.34\scriptstyle\pm.04$
-+Corr+Cov $0.68\scriptstyle\pm.04$ $0.82\scriptstyle\pm.02$ $0.39\scriptstyle\pm.03$ $0.50\scriptstyle\pm.03$ $0.58\scriptstyle\pm.07$ $0.71\scriptstyle\pm.07$ $0.48\scriptstyle\pm.05$ $0.65\scriptstyle\pm.05$
-[…]
-[…]
-+Corr $\textbf{0.78}\scriptstyle\pm.03$ 0.88±.02plus-or-minus0.88.020.$8$02 $0.35\scriptstyle\pm.02$ $0.52\scriptstyle\pm.02$ $0.73\scriptstyle\pm.06$ $0.84\scriptstyle\pm.03$ $0.34\scriptstyle\pm.05$ $0.41\scriptstyle\pm.06$
-+Corr+Cov $0.75\scriptstyle\pm.03$ $0.84\scriptstyle\pm.02$ $0.34\scriptstyle\pm.03$ $0.47\scriptstyle\pm.06$ $0.69\scriptstyle\pm.07$ $0.79\scriptstyle\pm.05$ $0.53\scriptstyle\pm.04$ $\textbf{0.68}\scriptstyle\pm.04$
-20 20 Mlp $\textbf{0.67}\scriptstyle\pm.05$ $\textbf{0.87}\scriptstyle\pm.02$ $0.28\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $60$06 $\textbf{0.84}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.04$ $0.51\scriptstyle\pm.04$
-Axial $0.65\scriptstyle\pm.03$ $0.86\scriptstyle\pm.01$ $\textbf{0.49}\scriptstyle\pm.05$ $\textbf{0.72}\scriptstyle\pm.03$ $0.58\scriptstyle\pm.05$ $0.83\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.81}\scriptstyle\pm.02$
-+Corr $0.62\scriptstyle\pm.03$ $0.83\scriptstyle\pm.01$ $0.25\scriptstyle\pm.05$ $0.47\scriptstyle\pm.05$ $0.58\scriptstyle\pm.05$ $0.81\scriptstyle\pm.02$ $0.21\scriptstyle\pm.02$ $0.37\scriptstyle\pm.03$
-+Corr+Cov $0.60\scriptstyle\pm.03$ $0.81\scriptstyle\pm.01$ $0.31\scriptstyle\pm.03$ $0.58\scriptstyle\pm.01$ $0.58\scriptstyle\pm.04$ $0.80\scriptstyle\pm.02$ $0.40\scriptstyle\pm.03$ $0.65\scriptstyle\pm.03$
-[…]
-Axial $0.74\scriptstyle\pm.05$ $0.90\scriptstyle\pm.02$ $\textbf{0.38}\scriptstyle\pm.03$ $\textbf{0.65}\scriptstyle\pm.01$ $0.68\scriptstyle\pm.08$ $0.87\scriptstyle\pm.04$ $\textbf{0.64}\scriptstyle\pm.03$ $\textbf{0.79}\scriptstyle\pm.03$
-[…]
-+Corr+Cov $0.73\scriptstyle\pm.05$ $0.87\scriptstyle\pm.02$ $0.36\scriptstyle\pm.03$ $0.58\scriptstyle\pm.03$ $0.62\scriptstyle\pm.07$ $0.82\scriptstyle\pm.03$ $0.43\scriptstyle\pm.04$ $0.67\scriptstyle\pm.02$
-Ablation studies
-We run ablation studies on the model architecture (Table 7) by re-training Cdn without pretrained aggregator weights, global statistics, or marginal estimates.
-We find that pretrained weights and global correlation are essential to model performance.
-The marginal estimates are of mixed importance: they are beneficial on the genome-wide (gw) dataset and Jurkat, but less on the others.
-Table 8 ablates variants of Cdn trained with different global statistics (but initialized with the inverse covariance pretrained weights) on synthetic data.
-The default summary statistic (Mlp, Axial) is inverse covariance.
-Corr denotes the correlation version, later finetuned on Perturb-seq data (since correlation is much easier to compute on large graphs, compared to inverse covariance).
-The Corr+Cov version aligns exactly with our theoretical intuition, and performs well in most settings (polynomial soft is the exception).
-This model modified the global layer to take an input of size $N\times N\times 2$, where each embedding was initialized to a distinct copy of the original.
-However, the inverse covariance version still performed the best, perhaps since there is no mismatch between the pretraining and finetuning statistics.
+For soft “shift” interventions, we set
+$f(x)$ $\leftarrow f(x)+\text{Sign}(z)\cdot z_{1}$
+$z_{1}$ $\sim\text{Uniform}(2,4)$
+$z$ $\sim\text{Uniform}(-1,1).$
+That is, we translate $f(x)$ by a scaling factor that is equal probability $\lessgtr 0$ (constant across all observations).
+Appendix E Additional analyses
+Figure 10: Ablation study: We compare Cdn trained on synthetic data, to the same model without any ground truth graph information (no pretraining or $\mathcal{L}_{G}$).
+We find that graph information leads to more stable and better results.
+Validation metrics and loss.
+One step is one batch (16 samples).
 Runtime
-We also compared the runtimes of various algorithms on the Perturb-seq and synthetic datasets.
+We compared the runtimes of various algorithms on the Perturb-seq and synthetic datasets.
 On the Perturb-seq data, all models finished running within minutes with the exception of Gears (Figure 11).
@@
 Due to space limitations, we report uncertainty estimates in Table 7.
-Multiple baselines produce deterministic results (Linear, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells.
+Multiple baselines produce deterministic results (Linear, Dge, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells.
 Specifically, for each perturbation with $M$ cells, we sample
@@
 cells uniformly at random, repeated 5 times.
+Perturb-seq trivial perturbations
+For the main results in Table 1, we focus on genetic perturbations for which the intended gene target was not the gene with the greatest log-fold change.
+This is because genetic perturbations are (unrealistically) specific, compared to drugs or transcription factors.
+In Table 8, we show that our model performs nearly perfectly in this easy setting.
+Graph loss
+We assess the value of predicting the causal graph, as opposed to only training a supervised classifier (Figure 10).
+Without modifying the architecture (which has inductive biases towards the graph modality), we train a version of Cdn without any graph labels (causal graph learner no longer pretrained; remove graph loss $\mathcal{L}_{G}$).
+Across the validation loss and various metrics, we find that the full Cdn is more stable, and leads to better results.
+This model corresponds to the ”no $\mathcal{L}_{G}$” test results in Table 3.
 Additional synthetic results
-Table 11 reports results on all synthetic test datasets, averaging over all interventions for each dataset.
+During model development, we realized that if the model is allowed to train on all types of perturbations, it can easily overfit and learn these distributional shifts very well.
+For example, in Table 9, the models trained on both types of soft interventions performs perfectly (unrealistic) in many cases, even though the graphs and causal mechanism parameters are completely novel.
+Therefore, to assess whether our model can generalize, we chose to hold out a class of soft intervention for our main results in Table 3.
+Table 10 reports results on all synthetic test datasets, averaging over all interventions for each dataset. We also include the “full” version of Cdn, i.e. trained on all mechanisms, in this analysis.
 Figure 11: Inference runtimes on Perturb-seq datasets.
@@
 Cdn 36 86 54 13
-Table 10: Uncertainty quantification on Perturb-seq datasets by sub-sampling to $80\%$ of cells per perturbation (or 50, whichever is higher).
-Standard deviation reported over 5 rounds of sub-sampling.
-Cdn is also the most robust to sub-sampling, compared to baselines.
+Table 7: Uncertainty quantification on Perturb-seq datasets, seen cell lines, by sub-sampling to $80\%$ of cells per perturbation (or 50, whichever is higher).
+Note that results may be slightly different from Table 1 due to sub-sampling.
+Standard deviation reported over 5 seeds.
 GenePt performance is highly variable.
+Dge is quite sensitive to sub-sampling on K562 genome-wide.
 K562 gw K562 es RPE1
-Model rank $R_{20}$ $r$ rank $R_{20}$ $r$ rank $R_{20}$ $r$
-Linear $0.50\scriptstyle\pm.00$ $0.17\scriptstyle\pm.01$ $0.20\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $10$01 $0.28\scriptstyle\pm.01$ $0.48\scriptstyle\pm.00$ $0.07\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$
-Mlp $0.48\scriptstyle\pm.00$ $0.21\scriptstyle\pm.00$ $0.20\scriptstyle\pm.00$ $0.48\scriptstyle\pm.00$ 0.11±.00plus-or-minus0.11.000.$1$00 $0.19\scriptstyle\pm.00$ $0.53\scriptstyle\pm.00$ $0.15\scriptstyle\pm.00$ $0.39\scriptstyle\pm.00$
-GenePt $0.54\scriptstyle\pm.09$ $0.28\scriptstyle\pm.06$ $0.29\scriptstyle\pm.08$ $0.50\scriptstyle\pm.08$ $0.21\scriptstyle\pm.09$ $0.39\scriptstyle\pm.10$ $0.51\scriptstyle\pm.10$ $0.20\scriptstyle\pm.08$ $0.51\scriptstyle\pm.07$
-Gears $0.50\scriptstyle\pm.01$ $0.13\scriptstyle\pm.01$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.51\scriptstyle\pm.00$ $0.07\scriptstyle\pm.00$ $0.23\scriptstyle\pm.00$ $0.46\scriptstyle\pm.00$ $0.07\scriptstyle\pm.01$ $0.41\scriptstyle\pm.02$
-PdG $0.49\scriptstyle\pm.01$ $0.21\scriptstyle\pm.01$ $0.25\scriptstyle\pm.00$ $0.50\scriptstyle\pm.01$ $0.16\scriptstyle\pm.01$ $0.32\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.08\scriptstyle\pm.01$ $0.47\scriptstyle\pm.01$
+Model rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$
+Linear $0.50\scriptstyle\pm.00$ $0.17\scriptstyle\pm.01$ $0.20\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $0.25\scriptstyle\pm.01$ $0.28\scriptstyle\pm.01$ $0.48\scriptstyle\pm.00$ $0.40\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$
+Mlp $0.48\scriptstyle\pm.00$ $0.17\scriptstyle\pm.00$ $0.20\scriptstyle\pm.00$ $0.48\scriptstyle\pm.00$ $0.16\scriptstyle\pm.00$ $0.19\scriptstyle\pm.00$ $0.53\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$ $0.39\scriptstyle\pm.00$
+Dge $0.72\scriptstyle\pm.01$ $0.48\scriptstyle\pm.02$ $0.52\scriptstyle\pm.02$ $0.86\scriptstyle\pm.00$ $0.59\scriptstyle\pm.02$ $0.61\scriptstyle\pm.02$ $0.74\scriptstyle\pm.00$ $0.56\scriptstyle\pm.00$ $0.64\scriptstyle\pm.00$
+GenePt $0.54\scriptstyle\pm.09$ $0.26\scriptstyle\pm.08$ $0.29\scriptstyle\pm.08$ $0.50\scriptstyle\pm.08$ $0.36\scriptstyle\pm.10$ $0.39\scriptstyle\pm.10$ $0.51\scriptstyle\pm.10$ $0.43\scriptstyle\pm.08$ $0.51\scriptstyle\pm.07$
+Gears $0.50\scriptstyle\pm.01$ $0.19\scriptstyle\pm.01$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.51\scriptstyle\pm.00$ $0.20\scriptstyle\pm.01$ $0.23\scriptstyle\pm.00$ $0.46\scriptstyle\pm.00$ 0.33±.01plus-or-minus0.33.010.$3$01 $0.41\scriptstyle\pm.02$
+PdG $0.49\scriptstyle\pm.01$ 0.22±.00plus-or-minus0.22.000.$2$00 $0.25\scriptstyle\pm.00$ $0.50\scriptstyle\pm.01$ $0.28\scriptstyle\pm.01$ $0.32\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.38\scriptstyle\pm.01$ $0.47\scriptstyle\pm.01$
+Cdn $\textbf{0.91}\scriptstyle\pm.01$ $\textbf{0.71}\scriptstyle\pm.02$ $\textbf{0.72}\scriptstyle\pm.02$ $\textbf{0.95}\scriptstyle\pm.00$ $\textbf{0.80}\scriptstyle\pm.03$ $\textbf{0.81}\scriptstyle\pm.03$ $\textbf{0.92}\scriptstyle\pm.00$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.74}\scriptstyle\pm.01$
+HepG2 Jurkat
+rank $\rho$ $r$ rank $\rho$ $r$
+Linear $0.49\scriptstyle\pm.00$ $0.34\scriptstyle\pm.01$ $0.42\scriptstyle\pm.01$ $0.50\scriptstyle\pm.00$ $0.23\scriptstyle\pm.01$ $0.27\scriptstyle\pm.01$
+Mlp $0.50\scriptstyle\pm.00$ $0.34\scriptstyle\pm.00$ $0.41\scriptstyle\pm.00$ $0.47\scriptstyle\pm.00$ $0.25\scriptstyle\pm.00$ $0.31\scriptstyle\pm.00$
+Dge $0.83\scriptstyle\pm.00$ $0.49\scriptstyle\pm.01$ $0.50\scriptstyle\pm.01$ $0.78\scriptstyle\pm.01$ 0.44±.02plus-or-minus0.44.020.$4$02 $0.49\scriptstyle\pm.02$
+GenePt $0.50\scriptstyle\pm.13$ $0.39\scriptstyle\pm.07$ $0.45\scriptstyle\pm.06$ $0.51\scriptstyle\pm.07$ $0.36\scriptstyle\pm.11$ $0.40\scriptstyle\pm.11$
+Gears $0.51\scriptstyle\pm.00$ $0.38\scriptstyle\pm.00$ $0.45\scriptstyle\pm.00$ $0.52\scriptstyle\pm.01$ $0.24\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$
+PdG $0.49\scriptstyle\pm.00$ $0.34\scriptstyle\pm.01$ $0.41\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ $0.29\scriptstyle\pm.00$ $0.34\scriptstyle\pm.00$
+Cdn $\textbf{0.93}\scriptstyle\pm.00$ $\textbf{0.67}\scriptstyle\pm.01$ $\textbf{0.73}\scriptstyle\pm.01$ $\textbf{0.96}\scriptstyle\pm.01$ $\textbf{0.83}\scriptstyle\pm.04$ $\textbf{0.85}\scriptstyle\pm.03$
+Table 8: Results on 5 Perturb-seq datasets, trivial perturbations (intended target has largest log-fold change, though not necessarily smallest adjusted p-value).
+Top: Train on all cell lines jointly.
+Bottom: Leave one cell line out of training.
+Test sets (unseen perturbations) are identical in both settings.
+Cdn (synthetic) and Dge do not require training on real Perturb-seq data.
+Metrics, from left to right: normalized rank of ground truth, top 1 Spearman and Pearson correlations.
+K562 (gw) K562 (es) RPE1 HepG2 Jurkat
+Setting Model rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$ rank $\rho$ $r$
+Seen cell line Linear $0.49$ $0.18$ $0.21$ $0.50$ $0.28$ $0.31$ $0.48$ $0.37$ $0.45$ $0.48$ $0.34$ $0.43$ $0.48$ $0.24$ $0.29$
+Mlp $0.48$ $0.16$ $0.20$ $0.47$ $0.17$ $0.21$ $0.54$ $0.29$ $0.37$ $0.51$ $0.35$ $0.42$ $0.48$ $0.26$ $0.32$
+GenePt $0.60$ $0.36$ $0.39$ $0.47$ $0.23$ $0.26$ $0.57$ $0.49$ $0.55$ $0.44$ $0.35$ $0.41$ $0.55$ $0.48$ $0.53$
+Gears $0.49$ $0.15$ $0.18$ $0.51$ $0.25$ $0.27$ $0.49$ $0.32$ $0.40$ $0.48$ $0.32$ $0.40$ $0.47$ $0.20$ $0.24$
+PdG $0.49$ $0.23$ $0.26$ $0.52$ $0.31$ $0.34$ $0.51$ $0.34$ $0.42$ $0.46$ $0.34$ $0.40$ $0.49$ $0.30$ $0.35$
+Cdn (Perturb-seq) 0.99 0.97 0.97 0.99 0.98 0.98 0.99 0.93 0.94 0.99 0.95 0.96 0.99 0.96 0.96
+Unseen cell line Dge $0.87$ $0.68$ $0.70$ $0.92$ $0.77$ $0.79$ $0.85$ $0.73$ $0.77$ $0.89$ $0.59$ $0.60$ $0.91$ $0.73$ $0.77$
+PdG $0.40$ $0.12$ $0.15$ $0.41$ $0.22$ $0.24$ $0.48$ $0.28$ $0.37$ $0.47$ $0.33$ $0.41$ $0.53$ $0.28$ $0.33$
+Cdn (synthetic) $0.91$ $0.72$ $0.74$ $0.92$ $0.83$ $0.84$ $0.94$ $0.83$ $0.86$ $0.90$ $0.73$ $0.77$ $0.95$ $0.82$ $0.84$
+Cdn (Perturb-seq) 0.98 0.95 0.95 0.99 0.95 0.96 0.99 0.92 0.94 0.99 0.95 0.96 0.99 0.96 0.96
+Table 9: Intervention target prediction results on synthetic datasets with $N=20,E=40$.
+Top: hard interventions (uniform); bottom: soft interventions (scale).
+We found that if the model is allowed to see the class of soft intervention (“all mechanisms”) during training, it can easily learn (overfit) the distribution shift perfectly, even though test datasets and graphs are novel.
+Thus, our main results in Table 3 treat “scale” as an unseen soft intervention.
+Linear (1) Linear (3) Polynomial (1) Polynomial (3)
 […]
-HepG2 Jurkat
-rank $R_{20}$ $r$ rank $R_{20}$ $r$
-Linear $0.49\scriptstyle\pm.00$ $0.06\scriptstyle\pm.01$ $0.42\scriptstyle\pm.01$ $0.50\scriptstyle\pm.00$ $0.08\scriptstyle\pm.00$ $0.27\scriptstyle\pm.01$
-Mlp $0.50\scriptstyle\pm.00$ $0.09\scriptstyle\pm.00$ $0.41\scriptstyle\pm.00$ $0.47\scriptstyle\pm.00$ $0.12\scriptstyle\pm.00$ $0.31\scriptstyle\pm.00$
-GenePt $0.50\scriptstyle\pm.13$ $\textbf{0.19}\scriptstyle\pm.08$ $0.45\scriptstyle\pm.06$ $0.51\scriptstyle\pm.07$ 0.22±.09plus-or-minus0.22.090.$2$09 $0.40\scriptstyle\pm.11$
-Gears $0.51\scriptstyle\pm.00$ 0.11±.00plus-or-minus0.11.000.$1$00 $0.45\scriptstyle\pm.00$ $0.52\scriptstyle\pm.01$ $0.08\scriptstyle\pm.00$ $0.30\scriptstyle\pm.00$
-PdG $0.49\scriptstyle\pm.00$ $0.05\scriptstyle\pm.01$ $0.41\scriptstyle\pm.01$ $0.49\scriptstyle\pm.01$ 0.11±.01plus-or-minus0.11.010.$1$01 $0.34\scriptstyle\pm.00$
-Cdn $\textbf{0.65}\scriptstyle\pm.01$ $0.17\scriptstyle\pm.01$ $\textbf{0.46}\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.01$ $\textbf{0.24}\scriptstyle\pm.02$ $\textbf{0.43}\scriptstyle\pm.01$
+Hard all mechanisms ($h_{\text{cat}}$) $0.86\scriptstyle\pm.09$ $0.97\scriptstyle\pm.02$ $0.91\scriptstyle\pm.06$ $0.97\scriptstyle\pm.03$ $0.86\scriptstyle\pm.04$ $0.97\scriptstyle\pm.01$ $0.91\scriptstyle\pm.03$ $0.97\scriptstyle\pm.01$
+all mechanisms ($h_{\text{diff}}$) $0.78\scriptstyle\pm.05$ $0.95\scriptstyle\pm.03$ $0.89\scriptstyle\pm.04$ $0.96\scriptstyle\pm.02$ $0.83\scriptstyle\pm.05$ $0.96\scriptstyle\pm.01$ $0.89\scriptstyle\pm.03$ $0.96\scriptstyle\pm.02$
+exclude scale $0.82\scriptstyle\pm.05$ $0.96\scriptstyle\pm.02$ 0.88±.04plus-or-minus0.88.040.$8$04 $0.95\scriptstyle\pm.02$ $0.85\scriptstyle\pm.05$ $0.97\scriptstyle\pm.01$ $0.90\scriptstyle\pm.04$ $0.96\scriptstyle\pm.02$
+Soft all mechanisms ($h_{\text{cat}}$) $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ 0.99±.01plus-or-minus0.99.010.$9$01 $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
+all mechanisms ($h_{\text{diff}}$) 0.99±.01plus-or-minus0.99.010.$9$01 $1.0\scriptstyle\pm.00$ $0.98\scriptstyle\pm.01$ 0.99±.00plus-or-minus0.99.000.$9$00 $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
+exclude scale $0.83\scriptstyle\pm.10$ $0.95\scriptstyle\pm.05$ $0.76\scriptstyle\pm.07$ $0.89\scriptstyle\pm.04$ $0.97\scriptstyle\pm.03$ $1.0\scriptstyle\pm.00$ $0.96\scriptstyle\pm.02$ 0.99±.01plus-or-minus0.99.010.$9$01
 Table 10: Intervention target prediction results on synthetic datasets, extended results.
-Uncertainty is standard deviation over 5 distinct datasets.
+Uncertainty is standard deviation over 5 i.i.d.
+datasets.
 Metrics are averaged over all $3N$ perturbations for a given dataset (1-3 targets).
-$N$ $E$ Model Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft)
+Cdn (full) denotes that we trained on all intervention mechanisms jointly, i.e.
+as in Table 9.
+$N$ $E$ Model Linear (Hard) Linear (Soft) Poly. (Hard) Poly. (Soft) Sigmoid. (Hard) Sigmoid (Soft)
 […]
-10 10 Ut-Igsp $0.29\scriptstyle\pm.02$ $0.57\scriptstyle\pm.03$ $0.30\scriptstyle\pm.01$ $0.63\scriptstyle\pm.02$ 0.33±.04plus-or-minus0.33.040.$3$04 $0.59\scriptstyle\pm.03$ $0.32\scriptstyle\pm.01$ $0.64\scriptstyle\pm.02$
-Dcdi-G $0.39\scriptstyle\pm.04$ $0.52\scriptstyle\pm.04$ $40$04 $0.51\scriptstyle\pm.04$ $0.38\scriptstyle\pm.03$ $0.49\scriptstyle\pm.02$ $0.36\scriptstyle\pm.03$ $0.49\scriptstyle\pm.03$
-Dcdi-Dsf $0.36\scriptstyle\pm.02$ $0.51\scriptstyle\pm.04$ $0.40\scriptstyle\pm.05$ $0.53\scriptstyle\pm.05$ $0.37\scriptstyle\pm.02$ $0.50\scriptstyle\pm.01$ $0.38\scriptstyle\pm.02$ $0.49\scriptstyle\pm.03$
-Dci 0.55±.08plus-or-minus0.55.080.$5$08 $0.79\scriptstyle\pm.04$ $\textbf{0.50}\scriptstyle\pm.08$ $\textbf{0.70}\scriptstyle\pm.05$ $0.47\scriptstyle\pm.09$ $0.71\scriptstyle\pm.06$ $0.41\scriptstyle\pm.02$ $0.67\scriptstyle\pm.01$
-BaCaDI-E $0.26\scriptstyle\pm.03$ $0.61\scriptstyle\pm.05$ $0.42\scriptstyle\pm.13$ $0.63\scriptstyle\pm.11$ $0.25\scriptstyle\pm.03$ $0.61\scriptstyle\pm.05$ $0.30\scriptstyle\pm.05$ $0.61\scriptstyle\pm.04$
-BaCaDI-M $0.23\scriptstyle\pm.01$ $0.56\scriptstyle\pm.03$ $0.37\scriptstyle\pm.11$ $0.62\scriptstyle\pm.11$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.56\scriptstyle\pm.04$ $0.30\scriptstyle\pm.05$ $0.61\scriptstyle\pm.04$
-[…]
-Cdn (Axial) $\textbf{0.73}\scriptstyle\pm.05$ $\textbf{0.85}\scriptstyle\pm.03$ $0.48\scriptstyle\pm.06$ $0.64\scriptstyle\pm.05$ $\textbf{0.62}\scriptstyle\pm.09$ 0.77±.07plus-or-minus0.77.070.$7$07 $\textbf{0.73}\scriptstyle\pm.07$ $\textbf{0.82}\scriptstyle\pm.06$
-10 20 Ut-Igsp $0.28\scriptstyle\pm.02$ $0.56\scriptstyle\pm.03$ $0.26\scriptstyle\pm.01$ $0.59\scriptstyle\pm.03$ $0.28\scriptstyle\pm.03$ $0.58\scriptstyle\pm.04$ $0.27\scriptstyle\pm.01$ $0.60\scriptstyle\pm.01$
-Dcdi-G $0.42\scriptstyle\pm.02$ $0.54\scriptstyle\pm.02$ $0.40\scriptstyle\pm.03$ $0.51\scriptstyle\pm.04$ $0.40\scriptstyle\pm.06$ $0.53\scriptstyle\pm.05$ $0.39\scriptstyle\pm.02$ $0.51\scriptstyle\pm.03$
-Dcdi-Dsf $0.41\scriptstyle\pm.03$ $0.52\scriptstyle\pm.03$ $0.39\scriptstyle\pm.05$ $0.52\scriptstyle\pm.05$ $0.39\scriptstyle\pm.02$ $0.51\scriptstyle\pm.02$ $0.39\scriptstyle\pm.03$ $50$05
-[…]
-BaCaDI-E $0.34\scriptstyle\pm.04$ $0.68\scriptstyle\pm.03$ $0.53\scriptstyle\pm.08$ $0.71\scriptstyle\pm.04$ 0.33±.03plus-or-minus0.33.030.$3$03 $0.72\scriptstyle\pm.04$ $0.64\scriptstyle\pm.07$ $0.78\scriptstyle\pm.05$
-BaCaDI-M $0.27\scriptstyle\pm.02$ $0.63\scriptstyle\pm.03$ $0.48\scriptstyle\pm.09$ $0.71\scriptstyle\pm.04$ $0.27\scriptstyle\pm.04$ $0.64\scriptstyle\pm.08$ $0.59\scriptstyle\pm.09$ 0.77±.06plus-or-minus0.77.060.$7$06
-[…]
-[…]
-20 20 Ut-Igsp $0.15\scriptstyle\pm.01$ $0.54\scriptstyle\pm.01$ $0.18\scriptstyle\pm.01$ $0.65\scriptstyle\pm.02$ $20$02 $0.60\scriptstyle\pm.02$ $0.21\scriptstyle\pm.02$ $0.67\scriptstyle\pm.02$
-Dcdi-G $0.24\scriptstyle\pm.02$ $0.49\scriptstyle\pm.02$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.02$ 0.22±.03plus-or-minus0.22.030.$2$03 $0.50\scriptstyle\pm.03$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.03$
-Dcdi-Dsf $0.24\scriptstyle\pm.02$ $0.50\scriptstyle\pm.02$ $0.23\scriptstyle\pm.02$ $0.50\scriptstyle\pm.03$ 0.22±.01plus-or-minus0.22.010.$2$01 $0.50\scriptstyle\pm.02$ $0.20\scriptstyle\pm.01$ $0.48\scriptstyle\pm.02$
-Dci $0.43\scriptstyle\pm.03$ 0.77±.01plus-or-minus0.77.010.$7$01 $0.45\scriptstyle\pm.08$ $\textbf{0.75}\scriptstyle\pm.04$ $0.48\scriptstyle\pm.05$ $0.76\scriptstyle\pm.03$ $0.50\scriptstyle\pm.02$ $0.79\scriptstyle\pm.02$
-BaCaDI-E $0.14\scriptstyle\pm.01$ $0.61\scriptstyle\pm.03$ $0.25\scriptstyle\pm.10$ $60$06 0.22±.06plus-or-minus0.22.060.$2$06 $0.76\scriptstyle\pm.06$ $0.30\scriptstyle\pm.06$ $0.74\scriptstyle\pm.05$
-BaCaDI-M $0.12\scriptstyle\pm.00$ $0.58\scriptstyle\pm.02$ $0.17\scriptstyle\pm.03$ $0.60\scriptstyle\pm.05$ $0.14\scriptstyle\pm.02$ 0.66±.05plus-or-minus0.66.050.$6$05 0.22±.05plus-or-minus0.22.050.$2$05 $0.71\scriptstyle\pm.06$
-Cdn (Mlp) $\textbf{0.67}\scriptstyle\pm.05$ $\textbf{0.87}\scriptstyle\pm.02$ $0.28\scriptstyle\pm.01$ $0.54\scriptstyle\pm.03$ $60$06 $\textbf{0.84}\scriptstyle\pm.03$ $0.34\scriptstyle\pm.04$ $0.51\scriptstyle\pm.04$
-Cdn (Axial) $0.65\scriptstyle\pm.03$ $0.86\scriptstyle\pm.01$ $\textbf{0.49}\scriptstyle\pm.05$ $0.72\scriptstyle\pm.03$ $0.58\scriptstyle\pm.05$ $0.83\scriptstyle\pm.02$ $\textbf{0.65}\scriptstyle\pm.02$ $\textbf{0.81}\scriptstyle\pm.02$
-20 40 Ut-Igsp $0.14\scriptstyle\pm.01$ $0.57\scriptstyle\pm.01$ $0.14\scriptstyle\pm.01$ $0.61\scriptstyle\pm.02$ $0.20\scriptstyle\pm.01$ $0.62\scriptstyle\pm.02$ $0.19\scriptstyle\pm.01$ $0.65\scriptstyle\pm.01$
-Dcdi-G 0.22±.03plus-or-minus0.22.030.$2$03 $0.49\scriptstyle\pm.03$ $0.21\scriptstyle\pm.02$ $0.48\scriptstyle\pm.03$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.49\scriptstyle\pm.02$ 0.22±.02plus-or-minus0.22.020.$2$02 $0.50\scriptstyle\pm.02$
-Dcdi-Dsf $0.21\scriptstyle\pm.02$ $0.46\scriptstyle\pm.03$ $0.21\scriptstyle\pm.02$ $0.47\scriptstyle\pm.02$ $0.23\scriptstyle\pm.02$ $0.50\scriptstyle\pm.03$ $0.21\scriptstyle\pm.01$ $0.49\scriptstyle\pm.03$
-[…]
-BaCaDI-E $0.23\scriptstyle\pm.05$ $0.70\scriptstyle\pm.04$ 0.33±.10plus-or-minus0.33.100.$3$10 $0.67\scriptstyle\pm.04$ $0.29\scriptstyle\pm.08$ $0.83\scriptstyle\pm.04$ $0.45\scriptstyle\pm.18$ $0.78\scriptstyle\pm.06$
-BaCaDI-M $0.15\scriptstyle\pm.01$ $0.64\scriptstyle\pm.02$ $0.19\scriptstyle\pm.04$ $0.63\scriptstyle\pm.02$ $0.17\scriptstyle\pm.02$ $0.70\scriptstyle\pm.05$ $0.31\scriptstyle\pm.10$ 0.77±.07plus-or-minus0.77.070.$7$07
-[…]
-Cdn (Axial) $0.74\scriptstyle\pm.05$ $0.90\scriptstyle\pm.02$ $0.38\scriptstyle\pm.03$ $0.65\scriptstyle\pm.01$ $0.68\scriptstyle\pm.08$ $0.87\scriptstyle\pm.04$ $\textbf{0.64}\scriptstyle\pm.03$ $0.79\scriptstyle\pm.03$
+10 10 Ut-Igsp .$29\scriptstyle\pm.02$$57\scriptstyle\pm.03$$30\scriptstyle\pm.01$$63\scriptstyle\pm.02$$3$04 .$59\scriptstyle\pm.03$$32\scriptstyle\pm.01$$64\scriptstyle\pm.02$$28\scriptstyle\pm.03$$58\scriptstyle\pm.04$$26\scriptstyle\pm.02$$59\scriptstyle\pm.03$
+Dcdi-G .$39\scriptstyle\pm.04$$52\scriptstyle\pm.04$$40\scriptstyle\pm.04$$51\scriptstyle\pm.04$$38\scriptstyle\pm.03$$49\scriptstyle\pm.02$$36\scriptstyle\pm.03$$49\scriptstyle\pm.03$$37\scriptstyle\pm.03$$49\scriptstyle\pm.03$$40\scriptstyle\pm.04$$52\scriptstyle\pm.06$
+Dcdi-Dsf .$36\scriptstyle\pm.02$$51\scriptstyle\pm.04$$40\scriptstyle\pm.05$$53\scriptstyle\pm.05$$37\scriptstyle\pm.02$$50\scriptstyle\pm.01$$38\scriptstyle\pm.02$$49\scriptstyle\pm.03$$37\scriptstyle\pm.02$$50\scriptstyle\pm.02$$37\scriptstyle\pm.02$$48\scriptstyle\pm.03$
+Bacadi-E .$26\scriptstyle\pm.03$$61\scriptstyle\pm.05$$42\scriptstyle\pm.13$$63\scriptstyle\pm.11$$25\scriptstyle\pm.03$$61\scriptstyle\pm.05$$30\scriptstyle\pm.05$$61\scriptstyle\pm.04$$28\scriptstyle\pm.04$$64\scriptstyle\pm.06$$43\scriptstyle\pm.12$$69\scriptstyle\pm.09$
+Bacadi-M .$23\scriptstyle\pm.01$$56\scriptstyle\pm.03$$37\scriptstyle\pm.11$$62\scriptstyle\pm.11$$2$02 .$56\scriptstyle\pm.04$$30\scriptstyle\pm.05$$61\scriptstyle\pm.04$$25\scriptstyle\pm.03$$61\scriptstyle\pm.05$$42\scriptstyle\pm.11$$69\scriptstyle\pm.09$
+Dci .55±.08plus-or-minus.55.08.$5$08 .$79\scriptstyle\pm.04$$50\scriptstyle\pm.08$$70\scriptstyle\pm.05$$47\scriptstyle\pm.09$$71\scriptstyle\pm.06$$41\scriptstyle\pm.02$$67\scriptstyle\pm.01$$59\scriptstyle\pm.07$$80\scriptstyle\pm.05$$51\scriptstyle\pm.05$$76\scriptstyle\pm.04$
+Mb+Ci .$53\scriptstyle\pm.05$$64\scriptstyle\pm.03$$57\scriptstyle\pm.10$$6$06 .$70\scriptstyle\pm.09$$79\scriptstyle\pm.07$$85\scriptstyle\pm.07$$89\scriptstyle\pm.05$$84\scriptstyle\pm.08$$8$06 .$96\scriptstyle\pm.05$$97\scriptstyle\pm.04$
+Cdn .$86\scriptstyle\pm.02$$93\scriptstyle\pm.01$$74\scriptstyle\pm.06$$85\scriptstyle\pm.06$$83\scriptstyle\pm.09$$91\scriptstyle\pm.05$$89\scriptstyle\pm.05$$94\scriptstyle\pm.03$$8$06 .$94\scriptstyle\pm.04$$85\scriptstyle\pm.07$$91\scriptstyle\pm.05$
+Cdn (no $\mathcal{L}_{G}$) .$67\scriptstyle\pm.09$$78\scriptstyle\pm.07$$82\scriptstyle\pm.05$$89\scriptstyle\pm.05$$82\scriptstyle\pm.08$$91\scriptstyle\pm.05$$91\scriptstyle\pm.04$$96\scriptstyle\pm.02$$89\scriptstyle\pm.05$$95\scriptstyle\pm.03$$94\scriptstyle\pm.03$$97\scriptstyle\pm.02$
+Cdn (full) .$87\scriptstyle\pm.01$$94\scriptstyle\pm.01$$9$01 $1.0\scriptstyle\pm.00$$82\scriptstyle\pm.08$$91\scriptstyle\pm.05$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$95\scriptstyle\pm.05$$98\scriptstyle\pm.02$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
+10 20 Ut-Igsp .$28\scriptstyle\pm.02$$56\scriptstyle\pm.03$$26\scriptstyle\pm.01$$59\scriptstyle\pm.03$$28\scriptstyle\pm.03$$58\scriptstyle\pm.04$$27\scriptstyle\pm.01$$60\scriptstyle\pm.01$$25\scriptstyle\pm.00$$5$03 .$25\scriptstyle\pm.02$$56\scriptstyle\pm.05$
+Dcdi-G .$42\scriptstyle\pm.02$$54\scriptstyle\pm.02$$40\scriptstyle\pm.03$$51\scriptstyle\pm.04$$40\scriptstyle\pm.06$$53\scriptstyle\pm.05$$39\scriptstyle\pm.02$$51\scriptstyle\pm.03$$37\scriptstyle\pm.02$$50\scriptstyle\pm.01$$35\scriptstyle\pm.03$$49\scriptstyle\pm.04$
+Dcdi-Dsf .$41\scriptstyle\pm.03$$52\scriptstyle\pm.03$$39\scriptstyle\pm.05$$52\scriptstyle\pm.05$$39\scriptstyle\pm.02$$51\scriptstyle\pm.02$$39\scriptstyle\pm.03$$50\scriptstyle\pm.05$$36\scriptstyle\pm.03$$50\scriptstyle\pm.05$$36\scriptstyle\pm.03$$49\scriptstyle\pm.03$
+Bacadi-E .$34\scriptstyle\pm.04$$68\scriptstyle\pm.03$$53\scriptstyle\pm.08$$71\scriptstyle\pm.04$$3$03 .$72\scriptstyle\pm.04$$64\scriptstyle\pm.07$$78\scriptstyle\pm.05$$34\scriptstyle\pm.05$$71\scriptstyle\pm.04$$51\scriptstyle\pm.07$$73\scriptstyle\pm.04$
+Bacadi-M .$27\scriptstyle\pm.02$$63\scriptstyle\pm.03$$48\scriptstyle\pm.09$$71\scriptstyle\pm.04$$27\scriptstyle\pm.04$$64\scriptstyle\pm.08$$59\scriptstyle\pm.09$$7$06 .$28\scriptstyle\pm.03$$65\scriptstyle\pm.05$$45\scriptstyle\pm.08$$71\scriptstyle\pm.05$
+Dci .$59\scriptstyle\pm.03$$78\scriptstyle\pm.03$$57\scriptstyle\pm.04$$7$03 .$68\scriptstyle\pm.07$$82\scriptstyle\pm.04$$65\scriptstyle\pm.08$$84\scriptstyle\pm.05$$70\scriptstyle\pm.03$$84\scriptstyle\pm.02$$6$06 .$83\scriptstyle\pm.05$
+Mb+Ci .$54\scriptstyle\pm.11$$6$09 .$49\scriptstyle\pm.06$$62\scriptstyle\pm.04$$83\scriptstyle\pm.04$$87\scriptstyle\pm.03$$96\scriptstyle\pm.02$$96\scriptstyle\pm.02$$80\scriptstyle\pm.09$$84\scriptstyle\pm.08$$96\scriptstyle\pm.02$$96\scriptstyle\pm.02$
+Cdn .88±.04plus-or-minus.88.04.$8$04 .$95\scriptstyle\pm.02$$56\scriptstyle\pm.07$$71\scriptstyle\pm.07$$96\scriptstyle\pm.02$$98\scriptstyle\pm.01$$95\scriptstyle\pm.04$$97\scriptstyle\pm.02$$94\scriptstyle\pm.04$$98\scriptstyle\pm.02$$68\scriptstyle\pm.02$$79\scriptstyle\pm.02$
+Cdn (no $\mathcal{L}_{G}$) .$79\scriptstyle\pm.06$$90\scriptstyle\pm.04$$7$04 .$86\scriptstyle\pm.02$$94\scriptstyle\pm.01$$97\scriptstyle\pm.01$$95\scriptstyle\pm.05$$96\scriptstyle\pm.04$$87\scriptstyle\pm.02$$93\scriptstyle\pm.02$$94\scriptstyle\pm.02$$96\scriptstyle\pm.01$
+Cdn (full) .$89\scriptstyle\pm.03$$94\scriptstyle\pm.02$$98\scriptstyle\pm.02$$9$01 .$97\scriptstyle\pm.01$$9$00 $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$96\scriptstyle\pm.03$$98\scriptstyle\pm.01$$98\scriptstyle\pm.02$$9$01
+20 20 Ut-Igsp .$15\scriptstyle\pm.01$$54\scriptstyle\pm.01$$18\scriptstyle\pm.01$$65\scriptstyle\pm.02$$20\scriptstyle\pm.02$$60\scriptstyle\pm.02$$21\scriptstyle\pm.02$$67\scriptstyle\pm.02$$15\scriptstyle\pm.01$$57\scriptstyle\pm.02$$17\scriptstyle\pm.01$$65\scriptstyle\pm.01$
+Dcdi-G .$24\scriptstyle\pm.02$$49\scriptstyle\pm.02$$2$02 .$49\scriptstyle\pm.02$$2$03 .$50\scriptstyle\pm.03$$2$02 .$49\scriptstyle\pm.03$$2$02 .$50\scriptstyle\pm.04$$2$02 .$50\scriptstyle\pm.03$
+Dcdi-Dsf .$24\scriptstyle\pm.02$$50\scriptstyle\pm.02$$23\scriptstyle\pm.02$$50\scriptstyle\pm.03$$2$01 .$50\scriptstyle\pm.02$$20\scriptstyle\pm.01$$48\scriptstyle\pm.02$$2$02 .$50\scriptstyle\pm.04$$21\scriptstyle\pm.02$$47\scriptstyle\pm.03$
+Bacadi-E .$14\scriptstyle\pm.01$$61\scriptstyle\pm.03$$25\scriptstyle\pm.10$$60\scriptstyle\pm.06$$2$06 .$76\scriptstyle\pm.06$$30\scriptstyle\pm.06$$74\scriptstyle\pm.05$$15\scriptstyle\pm.02$$65\scriptstyle\pm.03$$17\scriptstyle\pm.05$$57\scriptstyle\pm.05$
+Bacadi-M .$12\scriptstyle\pm.00$$58\scriptstyle\pm.02$$17\scriptstyle\pm.03$$60\scriptstyle\pm.05$$14\scriptstyle\pm.02$$6$05 .22±.05plus-or-minus.22.05.$2$05 .$71\scriptstyle\pm.06$$12\scriptstyle\pm.01$$59\scriptstyle\pm.04$$14\scriptstyle\pm.02$$56\scriptstyle\pm.05$
+Dci .$43\scriptstyle\pm.03$$7$01 .$45\scriptstyle\pm.08$$75\scriptstyle\pm.04$$48\scriptstyle\pm.05$$76\scriptstyle\pm.03$$50\scriptstyle\pm.02$$79\scriptstyle\pm.02$$46\scriptstyle\pm.05$$75\scriptstyle\pm.02$$38\scriptstyle\pm.12$$73\scriptstyle\pm.06$
+Mb+Ci .$48\scriptstyle\pm.09$$6$05 .$57\scriptstyle\pm.08$$70\scriptstyle\pm.05$$72\scriptstyle\pm.10$$84\scriptstyle\pm.06$$92\scriptstyle\pm.04$$94\scriptstyle\pm.02$$68\scriptstyle\pm.06$$80\scriptstyle\pm.04$$89\scriptstyle\pm.03$$92\scriptstyle\pm.02$
+Cdn .$74\scriptstyle\pm.02$$90\scriptstyle\pm.01$$85\scriptstyle\pm.06$$96\scriptstyle\pm.02$$78\scriptstyle\pm.09$$92\scriptstyle\pm.04$$92\scriptstyle\pm.02$$98\scriptstyle\pm.00$$75\scriptstyle\pm.06$$91\scriptstyle\pm.03$$78\scriptstyle\pm.06$$92\scriptstyle\pm.04$
+Cdn (no $\mathcal{L}_{G}$) .$50\scriptstyle\pm.08$$73\scriptstyle\pm.07$$87\scriptstyle\pm.03$$96\scriptstyle\pm.01$$76\scriptstyle\pm.09$$90\scriptstyle\pm.05$$92\scriptstyle\pm.03$$98\scriptstyle\pm.01$$71\scriptstyle\pm.04$$89\scriptstyle\pm.02$$93\scriptstyle\pm.02$$9$00
+Cdn (full) .$75\scriptstyle\pm.03$$91\scriptstyle\pm.02$$9$01 $1.0\scriptstyle\pm.00$$79\scriptstyle\pm.08$$93\scriptstyle\pm.04$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$76\scriptstyle\pm.05$$92\scriptstyle\pm.02$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$
+20 40 Ut-Igsp .$14\scriptstyle\pm.01$$57\scriptstyle\pm.01$$14\scriptstyle\pm.01$$61\scriptstyle\pm.02$$20\scriptstyle\pm.01$$62\scriptstyle\pm.02$$19\scriptstyle\pm.01$$65\scriptstyle\pm.01$$13\scriptstyle\pm.01$$56\scriptstyle\pm.02$$14\scriptstyle\pm.00$$60\scriptstyle\pm.01$
+Dcdi-G .22±.03plus-or-minus.22.03.$2$03 .$49\scriptstyle\pm.03$$21\scriptstyle\pm.02$$48\scriptstyle\pm.03$$2$02 .$49\scriptstyle\pm.02$$2$02 .$50\scriptstyle\pm.02$$21\scriptstyle\pm.01$$48\scriptstyle\pm.02$$23\scriptstyle\pm.01$$51\scriptstyle\pm.02$
+Dcdi-Dsf .$21\scriptstyle\pm.02$$46\scriptstyle\pm.03$$21\scriptstyle\pm.02$$47\scriptstyle\pm.02$$23\scriptstyle\pm.02$$50\scriptstyle\pm.03$$21\scriptstyle\pm.01$$49\scriptstyle\pm.03$$21\scriptstyle\pm.02$$49\scriptstyle\pm.02$$21\scriptstyle\pm.03$$49\scriptstyle\pm.02$
+Bacadi-E .$23\scriptstyle\pm.05$$70\scriptstyle\pm.04$$3$10 .$67\scriptstyle\pm.04$$29\scriptstyle\pm.08$$83\scriptstyle\pm.04$$45\scriptstyle\pm.18$$78\scriptstyle\pm.06$$20\scriptstyle\pm.05$$73\scriptstyle\pm.05$$51\scriptstyle\pm.13$$79\scriptstyle\pm.04$
+Bacadi-M .$15\scriptstyle\pm.01$$64\scriptstyle\pm.02$$19\scriptstyle\pm.04$$63\scriptstyle\pm.02$$17\scriptstyle\pm.02$$70\scriptstyle\pm.05$$31\scriptstyle\pm.10$$7$07 .$14\scriptstyle\pm.01$$65\scriptstyle\pm.03$$32\scriptstyle\pm.06$$7$04
+Dci .$60\scriptstyle\pm.06$$85\scriptstyle\pm.03$$50\scriptstyle\pm.02$$7$03 .$56\scriptstyle\pm.08$$81\scriptstyle\pm.05$$59\scriptstyle\pm.06$$84\scriptstyle\pm.04$$64\scriptstyle\pm.06$$87\scriptstyle\pm.02$$6$05 .88±.03plus-or-minus.88.03.$8$03
+Mb+Ci .$53\scriptstyle\pm.09$$70\scriptstyle\pm.06$$47\scriptstyle\pm.06$$64\scriptstyle\pm.03$$78\scriptstyle\pm.04$$8$02 .$95\scriptstyle\pm.03$$96\scriptstyle\pm.02$$83\scriptstyle\pm.03$$89\scriptstyle\pm.02$$96\scriptstyle\pm.04$$97\scriptstyle\pm.03$
+Cdn .$86\scriptstyle\pm.04$$96\scriptstyle\pm.02$$79\scriptstyle\pm.09$$93\scriptstyle\pm.05$$8$03 .$97\scriptstyle\pm.01$$96\scriptstyle\pm.02$$9$00 .88±.03plus-or-minus.88.03.$8$03 .$97\scriptstyle\pm.01$$71\scriptstyle\pm.05$$87\scriptstyle\pm.04$
+Cdn (no $\mathcal{L}_{G}$) .$69\scriptstyle\pm.09$$85\scriptstyle\pm.07$$80\scriptstyle\pm.06$$94\scriptstyle\pm.03$$84\scriptstyle\pm.03$$94\scriptstyle\pm.01$$95\scriptstyle\pm.02$$9$00 .$84\scriptstyle\pm.05$$93\scriptstyle\pm.03$$91\scriptstyle\pm.02$$98\scriptstyle\pm.00$
+Cdn (full) .$89\scriptstyle\pm.06$$97\scriptstyle\pm.02$$9$01 $1.0\scriptstyle\pm.00$$89\scriptstyle\pm.03$$97\scriptstyle\pm.01$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$$93\scriptstyle\pm.03$$98\scriptstyle\pm.01$ $1.0\scriptstyle\pm.00$ $1.0\scriptstyle\pm.00$

### v4

@@
 This approach consistently outperforms baselines for perturbation modeling on seven single-cell transcriptomics datasets.
-We also demonstrate significant improvements over current causal discovery methods for predicting soft and hard intervention targets across a variety of synthetic data.
+We also demonstrate significant improvements over current causal discovery methods for predicting soft and hard intervention targets across a variety of synthetic data.111 Code is available at https://github.com/rmwu/cdn
 perturbation, transcriptomics, causality
@@
 Machine learning methods have primarily been designed to infer the effects of perturbations on cells, with the goal of generalizing to unseen perturbations (Roohani et al., 2023), or unseen cell distributions (Lotfollahi et al., 2023; Bunne et al., 2023).
-In principle, these models can be used within an active learning framework to discover intervention targets (Huang et al., 2023; Zhang et al., 2023a).
+In principle, these models can be used within an active learning framework to discover intervention targets (Huang et al., 2024; Zhang et al., 2023a).
 However, the number of candidate sets scales exponentially with the number of variables under consideration, and there are limited training data for combinatorial perturbations ($<150$ pairs in Norman et al.
@@
 The pair of graphs is input to an attention-based classifier, which predicts whether each variable was subject to intervention.
-To address the challenges of data noise and sparsity, and to simulate the structure of biological interventions, these models are trained jointly in a supervised (amortized) framework (Ke et al., 2022; Petersen et al., 2023), over thousands of synthetic or real datasets.
+To address the challenges of data noise and sparsity, and to simulate the structure of biological interventions, these models are trained jointly in a supervised (amortized) framework (Ke et al., 2023; Petersen et al., 2023), over thousands of synthetic or real datasets.
 We evaluate Cdn on real transcriptomic data and synthetic settings.
-Cdn outperforms the state-of-the-art in perturbation modeling (deep learning and statistical approaches) on the five largest Perturb-seq datasets at the time of publication  (Replogle et al., 2022; Nadig et al., 2024), without using any external knowledge.
+Cdn outperforms the state-of-the-art in perturbation modeling (deep learning and statistical approaches), evaluated on the five largest Perturb-seq datasets at the time of publication  (Replogle et al., 2022; Nadig et al., 2024) without using any external knowledge.
 Furthermore, Cdn generalizes with minimal performance drop to unseen cell lines, which have different supports (genes), causal mechanisms (gene regulatory networks), and data distributions.
-Finally, on synthetic settings, Cdn outperforms causal discovery approaches for estimating unknown intervention targets.
+On synthetic settings, Cdn outperforms causal discovery approaches for estimating unknown intervention targets.
 Figure 1: (A) Biological applications.
@@
 These three objects may share similarities across cell lines, as certain genes or pathways are essential for survival.
-However, the empirical discrepancy is quite large between experiments.
+However, the empirical discrepancy can be quite large between experiments.
 For example, Nadig et al.
 (2024) reports that the median correlation in log-fold change of genes under the same perturbation, across two experiments in the same cell line, by the same lab, was 0.16.
-There are two directions for modeling perturbations.
+Early efforts to map the space of cells measured the distribution of gene levels as a surrogate for cell state (Lamb et al., 2006).
+Diseases, drugs, or genetic perturbations that induce similar changes in this distribution were thought to act through similar mechanisms.
+However, it is intractable to map all combinations of starting cell states and perturbations through experiment alone.
+This bottleneck has led to two directions in the in-silico modeling of perturbations.
 To reduce experimental costs, which scale with the number of cell lines and perturbations, one option is to infer the post-intervention distribution of cells.
@@
 unseen cell lines or patients.
-Alternatively, models like GEARS (Roohani et al., 2023), AttentionPert (Bai et al., 2024), and GIM (Schneider et al., 2024) are designed to generalize to unseen perturbations.
-These approaches do not explicitly predict targets, but if the space of perturbations $\mathcal{I}$ can be efficiently enumerated, these models are useful for comparing different $I$ (Zhang et al., 2023a; Huang et al., 2023).
+Alternatively, models like GEARS (Roohani et al., 2023), graphVCI (Wu et al., 2023), AttentionPert (Bai et al., 2024), and GIM (Schneider et al., 2024) infer the effects of unseen perturbations.
+This setting is also a common application for single cell foundation models (Theodoris et al., 2023; Cui et al., 2024; Hao et al., 2024).
+These latter methods generalize by modeling the relationships between seen and unseen perturbations.
+For example, GEARS leverages literature-derived relationships in the Gene Ontology knowledge graph (Ashburner et al., 2000), and graphVCI learns to refine the gene regulatory relationships implied by ATAC-seq data (Grandi et al., 2022).
+Approaches for inferring the effects of interventions do not explicitly predict targets, but if the space of perturbations $\mathcal{I}$ can be efficiently enumerated, these models are useful for comparing different $I$ (Zhang et al., 2023a; Huang et al., 2024).
 For example, given an oracle $\phi:D_{\text{obs}},I\to\tilde{P}$, the $I^{*}$ associated with $\tilde{P}^{*}$ is
@@
 for some dissimilarity $d$.
-However, recent works have reported that the predictive ability of current models does not exceed that of naive baselines (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024), reflecting potential barriers to this approach.
+However, recent works have reported that the predictive ability of current models does not exceed that of naive baselines (Kernfeld et al., 2023; Ahlmann-Eltze et al., 2024; Märtens et al., 2024), reflecting that phenotype prediction is currently unsolved.
 Perturbation datasets can also be used to train models for predicting targets of interest, as is the focus of this work.
@@
 pathways), in which the targets are involved.
-To draw this connection, current approaches use external information such as knowledge graphs (Cosgrove et al., 2008; Gonzalez et al., 2024) and natural language (Roohani et al., 2024).
+To draw this connection, current approaches use external information such as knowledge graphs (Cosgrove et al., 2008; Gonzalez et al., 2024) and natural language (Roohani et al., 2025).
 However, these information are collected from highly inhomogenous sources (Ashburner et al., 2000), and it is difficult to quantify their correctness and completeness in each new setting.
-In contrast to these approaches, we choose to infer the mechanistic structure directly from data.
+An alternative is to infer the mechanistic structure directly from data, either explicitly (di Bernardo et al., 2005) or as a model parameter (Noh & Gunawan, 2016).
+In this work, we follow the latter strategy, while allowing for a richer (non-linear) class of relationships.
 2.3 Causal discovery
@@
 However, the former strictly assumes linearity, while the latter requires at least as many environments as $\|I\|$ (we only consider before and after perturbation).
-Finally, Dci (Belyaeva et al., 2021) infers edge-level differences between linear Gaussian models, rather than identifying the full causal structure.
+Finally, several works aim to infer differences between causal mechanisms (edges), rather than identifying the full causal structure (Wang et al., 2018; Belyaeva et al., 2021; Chen et al., 2023; Malik et al., 2024).
 Our work is also motivated by the idea that differences between summary statistics (e.g.
@@
 While the majority of causal discovery algorithms must be fit or run from scratch for each data distribution, amortized causal discovery algorithms have addressed this task as a supervised machine learning problem.
-Using large numbers of synthetic datasets, generated by synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Li et al., 2020; Ke et al., 2022; Lorch et al., 2022; Petersen et al., 2023; Wu et al., 2024).
+Using large numbers of synthetic datasets, generated by synthetic graphs, a neural network is trained to map $D$ directly to $G$ (Li et al., 2020; Ke et al., 2023; Lorch et al., 2022; Petersen et al., 2023; Wu et al., 2025).
 Compared to traditional approaches, these algorithms trade explicit assumptions for implicit characteristics of the data simulation process (Montagna et al., 2024).
@@
 In principle, this module can be implemented with any data-efficient causal discovery model.
-Here, we based the architecture on Sea (Wu et al., 2024), as its runtime complexity does not explicitly depend on the number of samples (which can be large for $D_{\text{obs}}$).
+Here, we based the architecture on Sea (Wu et al., 2025), as its runtime complexity does not explicitly depend on the number of samples (which can be large for $D_{\text{obs}}$).
 Input featurization  We represent dataset $D$ through two types of statistics.
-First, we compute global statistics $\rho\in\mathbb{R}^{N\times N}$, like pairwise correlation, over all variables.
-Second, we obtain estimates of local causal structure by running classical causal discovery algorithms like FCI (Spirtes et al., 1995) over small subsets of variables.
-We denote the set of estimates as $E^{\prime}\in\mathcal{E}^{k\times k\times T}$, where $\mathcal{E}$ is the space of possible edge types (e.g.
-,$\leftrightarrow$,-$\rightarrow$ , - , →), and $k,T$ are hyperparameters that trade off accuracy vs.
-runtime.
+First, we compute global statistics $\rho\in\mathbb{R}^{N\times N}$ (pairwise correlation) over all variables.
+This statistic is an inexpensive way for the model to filter out edges that are not likely to occur.
+Second, we obtain estimates of local causal structure to orient edges, by running classical causal discovery algorithms over small subsets of variables.
+Specifically, we run the FCI algorithm $T$ times over subsets of $k$ variables (Spirtes et al., 1995), which are sampled proportional to their pairwise correlation.
+We denote the set of estimates as $E^{\prime}\in\mathcal{E}^{k\times k\times T}$, where $\mathcal{E}$ is the space of possible edge types inferred by FCI (e.g.
+,$\leftrightarrow$,-$\rightarrow$ , - , →).
+The size and number of local estimates, $k$ and $T$, are hyperparameters that trade off accuracy vs.
+runtime (Section 3.3).
 To compare $\rho$ and $E^{\prime}$ in the same space, we linearly project elements of $\rho$ from $\mathbb{R}$ to $\mathbb{R}^{d}$, and map categorical edge types into $\mathbb{R}^{d}$ using a learned embedding, i.e.
@@
 This design takes after Transformer models for higher dimensional data in other domains, e.g.
-Axial Transformers for videos (Ho et al., 2020) and MSA Transformers for aligned protein sequences (Rao et al., 2021).
+Axial Transformers for videos (Ho et al., 2019) and MSA Transformers for aligned protein sequences (Rao et al., 2021).
 A key difference is the graph modality: when making edge or node-level predictions, the model should be equivariant to node labeling.
@@
 Details are available in Section 4.2 and Appendix D.2.
-Model parameters  Marginal estimates $E^{\prime}$ are computed using the classic FCI algorithm (Spirtes et al., 1995), and global statistic $\rho$ is pairwise correlation, as it is fast to compute and numerically stable.
+Model parameters  Marginal estimates $E^{\prime}$ are computed using the classic FCI algorithm (Spirtes et al., 1995) with $k=5$ and $T=100$, and global statistic $\rho$ is pairwise correlation, as it is fast to compute and numerically stable.
 We swept over the number of differential network layers (Figure 4) on synthetic data, and we used 3 layers for $h_{\text{cat}}$ and 2 layers for $h_{\text{diff}}$.
@@
 4 Experiments
-We evaluated Cdn on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1).
+While our method is intended to be used in cases where true targets are unknown, we can only evaluate on datasets where the target is known.
+Thus, we demonstrate that Cdn can recover perturbation targets on seven transcriptomics datasets, with comparisons to state-of-the-art models for these applications (Section 4.1).
 For completeness, we also benchmarked Cdn against multiple causal discovery algorithms for unknown interventions in variety of controlled settings (Section 4.2).
 4.1 Biological experiments
-Datasets  We evaluate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al.
+Datasets  We validate Cdn on five Perturb-seq (Dixit et al., 2016) datasets (genetic perturbations) from Replogle et al.
 (2022) and Nadig et al.
@@
 Thus, we propose the following metrics.
-  • We compute the rank of the true target, normalized by the number of candidate genes. Rank ranges from 0 (worst, bottom of list) to 1 (best, top of list).
+  • We order genes based on the similarity of their predicted effect to the ground truth target effect, or based on their predicted probability of being a target.
+The rank is the index of the true target, normalized by the number of candidate genes.
+Rank ranges from 0 (worst, bottom of list) to 1 (best, top of list).
   • We measure the similarity between the perturbation effect (log-fold change) of the top 1 predicted target and the true target with the Spearman rank correlation $\rho$ and the Pearson correlation $r$.
 These range from -1 to 1, where 1 is best.
-  • To emulate “virtual screening,” we plot recall at p: the fraction of targets recovered within the $p\in(0,1)$ candidates. Recall ranges from 0 to 1, where 1 is best.
+  • To emulate “virtual screening,” we plot recall at p: the fraction of targets recovered within the $p\in(0,1)$ candidates.
+Recall ranges from 0 to 1, where 1 is best.
+Rank is equivalent to the expected recall at $p$.
 Baselines  We compare to both algorithms that infer the effects of perturbations on cells (used as scoring functions, Equation 3), and algorithms that directly predict targets.
@@
 […]
+exclude scale $0.32\scriptstyle\pm.07$ $0.64\scriptstyle\pm.04$ $0.34\scriptstyle\pm.09$ $0.65\scriptstyle\pm.08$ $0.06\scriptstyle\pm.02$ $0.63\scriptstyle\pm.02$ $0.08\scriptstyle\pm.02$ $0.61\scriptstyle\pm.06$
+exclude scale 0.44±.04plus-or-minus0.44.040.$4$04 $0.92\scriptstyle\pm.01$ $0.43\scriptstyle\pm.09$ $0.90\scriptstyle\pm.03$ $0.06\scriptstyle\pm.01$ $0.72\scriptstyle\pm.02$ $0.10\scriptstyle\pm.02$ $0.73\scriptstyle\pm.02$
 Baselines  We compare against discrete and continuous causal discovery algorithms for unknown interventions.
@@
 In the case of Dcdi and BaCaDI, this may be because it is hard to select a single sparsity threshold for varying sizes of intervention sets, and to balance the sparsity regularizer with the generative modeling objective.
+5 Discussion
+This work introduces causal differential networks (Cdn), which focuses on the application of predicting perturbation targets for efficient experimental design.
+On a high level, our goal is to lower the cost and scale of experiments, by reducing the search space.
+Our key technical contribution is the idea of using supervised causal discovery as a pretraining objective, to obtain dataset representations useful for downstream tasks like inferring targets.
+Specifically, Cdn trains a causal graph learner that maps a pair of observational and interventional datasets to their data-generating mechanisms, and a differential network classifier to identify variables whose conditional independencies have changed.
+We demonstrated that our approach outperforms state-of-the-art methods in perturbation modeling and causal discovery, across seven transcriptomics datasets and a variety of synthetic settings.
+Overall, these results highlight the potential of our method to design more efficient experiments, by reducing the search space of perturbations.
+There are several limitations of this work and directions for future research.
+First, our implementation does not address cyclic, feedback relationships or time-resolved, dynamic systems – both of which are common in biology.
+Modeling these cases may require data simulation schemes that can sample from the steady-state of a cyclic system, or architectures that directly predict the parameters of a system of stochastic differential equations (Lorch et al., 2024).
+In addition, current supervised causal discovery algorithms, which we use for the causal graph learner, tend to assume causal sufficiency, i.e.
+that there are no unseen confounders.
+This assumption is unrealistic in biology, as it is not physically possible to measure all relevant variables.
+Thus, there is an opportunity to extend this work for the latent confounding setting.
+Finally, since it is difficult to simulate biology, there will be an inevitable domain shift between synthetic and real data.
+However, there does exist plentiful unlabeled single-cell data (Consortium* et al., 2022), which could be used with self-supervised objectives to bridge the gap (Sun et al., 2020).
+In conclusion, we hope that this work will enable efficient experimental design and interpretation, as well as future machine learning efforts towards these tasks.
 Impact statement
@@
 (2025).
-The causal structure learner closely resembles the architecture of Sea (Wu et al., 2024), while the differential network is slightly different (since the input space is different).
+The causal structure learner closely resembles the architecture of Sea (Wu et al., 2025), while the differential network is slightly different (since the input space is different).
 Briefly, the input to the causal structure learner is a matrix of shape $N\times N\times T\times d$.
@@
 Cdn took around 6 hours to train on synthetic data, and 1 hour to finetune on real data.
-Unless otherwise noted, the causal structure learner was initialized from the Sea inverse covariance, FCI pretrained weights (Wu et al., 2024).
+Unless otherwise noted, the causal structure learner was initialized from the Sea inverse covariance, FCI pretrained weights (Wu et al., 2025).
 Figure 4: Hyperparameter search: number of differential network layers. We selected 3 layers based on validation metrics.
@@
 In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$.
-The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2024).
+The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2025).
 We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
@@
 D.1 Biological datasets
+Data processing
 We converted all single cell datasets to log-normalized, transcripts per 10,000 UMIs.
@@
 Figure 8 plots the full distribution of number of cells and DEGs per perturbation.
+Top differentially expressed genes
+Here, we limited our analysis to the top 1000 DEGs.
+In the literature, it is quite common to restrict analysis to subsets of genes, though the threshold and criteria may vary (Luecken & Theis, 2019).
+Selecting the top 1000 genes (based on differential expression p-value) is reasonably permissive, since these genes are selected per perturbation.
+For a sense of scale, Nadig et al.
+(2024) writes:
+“In a genome-scale Perturb-seq screen, we find that a typical gene perturbation affects an estimated 45 genes, whereas a typical essential gene perturbation affects over 500 genes.”
+A classic approach for identifying drug targets argues that targets are not highly differentially expressed, but network-based approaches are useful for predicting them from gene expression (Isik et al., 2015).This work also samples 1000 candidate targets as decoys.
+Gene co-expression
 Figures 6 and 7 depict the differences in gene co-expression matrices between control and perturbed cells.
 Genetic perturbations tend to have much clearer phenotypes compared to chemical perturbations, whose effects are more diffuse.
+Sci-Plex drug targets
 The full names of the Sci-Plex drugs and their targets are as follows. These targets were collated by the authors (McFaline-Figueroa et al., 2024), and are each are well-documented in the literature.
@@
 Perturb-seq uncertainty quantification
-Due to space limitations, we report uncertainty estimates in Table 7.
-Multiple baselines produce deterministic results (Linear, Dge, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells.
+Due to space limitations, we report uncertainty estimates in Tables 7 and 8.
+Multiple baselines produce deterministic results (Linear, Dge, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells (Table 7).
 Specifically, for each perturbation with $M$ cells, we sample
@@
 cells uniformly at random, repeated 5 times.
+We also consider the robustness of our model to samplings of the test set (Table 8).
+Typically, the largest cluster (assigned to train, e.g.
+nearly 1/3 of K562) contains weak perturbations with minimal phenotype, so different clusterings don’t introduce significant heterogeneity.
+Instead, we can partition all perturbations uniformly into 5 sets for the unseen cell line setting, across which the variation in results is minimal.
 Perturb-seq trivial perturbations
@@
 Therefore, to assess whether our model can generalize, we chose to hold out a class of soft intervention for our main results in Table 3.
+Table 11 reports results on larger graphs.
+There is some drop in performance, especially Polynomial.
+This could be attributed to error propagation from the causal graph predictor, which performs less well (at predicting graphs) on larger graphs, and could potentially be ameliorated by finetuning on larger synthetic data (similar to how Perturb-seq models were finetuned to be larger).
+Linear + Soft remains quite good, perhaps because the theory (Appendix C) suggests that the summary statistics are sufficient for predicting targets, without the graph prediction.
 Table 12 reports results on all synthetic test datasets, averaging over all interventions for each dataset. We also include the “full” version of Cdn, i.e. trained on all mechanisms, in this analysis.
@@
 Cdn $\textbf{0.93}\scriptstyle\pm.00$ $\textbf{0.67}\scriptstyle\pm.01$ $\textbf{0.73}\scriptstyle\pm.01$ $\textbf{0.96}\scriptstyle\pm.01$ $\textbf{0.83}\scriptstyle\pm.04$ $\textbf{0.85}\scriptstyle\pm.03$
+Table 8: To quantify the uncertainty across draws of the test set (perturbations, instead of cells), we partition all perturbations uniformly into 5 sets for the unseen cell line setting on K562 genome-wide.
+The variation across test sets is minimal.
+Test Split rank $\rho$ $r$ Recall@1 Recall@5 Recall@20
+Original ($n=678$) 0.974 0.915 0.921 0.699 0.900 0.957
+New splits ($n=353,\times 5$) 0.973 ±$\scriptstyle\pm 0.009$ 0.917 ±$\scriptstyle\pm 0.017$ 0.923 ±$\scriptstyle\pm 0.016$ 0.696 ±$\scriptstyle\pm 0.031$ 0.908 ±$\scriptstyle\pm 0.032$ 0.961 ±$\scriptstyle\pm 0.021$
 Table 9: Results on 5 Perturb-seq datasets, trivial perturbations (intended target has largest log-fold change, though not necessarily smallest adjusted p-value).
@@
 exclude scale $0.83\scriptstyle\pm.10$ $0.95\scriptstyle\pm.05$ $0.76\scriptstyle\pm.07$ $0.89\scriptstyle\pm.04$ $0.97\scriptstyle\pm.03$ $1.0\scriptstyle\pm.00$ $0.96\scriptstyle\pm.02$ 0.99±.01plus-or-minus0.99.010.$9$01
+Table 11: While the majority of classic causal discovery models struggle to scale to larger graphs, we show that Cdn still performs reasonably on $N=100,E=200$ graphs.
+AUC reported over 10 i.i.d.
+graphs.
+Intervention Linear (1) Linear (3) Sigmoid (1) Sigmoid (3) Polynomial (1) Polynomial (3)
+Hard .$64\pm.04$$65\pm.08$$63\pm.02$$61\pm.06$$50\pm.01$$49\pm.03$
+Soft .$92\pm.01$$90\pm.03$$72\pm.02$$73\pm.02$$53\pm.03$$52\pm.05$
 Table 12: Intervention target prediction results on synthetic datasets, extended results.